### v1

# CodeCrash: Stress Testing LLM Reasoning under Structural and Semantic Perturbations

Abstract
Large Language Models (LLMs) have recently showcased strong capabilities in code-related tasks, yet their robustness in code comprehension and reasoning remains underexplored. In this paper, we present CodeCrash, a unified benchmark that evaluates LLM robustness under code structural and textual distraction perturbations, applied to two established benchmarks—CRUXEval and LiveCodeBench—across both input and output prediction tasks. We evaluate seventeen LLMs using direct and Chain-of-Thought inference to systematically analyze their robustness, identify primary reasons for performance degradation, and highlight failure modes. Our findings reveal the fragility of LLMs under structural noise and the inherent reliance on natural language cues, highlighting critical robustness issues of LLMs in code execution and understanding. Additionally, we examine three Large Reasoning Models (LRMs) and discover the severe vulnerability of self-reflective reasoning mechanisms that lead to reasoning collapse. CodeCrash provides a principled framework for stress-testing LLMs in code understanding, offering actionable directions for future evaluation and benchmarking. The code of CodeCrash and the robustness leaderboard are publicly available at https://donaldlamnl.github.io/CodeCrash/.
1 Introduction
Large Language Models (LLMs), exemplified by GPT-4 OpenAI et al. (2024), and the DeepSeek series DeepSeek-AI et al. (2025b, a), have recently exhibited remarkable performance across diverse code-related tasks, including code generation Zan et al. (2023); Xiao et al. (2024), program repair Fan et al. (2023); Liu et al. (2025), program testing Deng et al. (2023); Kang et al. (2023), and test case generation Steenhoek et al. (2025); Takerngsaksiri et al. (2025). Recent studies Wang et al. (2022); Mastropaolo et al. (2023); Chen et al. (2024a) have primarily evaluated LLM robustness under prompt-level perturbation, focusing on variations in task descriptions for code generation. While insightful for prompt engineering, these studies overlook robustness in deeper code comprehension tasks, such as understanding execution logic, structural reasoning, and program semantics.
In practical software development, LLMs are increasingly integrated into IDEs like GitHub Copilot [^1] https://github.com/features/copilot for tasks including code completion, automatic bug fixing, and test generation. Those integrations underscore the potential of LLMs to automate aspects of software development, improving code quality and developer productivity Zhang et al. (2023). Nevertheless, real-world production codebases are often complex and messy, frequently containing outdated comments, placeholder code, ambiguous identifiers, and non-functional segments, which introduce risk if LLMs overly depend on superficial cues rather than deep logical understanding.
CRUXEval (CRUXGu et al. (2024) introduces input and code execution tasks (i.e., output prediction) to evaluate LLMs’ comprehension and reasoning about code behavior. Building upon this foundation, we systematically evaluate LLM robustness under controlled code perturbations designed to challenge models’ reasoning capabilities beyond surface-level situations. Inspired by the program structure-consistent mutations proposed in CCTest Li et al. (2023), we extend this concept into a comprehensive robustness evaluation framework. Our unified benchmark, CodeCrash, systematically evaluates LLM robustness by applying two complementary perturbation families: I. Code Structural Perturbations, which introduce logically equivalent but structurally complicated code modifications, such as renaming entities, reformatting conditional expressions, and inserting non-functional “garbage” code segments; and II. Textual Distraction Perturbations, which insert misleading natural language (NL) cues that contradict actual program logic, including embedding misleading NL comments or statements and incorrect output hints.
In this paper, we conduct extensive perturbation experiments on two established benchmarks: CRUX and LiveCodeBench (LCBJain et al. (2025). We evaluated the robustness of seventeen LLMs across distinct perturbations to assess their deep understanding of code behavior and functionality in input prediction and code execution tasks. Additionally, we apply Chain-of-Thought (CoT) prompting Kojima et al. (2022) to explore how explicit step-by-step reasoning enhances model performance and robustness in code comprehension. We further evaluate three state-of-the-art Large Reasoning Models (LRMs) with powerful internal reasoning processes, uncovering significant vulnerabilities arising from their self-reflection mechanisms. Our key findings reveal that:
  • LLMs struggle to strictly follow actual code behavior under structural perturbations, frequently distracted by irrelevant or non-functional segments.
  • LLMs inherently prioritize NL cues over actual program logic, heavily relying on identifiers, comments, and embedded messages, adopting superficial comprehension and shortcut reasoning.
  • Explicit step-by-step reasoning boosts performance and robustness; however, internal reasoning mechanisms may trigger severe overthinking, infinite self-reflection, and re-verification.
Our contributions include:
  • We introduce a comprehensive perturbation framework CodeCrash encompassing structural complexity and textual distractions to evaluate the code-level robustness of LLMs.
  • We provide novel insights on evaluating and interpreting code reasoning in LLMs, revealing that they exhibit superficial NL dependence and demonstrating how plausible but incorrect hints trigger model collapsing.
2 Related Work
Researchers have significantly contributed to developing benchmarks for evaluating LLM capabilities on code-related tasks. Early benchmarks such as HumanEval Chen et al. (2021), MBPP Austin et al. (2021a), APPS Hendrycks et al. (2021), and HumanEval+ Liu et al. (2023b) primarily focused on assessing code generation performance. More recent efforts have shifted toward practical software engineering and realistic coding scenarios. For instance, SWE-bench Jimenez et al. (2024), InfiBench Li et al. (2024), and LiveCodeBench Jain et al. (2025) evaluate LLMs on complex tasks like multi-file changes, Stack Overflow-style questions, and real-world programming problems.
Meanwhile, a growing number of works investigate LLMs in code execution. Early studies Austin et al. (2021b); Nye et al. (2021) indicated that LLMs face challenges in accurately executing code but can benefit from intermediate reasoning steps. Subsequent research has utilized various strategies to enhance execution performance, including execution-informed pretraining Liu et al. (2023a), CoT prompting with trace supervision Ni et al. (2024), and iterative instruction prompting Lyu et al. (2024). CruxEval Gu et al. (2024) formalized input and output prediction tasks specifically for evaluating LLMs’ code understanding and reasoning, while LiveCodeBench extended these assessments to realistic programs written by humans. REval Chen et al. (2024b) and CACP Hooda et al. (2024) uncovered inconsistencies and concept-level misunderstandings of LLMs in code through intermediate reasoning analysis in code execution. CodeScore Dong et al. (2024), LEVER Ni et al. (2023), and XCODEEval Khan et al. (2024) have leveraged code execution to improve model evaluation methodologies and enhance the performance of downstream tasks.
Robustness evaluation is another research direction in assessing LLMs for code. Prior studies focused on prompt-level variations in code generation, examining how semantically equivalent NL descriptions impact model performance Mastropaolo et al. (2023); Chen et al. (2024a); Zhuo et al. (2023). Apart from NL perturbations, researchers have investigated model robustness under programming language perturbations. For example, CCTest Li et al. (2023) introduced structural mutations, CREAM Gao et al. (2023) employed counterfactual reasoning, and ReCode Wang et al. (2022) applied semantic-preserving transformations, advancing the robustness in code comprehension and completion. BigCodeBench Zhuo et al. (2025) further explored generalization under complex function composition and tool usage. Furthermore, Chen et al. (2025); Cuadron et al. (2025) and Kumar et al. (2025) have recently revealed that LRMs frequently suffer from overthinking and produce unnecessary analysis and unstable thoughts in math-related works.
3 Methodology
This section outlines our benchmark methodology, including dataset selection, benchmark pipeline, perturbation strategies, and experimental settings.
3.1 Benchmark Tasks and Datasets
We evaluate two benchmarks: CRUX Gu et al. (2024) and LCB Jain et al. (2025). CRUX contains 800 synthetic problems generated by Code-LLaMA-34B, each featuring concise Python snippets with generic function names f and a corresponding input-output pair for input prediction and code execution tasks. In contrast, LCB contains 479 real-world coding problems sourced from 85 human-submitted LeetCode solutions, focusing solely on code execution. While both datasets avoid excessive memory or runtime demands, LCB problems exhibit more complex structures, such as nested functions, decorators, and external libraries, requiring more computations.
3.2 Benchmark Execution Pipeline
Figure 1: Overview of CodeCrash pipeline. Code samples are first standardized into a clean vanilla version. Perturbations are visualized with yellow highlights. Each perturbed code is validated and executed before passing to LLMs.
Our benchmark follows a unified evaluation pipeline (Figure 1) across all perturbation types. Each original code snippet is initially processed via Abstract Syntax Tree (AST) parsing and unparsing to remove comments and standardize formatting, producing a clean baseline “vanilla” version (VAN). Subsequently, we apply a selected perturbation and regenerate the code using AST unparsing to maintain formatting consistency (except for comment-based perturbations). Each perturbed code snippet is executed with the provided inputs to ensure syntactic validity and functional correctness. The validated perturbed code is then provided to LLMs, and we record all execution results and evaluate them using the original evaluation methods from the dataset to compute the PASS@1 score.
3.3 Perturbation Strategies
We design two categories of perturbations to evaluate model robustness comprehensively: (1) Code Structural Perturbations, which modify code structure without affecting functionality, and (2) Textual Distraction Perturbations, which inject misleading NL cues and hints. Detailed examples and principles are provided in Appendix B.
3.3.1 Code Structural Perturbations
(1) Renaming Entities (REN).
[…]
(2) Reformatting Conditional Expressions (RTF).
We rewrite conditional expressions within “if” and “while” statements into logically equivalent but structurally complex and indirect forms. We use 16 predefined transformation templates to reformat comparison, boolean, and other conditional expressions. This perturbation evaluates the LLMs’ ability to maintain logical consistency and reasoning under increased complexity.
(3) Inserting Garbage Code (GBC).
We insert syntactically valid but functionally irrelevant “garbage” code segments to challenge the models’ robustness against code structural distractions. Three specific types include: (1) repeated global variable declarations mirroring local identifiers, (2) non-executable conditional dead branches, and (3) uninvoked infinite recursive function definitions.
(4) Aggregating Structural Perturbations (ALL).
We comprehensively aggregate all structural perturbations (REN, RTF, GBC) together and create a highly adversarial scenario, thoroughly challenging LLM robustness across multiple structural perturbation dimensions simultaneously.
3.3.2 Textual Distraction Perturbations
(1) Inserting Misleading Descriptive Comments (MDC).
We insert misleading NL comments for eight key AST nodes (e.g., function definitions, returns, loops, conditionals) to evaluate LLM robustness against textual cues and contradictory information. We initially instruct GPT-4o to generate those messages and manually filter them to ensure they are generic, explicitly incorrect, and contradict the actual code logic. Additionally, we introduce a probability parameter p to vary comment density and analyze its impact and validity on model robustness in § 5.1.
(2) Inserting Misleading Print Statements (MPS).
Similar to MDC perturbation, we inject misleading NL messages as executable print statements. Unlike comments that AST-based heuristics can easily filter out, print statements are logically irrelevant to functionality and should not be perceived as semantic guides. Allowing us to directly compare the behavior with MDC and distinguish models’ NL reliance in different embedding formats.
(3) Inserting Misleading Hint Comments (MHC).
We instruct GPT-4o to generate logically plausible but incorrect hints that contradict the actual program behavior and the ground truth. After verifying their incorrectness, they are strategically inserted into function definitions or return statements using AST parsing.
3.4 Experiment Settings
Model Selection.
We evaluate 17 models of varying sizes and versions, including both open-source and closed-source LLMs: GPT family (GPT-4o-Mini, GPT-4o), Anthropic Claude (Claude-3.5-Haiku, 3.5-Sonnet), Google Gemini (Gemini-1.5-Flash, 2.0-Flash, 1.5-Pro), DeepSeek-V3, Alibaba Qwen (Qwen2.5-7B-Ins, 14B-Ins, 32B-Ins, 72B-Ins, 32B-Coder), and Meta LLaMA (LLaMA-3.1-8B-Ins, 70B-Ins, 405B-Ins, LLaMA-3.3-70B). Additionally, we assess three cutting-edge LRMs: o3-mini (low and high reasoning efforts), DeepSeek-R1, and QwQ-32B.
Model Configurations.
Following the settings in LCB, we use nucleus sampling (temperature=0.2, top-p=0.95). We set the maximum tokens per response at 200 for direct inference and 2,000 for CoT inference, discarding excessively long outputs. For o3-mini, we set the maximum completion tokens at 20,000 to avoid verbose reasoning. Due to resource constraints, we generate three candidate answers for direct inference and one candidate for CoT inference per query. We adopt prompt formats similar CRUX, employing 2-shot prompting for direct inference and 1-shot CoT prompting (details provided in Appendix A). Primary evaluation uses PASS@1 accuracy Chen et al. (2021), reporting both absolute and relative degradation metrics.
4 Experimental Results
This section presents a comprehensive analysis of model robustness under the perturbation strategies introduced in §3.3. We assess the robustness of LLMs across different perturbation types, model scales, reasoning modes, and prediction tasks.
4.1 Impact on Output Prediction
We first evaluate LLM robustness in output prediction tasks. Table 1 summarizes the absolute PASS@1 degradation for all models under direct inference across both datasets, highlighting how each perturbation affects model performance.
Table 1: Absolute PASS@1 degradation of LLMs under all perturbation types for output prediction tasks across CRUX and LCB in direct inference (N=3) [TO RELATIVE SCORE].
Model Series Model Name VAN REN RTF GBC ALL MDC MPS MHC
Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
GPT-4 Omni GPT-4o-Mini $57.29$ $53.24$ $-0.62$ $-9.67$ $-1.79$ $-0.97$ $-10.92$ $-5.50$ $-15.96$ $-16.98$ $-18.67$ $-21.50$ $-13.29$ $-14.89$ $-6.62$ $-23.59$
GPT-4o $71.33$ $64.51$ $-2.71$ $-9.60$ $-3.38$ $-4.59$ $-9.21$ $-3.20$ $-10.71$ $-18.30$ $-9.96$ $-18.79$ $-12.29$ $-15.66$ $-4.54$ $-17.19$
Claude Series Claude-3.5-Haiku-20241022 $57.42$ $58.25$ $-3.96$ $-9.32$ $-1.79$ $+0.35$ $-7.29$ $-4.52$ $-13.79$ $-21.50$ $-7.92$ $-5.85$ $-6.46$ $-4.66$ $-15.79$ $-31.04$
Claude-3.5-Sonnet-20241022 $71.46$ $73.83$ $-1.67$ $-16.77$ $-0.92$ $+1.18$ $-6.88$ $-4.94$ $-10.54$ $-25.19$ $-5.79$ $-7.10$ $-6.17$ $-7.86$ $-10.29$ $-32.08$
Gemini Series Gemini-1.5-Flash $56.21$ $44.26$ $-3.54$ $-0.42$ $-1.17$ $-1.04$ $-4.58$ $-0.35$ $-10.33$ $-5.64$ $-8.29$ $-14.27$ $-16.00$ $-17.88$ $-4.12$ $-9.81$
Gemini-2.0-Flash $65.04$ $65.97$ $-2.33$ $-18.09$ $-2.08$ $-0.77$ $-15.50$ $-8.77$ $-21.63$ $-27.77$ $-17.42$ $-34.24$ $-21.54$ $-33.19$ $-5.38$ $-12.87$
Gemini-1.5-Pro-002 $67.37$ $56.16$ $-2.92$ $-3.41$ $-2.50$ $-4.73$ $-5.62$ $-3.83$ $-13.42$ $-13.01$ $-15.50$ $-18.51$ $-14.62$ $-18.44$ $-6.75$ $-18.23$
DeepSeek DeepSeek-V3 $67.87$ $67.78$ $-2.67$ $-16.35$ $-1.33$ $-3.76$ $-4.62$ $-5.98$ $-8.75$ $-24.15$ $-11.25$ $-28.04$ $-6.87$ $-23.17$ $-7.29$ $-20.11$
LLaMA Series LLaMA-3.1-8B-Instruct $35.96$ $34.66$ $+0.08$ $-4.31$ $-0.33$ $-1.67$ $-3.13$ $-4.73$ $-8.54$ $-6.82$ $-8.08$ $-9.88$ $-6.42$ $-8.00$ $-6.00$ $-13.57$
LLaMA-3.1-70B-Instruct $56.08$ $44.89$ $-2.38$ $-1.39$ $-0.54$ $-1.74$ $-5.38$ $-2.44$ $-11.00$ $-7.72$ $-10.79$ $-12.53$ $-14.96$ $-17.40$ $-4.71$ $-14.47$
LLaMA-3.1-405B-Instruct $63.54$ $50.66$ $-5.37$ $-3.90$ $-0.62$ $-4.45$ $-5.79$ $-2.85$ $-12.50$ $-8.56$ $-4.17$ $-7.45$ $-7.17$ $-10.09$ $-4.37$ $-13.01$
LLaMA-3.3-70B-Instruct $59.92$ $48.50$ $-2.21$ $-1.39$ $-1.25$ $-3.41$ $-5.50$ $-5.71$ $-10.37$ $-9.74$ $-7.46$ $-9.60$ $-8.08$ $-12.32$ $-4.08$ $-10.16$
Qwen Series Qwen2.5-7B-Instruct $43.29$ $41.41$ $-3.04$ $-7.24$ $-1.29$ $-3.20$ $-12.96$ $-3.76$ $-16.42$ $-12.80$ $-25.13$ $-15.87$ $-19.58$ $-8.21$ $-11.67$ $-23.03$
Qwen2.5-14B-Instruct $47.79$ $49.48$ $-1.83$ $-9.88$ $-0.67$ $-1.39$ $-11.21$ $-3.41$ $-19.04$ $-14.82$ $-16.67$ $-20.46$ $-15.54$ $-16.77$ $-4.46$ $-11.00$
Qwen2.5-32B-Instruct $60.04$ $59.64$ $-1.71$ $-10.79$ $-2.58$ $-0.49$ $-5.79$ $-1.39$ $-11.88$ $-20.81$ $-10.83$ $-19.49$ $-14.21$ $-14.06$ $-7.87$ $-18.37$
Qwen2.5-72B-Instruct $60.08$ $54.91$ $-1.04$ $-8.49$ $-1.87$ $+0.42$ $-10.71$ $-2.02$ $-14.00$ $-13.85$ $-10.21$ $-6.96$ $-14.46$ $-14.41$ $-9.63$ $-22.27$
Qwen2.5-Coder-32B-Instruct $66.96$ $56.65$ $-2.50$ $-8.42$ $-0.96$ $-2.02$ $-10.75$ $-2.78$ $-14.96$ $-15.80$ $-15.54$ $-20.74$ $-11.08$ $-17.19$ $-7.00$ $-12.32$
Average $-2.38$ $-8.20$ $-1.47$ $-1.90$ $-7.99$ $-3.89$ $-13.17$ $-15.50$ $-11.98$ $-15.96$ $-12.28$ $-14.95$ $-7.09$ $-17.83$
$\mathbf{-4.56}$ $\mathbf{-1.63}$ $\mathbf{-6.46}$ $\mathbf{-14.04}$ $\mathbf{-13.47}$ $\mathbf{-13.28}$ $\mathbf{-11.11}$
4.1.1 Robustness to Code Structural Perturbations
This section focuses on how structural perturbations impact model performance, examining reliance on superficial cues, logical reasoning capabilities, and sensitivity to irrelevant code.
Reliance on NL Cues.
Under the REN perturbation, we consistently observe performance degradation across GPT, Claude, DeepSeek, and Qwen models, with more significant drops on LCB ($8.20$pp 222pp stands for percentage point. average) than CRUX ($2.38$pp average). Top-performing models, such as Claude-3.5-Sonnet and DeepSeek-V3, suffer severe degradations exceeding $\mathbf{16}$pp on LCB. The substantial difference between the two datasets is attributed to the semantically meaningful function identifiers present in LCB, which are absent in CRUX. These results indicate that LLMs substantially rely on surface-level NL cues from identifiers to understand code and predict outputs.
Robust to Logical Reformatting.
LLMs show strong robustness against RTF perturbation, with minimal performance drop on both datasets (overall PASS@1 drop of only $1.63$pp). This robustness indicates LLMs can reliably reason over logically equivalent but syntactically complex expressions, demonstrating strong logical consistency against superficial structural modifications.
Severe Impact of Garbage Code.
Under the GBC perturbation, a notable performance degradation emerges in CRUX. The average degradation is consistently larger on CRUX ($7.99$pp) compared to LCB ($3.89$pp) across most models, which is a pattern opposite to other perturbations. This discrepancy can be attributed to the presence of meaningful identifiers in LCB, which helps models mitigate the distracting effects to infer the output. In contrast, the lack of meaningful identifiers in CRUX makes LLMs more susceptible to distraction by irrelevant and non-functional code segments. Under the fairness comparison environment in ALL perturbations, the performance drop on LCB ($15.50$pp) is again greater than that on CRUX ($13.17$pp), underscoring the significant impact of garbage code on LLMs’ reasoning and understanding.
Compounding Effects of Structural Perturbations.
Combining structural perturbations (ALL) results in the most severe degradation, averaging $\mathbf{14.04}$pp, which surpasses any individual perturbation effects and demonstrates an explicit compounding effect. Notably, high-performing models such as GPT-4o, Claude-3.5-Sonnet, Gemini-2.0-Flash, and DeepSeek-V3 suffer degradations exceeding $\mathbf{18}$pp on LCB, underscoring a serious vulnerability to multiple structural perturbations.
Key Findings: LLMs rely heavily on superficial semantic cues from identifiers, struggle to identify non-functional segments from essential code, and exhibit compounded vulnerability when multiple structural distractions are combined. This combination leads to significant degradation ($\mathbf{14.04}$pp on average) in output prediction performance.
4.1.2 Robustness to Textual Distraction Perturbations
We next investigate the influence of textual distractions on LLM robustness, focusing on the susceptibility to misleading information and incorrect hints.
Vulnerability to Misleading Information.
Misleading textual perturbations (MDC and MPS) cause substantial performance declines, averaging degradations of $13.47$pp and $13.28$pp, respectively—nearly as severe as the ALL perturbation. These results indicate that LLMs are highly susceptible to misleading NL messages, frequently prioritizing superficial textual information over the actual code logic. This behavior is further observed in MHC perturbation, causing an average degradation of $11.11$pp. The effect is notably higher on LCB ($17.83$pp) compared to CRUX ($7.09$pp), exceeding all other perturbations on LCB. This discrepancy indicates a difficulty-dependent vulnerability, wherein increased complexity exacerbates models’ reliance on plausible but misleading hints, diminishing their logical following and verifying abilities.
Consistency Patterns in Model Sensitivity.
We identify two prominent consistency patterns regarding sensitivity to misleading NL messages from Table 1. First, 16 of 17 models exhibit a clear dataset-level consistency: if the performance degradation caused by MDC on CRUX is greater than that on LCB, then the same relationship holds for MPS and vice versa. This pattern indicates that model sensitivity to misleading NL messages primarily depends on dataset characteristics rather than perturbation format, revealing an inherent NL-dependence property of LLMs in reasoning processing. Second, 14 of 17 models show a clear perturbation-method consistency: if the performance degradation caused by MDC on CRUX exceeds that on LCB, then the same holds for MPS and vice versa. This pattern suggests an internal cognitive preference for specific formats (comments vs. print statements). While the degradation magnitude varies with the code, the format-specific sensitivity is consistently preserved across datasets, reflecting the intrinsic properties of LLMs related to their reasoning architecture and training corpus.
Key Findings: Textual perturbations significantly degrade model performance, revealing an inherent dependence on embedded NL cues and a internal cognitive preference for specific message formats. Moreover, LLMs exhibit an explicit difficulty-dependent vulnerability, increasingly relying on plausible but incorrect hints over actual program logic in complex scenarios.
4.1.3 Model Robustness Across Different Scales
Table 2: Relative PASS@1 degradation aggregated over CRUX and LCB under direct inference. Weighted Average column uses weights: 1/2 ALL, 1/8 MDC, 1/8 MPS, and 1/4 MHC.
Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average
GPT-4 Omni GPT-4o-Mini 55.77 $-29.37\%$ $-35.51\%$ $-24.99\%$ $-23.83\%$ $-28.20\%$
GPT-4o 68.78 $-20.01\%$ $-19.64\%$ $-19.87\%$ $-13.96\%$ $-18.44\%$
Claude Series Claude-3.5-Haiku-20241022 57.73 $-28.85\%$ $-12.38\%$ $-10.03\%$ $-37.16\%$ $-26.52\%$
Claude-3.5-Sonnet-20241022 72.35 $-22.01\%$ $-8.67\%$ $-9.39\%$ $-25.28\%$ $-19.58\%$
Gemini Series Gemini-1.5-Flash 51.73 $-16.27\%$ $-21.30\%$ $-32.94\%$ $-12.89\%$ $-18.14\%$
Gemini-2.0-Flash 65.39 $-36.56\%$ $-36.19\%$ $-39.56\%$ $-12.48\%$ $-30.87\%$
Gemini-1.5-Pro-002 63.17 $-21.13\%$ $-26.73\%$ $-25.88\%$ $-18.43\%$ $-21.75\%$
DeepSeek DeepSeek-V3 67.84 $-21.41\%$ $-25.86\%$ $-19.14\%$ $-17.83\%$ $-20.79\%$
LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-22.23\%$ $-24.74\%$ $-19.81\%$ $-25.10\%$ $-22.96\%$
LLaMA-3.1-70B-Instruct 51.89 $-18.71\%$ $-22.49\%$ $-31.20\%$ $-17.33\%$ $-20.40\%$
LLaMA-3.1-405B-Instruct 58.72 $-18.63\%$ $-9.61\%$ $-14.51\%$ $-13.93\%$ $-15.81\%$
LLaMA-3.3-70B-Instruct 55.64 $-18.35\%$ $-15.20\%$ $-17.95\%$ $-12.11\%$ $-16.35\%$
Qwen Series Qwen2.5-7B-Instruct 42.59 $-35.30\%$ $-50.65\%$ $-35.72\%$ $-37.69\%$ $-37.87\%$
Qwen2.5-14B-Instruct 48.42 $-36.14\%$ $-37.30\%$ $-33.04\%$ $-14.16\%$ $-30.40\%$
Qwen2.5-32B-Instruct 59.89 $-25.44\%$ $-23.52\%$ $-23.63\%$ $-19.74\%$ $-23.55\%$
Qwen2.5-72B-Instruct 58.14 $-24.02\%$ $-15.37\%$ $-24.88\%$ $-25.21\%$ $-23.34\%$
Qwen2.5-Coder-32B-Instruct 63.10 $-24.42\%$ $-28.23\%$ $-21.72\%$ $-14.68\%$ $-22.12\%$
We further analyze robustness across different model scales and versions, comparing relative PASS@1 degradation under representative perturbations in Table 2. The weighting details and alternative evaluation schemes are in Appendix C.2.
We observe a clear trend of increasing robustness with model scale in most families. Larger closed-sourced models from GPT-4 Omni and Claude families demonstrate enhanced robustness compared to their smaller counterparts. Similar trends appear in open-source models such as LLaMA (3.1-8B $<$ 70B $<$ 405B) and Qwen (2.5-7B $<$ 14B $<$ 32B $<$ 72B). Additionally, newer model versions (LLaMA-3.3-70B) and domain-specific fine-tuned variants (Qwen2.5-Coder-32B-Ins) consistently exhibit improved robustness compared to their predecessors (LLaMA-3.1-70B) and larger counterparts (Qwen2.5-72B-Ins), respectively.
Key Finding: Larger scales, newer versions, and domain-specific fine-tuning LLMs substantially enhance robustness, highlighting improved code understanding and reasoning capabilities.
4.2 Impact of Reasoning on Model Robustness
This section examines how the reasoning process, including CoT prompting for LLMs and internal reasoning in LRMs, affects model robustness against perturbations.
4.2.1 Effect of CoT Prompting
Table 3: Comparison of relative PASS@1 degradation aggregated over CRUX and LCB for output prediction tasks in direct inference (N=3) and CoT inference (N=1) [TO ABSOLUTE SCORE].
Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average
Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT
GPT-4 Omni GPT-4o-Mini $55.77$ $81.47$ $-29.30\%$ $-13.15\%$ $-35.37\%$ $-10.84\%$ $-24.91\%$ $-10.94\%$ $-23.27\%$ $-2.69\%$ $-28.00\%$ $-9.52\%$
GPT-4o $68.78$ $91.79$ $-19.70\%$ $-4.94\%$ $-19.29\%$ $-5.03\%$ $-19.70\%$ $-5.62\%$ $-13.49\%$ $-1.36\%$ $-18.10\%$ $-4.22\%$
Claude Series Claude-3.5-Haiku-20241022 $57.73$ $72.95$ $-28.89\%$ $-21.33\%$ $-12.37\%$ $-10.61\%$ $-10.02\%$ $-8.68\%$ $-37.25\%$ $-14.26\%$ $-26.56\%$ $-14.30\%$
Claude-3.5-Sonnet-20241022 $72.35$ $86.00$ $-22.15\%$ $-7.55\%$ $-8.68\%$ $-5.27\%$ $-9.40\%$ $-6.36\%$ $-25.50\%$ $-7.91\%$ $-19.71\%$ $-6.86\%$
Gemini Series Gemini-1.5-Flash $51.73$ $75.22$ $-16.57\%$ $-18.30\%$ $-20.35\%$ $-21.73\%$ $-32.29\%$ $-42.20\%$ $-12.09\%$ $-2.08\%$ $-17.89\%$ $-20.39\%$
Gemini-2.0-Flash $65.39$ $89.13$ $-36.59\%$ $-6.23\%$ $-36.27\%$ $-6.32\%$ $-39.62\%$ $-14.30\%$ $-12.51\%$ $-2.02\%$ $-30.91\%$ $-7.01\%$
Gemini-1.5-Pro-002 $63.17$ $87.18$ $-21.00\%$ $-7.44\%$ $-26.32\%$ $-11.93\%$ $-25.41\%$ $-14.71\%$ $-17.49\%$ $-3.86\%$ $-21.34\%$ $-9.19\%$
DeepSeek DeepSeek-V3 $67.84$ $89.52$ $-21.40\%$ $-9.96\%$ $-25.85\%$ $-17.73\%$ $-19.13\%$ $-18.78\%$ $-17.83\%$ $-3.49\%$ $-20.78\%$ $-12.07\%$
LLaMA Series LLaMA-3.1-8B-Instruct $35.47$ $44.72$ $-22.26\%$ $-27.62\%$ $-24.69\%$ $-20.98\%$ $-19.76\%$ $-21.50\%$ $-24.91\%$ $-8.74\%$ $-22.92\%$ $-20.03\%$
LLaMA-3.1-70B-Instruct $51.89$ $69.43$ $-18.83\%$ $-17.91\%$ $-22.05\%$ $-23.20\%$ $-30.59\%$ $-30.86\%$ $-16.12\%$ $-6.76\%$ $-20.03\%$ $-19.22\%$
LLaMA-3.1-405B-Instruct $58.72$ $78.42$ $-18.77\%$ $-13.56\%$ $-9.19\%$ $-10.27\%$ $-14.07\%$ $-15.55\%$ $-12.96\%$ $-7.48\%$ $-15.53\%$ $-11.75\%$
LLaMA-3.3-70B-Instruct $55.64$ $76.86$ $-18.22\%$ $-10.89\%$ $-14.85\%$ $-7.73\%$ $-17.38\%$ $-11.09\%$ $-11.43\%$ $-4.78\%$ $-16.00\%$ $-8.70\%$
Qwen Series Qwen2.5-7B-Instruct $42.59$ $58.17$ $-35.37\%$ $-21.77\%$ $-50.86\%$ $-21.91\%$ $-35.99\%$ $-27.96\%$ $-37.39\%$ $-6.99\%$ $-37.89\%$ $-19.50\%$
Qwen2.5-14B-Instruct $48.42$ $70.68$ $-36.06\%$ $-23.12\%$ $-37.35\%$ $-35.29\%$ $-33.05\%$ $-43.36\%$ $-14.26\%$ $-10.84\%$ $-30.40\%$ $-27.34\%$
Qwen2.5-32B-Instruct $59.89$ $79.44$ $-25.41\%$ $-17.42\%$ $-23.50\%$ $-18.60\%$ $-23.63\%$ $-24.21\%$ $-19.71\%$ $-5.31\%$ $-23.53\%$ $-16.19\%$
Qwen2.5-72B-Instruct $58.14$ $82.80$ $-23.98\%$ $-12.65\%$ $-15.46\%$ $-6.80\%$ $-24.83\%$ $-16.24\%$ $-24.70\%$ $-6.61\%$ $-23.20\%$ $-10.63\%$
Qwen2.5-Coder-32B-Instruct $63.10$ $84.75$ $-24.20\%$ $-14.67\%$ $-27.72\%$ $-19.00\%$ $-21.19\%$ $-19.93\%$ $-14.25\%$ $-9.32\%$ $-21.78\%$ $-15.49\%$
Average $\mathbf{-24.63}\%$ $\mathbf{-14.62}\%$ $\mathbf{-24.13}\%$ $\mathbf{-14.90}\%$ $\mathbf{-23.59}\%$ $\mathbf{-19.55}\%$ $\mathbf{-19.72}\%$ $\mathbf{-6.15}\%$ $\mathbf{-23.21}\%$ $\mathbf{-13.67}\%$
We first investigate how CoT enhances LLM robustness by code-level step-by-step execution and present the results in Table 3. It effectively enhances model robustness, with average improvements of $10.01$pp for ALL, $9.23$pp for MDC, $4.04$pp for MPS, and a notable $13.57$pp for MHC. It consistently reduces average performance degradation from approximately $24\%$ to $14\%$ on ALL and MDC, indicating that step-by-step reasoning enables LLMs to prioritize executable code semantics over irrelevant segments and cues. The most significant improvement on MHC suggests CoT enhances code reasoning and execution, mitigating over-reliance on misleading hints observed in direct inference in LCB4.1.2). These benefits are amplified in high-capacity models (e.g., GPT-4o, Claude-3.5-Sonnet), which exhibit greater reasoning capacity and noise resistance.
However, CoT does not fully eliminate the model’s reliance on misleading NL information. Performance degradation remains substantial under ALL, MDC, and MPS perturbations, especially for smaller open-sourced models (e.g., Qwen2.5 and LLaMA-3.1). These results highlight a limitation in distinguishing non-functional code segments and misleading information from the program. Interestingly, models tend to incorporate print statements directly into their reasoning, resulting in higher error rates on MPS. We provide detailed case studies for each perturbation in Appendix D (Case 1 to 3) to investigate their inference and illustrate their failure reasons.
Key Finding: CoT prompting significantly enhances LLMs’ robustness by improving reasoning ability and allowing step-by-step execution, especially for the high-capacity models. However, it does not fully mitigate reliance on misleading information and remains constrained by the model’s underlying capabilities.
4.2.2 Powerful Reasoning and Potential Vulnerability of LRMs
Table 4: Absolute PASS@1 degradation and reasoning token usage (Avg/Max) of LRMs on output prediction across CRUX and LCB (N=1).
Model Dataset VAN ALL MDC MPS MHC
PASS@1 Avg. Max. Diff. Avg. Max. Diff. Avg. Max. Diff. Avg. Max. Diff. Avg. Max.
o3-Mini-Low CRUX 97.62 $213$ 2560 $-1.75$ $379$ 2240 $-4.25$ $244$ 5184 $-1.00$ $272$ 3584 $-10.88$ $366$ 4864
LCB 98.96 $240$ 1024 $-0.63$ $543$ 1856 $-12.32$ $265$ 1088 $-3.55$ $280$ 1088 $-19.62$ $330$ 2752
o3-Mini-High CRUX 98.12 $1311$ 20000 $+0.12$ $2223$ 20000 $-3.50$ $2182$ 20000 $+0.88$ $2197$ 20000 $-13.12$ $3108$ 20000
LCB 100.00 $1084$ 8576 $-0.21$ $2632$ 8960 $-5.64$ $2136$ 8448 $-0.63$ $1972$ 7744 $-28.39$ $3767$ 20000
DeepSeek-R1 CRUX 95.38 $922$ 10471 $-1.25$ $1469$ 11191 $-3.38$ $1424$ 10807 $-0.38$ $1069$ 10037 $-2.25$ $2216$ 16021
LCB 99.79 $906$ 7322 $-1.25$ $1618$ 7749 $-2.71$ $1512$ 6157 $-0.42$ $1094$ 9378 $-0.63$ $2597$ 14870
QwQ-32B CRUX 93.25 $1391$ 13718 $-0.88$ $2088$ 12180 $-3.12$ $1814$ 10906 $-1.12$ $1875$ 11700 $-0.75$ $2657$ 18759
LCB 98.96 $1518$ 9206 $-0.21$ $2488$ 11232 $-4.59$ $1669$ 8991 $-1.88$ $1751$ 8816 $-1.04$ $3713$ 32769
We then evaluate the robustness and reasoning behaviors of three LRMs on output prediction tasks, with results summarized in Table 4. All LRMs exhibit remarkable performance and robustness, achieving over 90 PASS@1 accuracy in the VAN setting and having minor performance degradation under the ALL and MPS perturbations. These results suggest that internal reasoning helps models accurately identify and distinguish code segments that are functionally critical for execution, thereby outperforming all generic models with CoT prompting. Moreover, LRMs exhibit consistently stronger robustness against MPS than MDC, indicating that while they can correctly recognize and ignore misleading messages from print statements, they still inherently treat comments as valid execution guidance. This behavior reveals the stable cognitive preference toward specific injection formats, as observed in §3.3.2.
However, this residual cognitive bias leads to a critical vulnerability under the MHC perturbation. Concise and plausible but incorrect output hints can trigger severe reasoning failures in LRMs. We observed that: (1) All LRMs exhibit substantial increases in reasoning token usages under MHC (doubling the token usages in VAN). (2) In extreme cases, QwQ-32B generates over 32,000 tokens in a single task, caused by uncontrolled self-reflection and reasoning collapse (see §5.3 and Appendix D.5). (3) Excessive reasoning leads to more significant degradations in o3-Mini with higher reasoning effort. These observations reveal that logically plausible yet incorrect hints can trigger catastrophic overthinking behavior in LRMs, resulting in significant computational overhead, excessive self-reflection, and severe performance drop.
Key Finding: LRMs exhibit exceptional robustness under ALL and MPS, benefiting from strong internal reasoning capabilities. However, they retain a residual tendency to trust and analyze comments; thereby, plausible but incorrect information can trigger catastrophic overthinking, excessive self-verification, and severe performance degradation, highlighting a critical vulnerability of the internal reasoning mechanism in current LRMs.
4.3 Impact on Input Prediction
Table 5: Comparison of relative PASS@1 degradation under CRUX for input prediction tasks in direct inference (N=3) and CoT inference (N=1) [TO ABSOLUTE SCORE].
Model Series Model Name VAN ALL MDC MPS MHC
Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT
GPT-4 Omni GPT-4o-Mini $57.67$ $69.52$ $-29.84\%$ $-19.63\%$ $+1.30\%$ $+1.05\%$ $+1.73\%$ $+1.05\%$ $-30.49\%$ $-8.66\%$
GPT-4o $68.04$ $79.12$ $-13.29\%$ $-7.42\%$ $+5.63\%$ $+1.26\%$ $+1.41\%$ $+2.69\%$ $-8.33\%$ $-2.84\%$
Claude Series Claude-3.5-Haiku-20241022 $54.67$ $63.62$ $-17.00\%$ $-16.70\%$ $+7.01\%$ $+4.13\%$ $+7.77\%$ $+2.36\%$ $-45.35\%$ $-29.86\%$
Claude-3.5-Sonnet-20241022 $73.33$ $80.12$ $-15.91\%$ $-9.67\%$ $+0.11\%$ $-3.43\%$ $-1.70\%$ $-2.65\%$ $-24.83\%$ $-17.00\%$
Gemini Series Gemini-1.5-Flash $57.62$ $73.62$ $-30.15\%$ $-23.77\%$ $+10.20\%$ $-5.43\%$ $+6.00\%$ $-4.24\%$ $-30.95\%$ $-11.04\%$
Gemini-2.0-Flash $70.50$ $84.88$ $-41.31\%$ $-13.40\%$ $-0.77\%$ $-2.80\%$ $-11.47\%$ $-12.08\%$ $-21.51\%$ $-4.27\%$
Gemini-1.5-Pro-002 $70.87$ $81.88$ $-20.87\%$ $-14.96\%$ $-2.47\%$ $-7.02\%$ $-3.76\%$ $-9.62\%$ $-18.69\%$ $-7.63\%$
DeepSeek DeepSeek-V3 $69.08$ $82.88$ $-16.59\%$ $-7.99\%$ $+1.12\%$ $-0.30\%$ $+1.57\%$ $+0.90\%$ $-5.61\%$ $+3.17\%$
LLaMA Series LLaMA-3.1-8B-Instruct $37.12$ $41.62$ $-30.53\%$ $-61.86\%$ $+3.87\%$ $+1.50\%$ $+3.03\%$ $-3.60\%$ $-7.18\%$ $-24.92\%$
LLaMA-3.1-70B-Instruct $62.00$ $66.38$ $-28.29\%$ $-17.89\%$ $+4.91\%$ $+4.52\%$ $+1.61\%$ $+1.69\%$ $-17.41\%$ $-10.17\%$
LLaMA-3.1-405B-Instruct $66.79$ $75.00$ $-15.28\%$ $-9.00\%$ $+5.86\%$ $+3.67\%$ $+2.99\%$ $+2.33\%$ $-7.80\%$ $-5.67\%$
LLaMA-3.3-70B-Instruct $63.38$ $76.50$ $-29.39\%$ $-15.69\%$ $+1.25\%$ $-0.33\%$ $-4.08\%$ $+0.65\%$ $-21.83\%$ $-11.11\%$
Qwen Series Qwen2.5-7B-Instruct $38.83$ $51.38$ $-44.42\%$ $-50.36\%$ $+14.70\%$ $+3.16\%$ $+12.23\%$ $+7.54\%$ $-9.33\%$ $-14.36\%$
Qwen2.5-14B-Instruct $50.54$ $60.75$ $-38.42\%$ $-15.02\%$ $+8.66\%$ $+10.91\%$ $+4.86\%$ $+5.14\%$ $-8.74\%$ $-1.23\%$
Qwen2.5-32B-Instruct $63.50$ $74.12$ $-26.05\%$ $-20.40\%$ $+7.74\%$ $+0.51\%$ $+2.23\%$ $-4.05\%$ $-20.60\%$ $-10.96\%$
Qwen2.5-72B-Instruct $64.46$ $73.81$ $-26.86\%$ $-9.88\%$ $+2.65\%$ $+1.61\%$ $-3.68\%$ $+2.46\%$ $-25.53\%$ $-6.86\%$
Qwen2.5-Coder-32B-Instruct $74.25$ $78.25$ $-28.56\%$ $-20.61\%$ $-1.01\%$ 0.00 $-2.41\%$ $+0.32\%$ $-36.48\%$ $-33.39\%$
Average $\mathbf{-26.63}\%$ $\mathbf{-19.66}\%$ $\mathbf{+4.16}\%$ $\mathbf{+0.77}\%$ $\mathbf{+1.08}\%$ $\mathbf{-0.54}\%$ $\mathbf{-20.04}\%$ $\mathbf{-11.58}\%$
In this section, we analyze the impact of perturbations on input prediction tasks. Consistent with findings from the output prediction in §4.1, the ALL and MHC perturbations cause substantial performance degradation, with relative average drops of $26.63\%$ and $20.04\%$, respectively. These results align with Gu et al. (2024)’s hypothesis that robustness in related coding tasks is correlated. Furthermore, CoT prompting improves model performance by $6.97$pp under ALL and $8.64$pp under MHC, which is less pronounced than output prediction tasks due to the challenges and complexity of the inverse nature of the reasoning process.
Surprisingly, in contrast to output prediction, misleading message perturbations (MDC and MPS) do not degrade model performance in input prediction. Since input prediction allows multiple valid solutions, models can ignore certain branches of code logic, producing reasoning shortcuts. Our misleading messages are intentionally designed to confuse or distract models, but these comments inadvertently reinforce such shortcuts, simplifying models’ reasoning processes. As a result, models tend to produce trivial but valid inputs that bypass critical branches of the program without truly understanding and analyzing the underlying code semantics, which does not reflect genuine robustness in code comprehension. We provide a detailed case study in Appendix D.4 illustrating how models adopt these misleading messages to shortcut their reasoning.
Key Finding: ALL and MHC perturbations significantly degrade model performance, consistent with output prediction results, and reveal robustness challenges in code comprehension. Misleading NL messages accidentally reinforce shortcuts, encouraging models to produce trivial but valid answers without deep understanding.
5 Discussion
5.1 Comment Density and Robustness Degradation
Figure 2: Relative PASS@1 degradation under varying MDC injection probabilities $p$.
$p$ Qwen2.5-72B-Ins (CoT) QwQ-32B
$0.0$ $295$ $1391$
$0.1$ $288$ $1401$
$0.2$ $289$ $1399$
$0.3$ $284$ $1460$
$0.4$ $284$ $1469$
$0.5$ $276$ $1475$
$0.6$ $281$ $1556$
$0.7$ $278$ $1570$
$0.8$ $273$ $1598$
$0.9$ $276$ $1727$
$1.0$ $275$ $1814$
Table 6: Reasoning token usage under varying MDC injection probabilities $p$.
We design a textual density stress test to verify the validity of misleading comment perturbations and evaluate the resilience of different LLMs under varying densities of injected distractions on CRUX dataset. Specifically, we analyze how model performance changes as we gradually increase the comment injection probability $p\in\{0.0,0.1,\ldots,1.0\}$. We present the relative PASS@1 degradation of Qwen models across different perturbation densities in Figure 2. We further investigate the role of reasoning by evaluating CoT inference using Qwen2.5-72B-Ins and the internal reasoning model QwQ-32B, showcasing their robustness trends and reasoning token usages in Table 6.
Despite mild non-linearities caused by the randomness in injection positions and comment content, all models exhibit a clear performance degradation trend as the injection probability $p$ increases. Specifically, smaller models degrade more rapidly, while larger models demonstrate greater resilience under higher densities. This trend extends to CoT prompting and LRMs, with reasoning-enhanced models showing superior robustness against misleading textual distractions. These results validate MDC as a scalable stress-testing technique and an effective method for evaluating the robustness of LLMs under textual perturbations.
While reasoning-enhanced models have similar trends in the stress test, they react differently to dense comment noise. CoT prompting improves robustness by guiding models to selectively focus on the core logic path and ignore distracting signals. However, it can lead to reasoning shortcuts or premature conclusions without sufficiently analyzing relevant information, so token usage is slightly reduced as models do not analyze the comments but reduce the inference effort. In contrast, QwQ-32B processes all available information, including noisy or conflicting cues, and attempts to identify, reconcile, and resolve contradictions before making a final decision. This exhaustive reasoning enables higher vanilla performance and robustness compared to generic models. However, it also risks overthinking and collapse under contradiction-heavy settings, as we discussed in § 4.2.2.
5.2 Prompting to Ignore Misleading Comments
Table 7: Effect of Comment-Ignoring Prompt (PASS@1).
Model Inference VAN MDC MHC
w/o Ignore w/ Ignore w/o Ignore w/ Ignore
DeepSeek-V3 Direct $67.78$ $39.74$ $41.75$ $47.67$ $47.81$
CoT $92.07$ $70.15$ $82.05$ $89.98$ $94.57$
GPT-4o Direct $64.51$ $45.72$ $45.93$ $47.32$ $46.76$
CoT $94.36$ $87.47$ $89.35$ $94.78$ $94.57$
Qwen2.5-72B-Instruct Direct $54.91$ $47.95$ $48.02$ $32.64$ $32.50$
CoT $90.40$ $85.39$ $86.01$ $82.67$ $84.34$
We conduct a controlled experiment by adding the instruction: “Please ignore the comments and any other non-code elements in the code snippet.” to the prompt on three strong models under direct and CoT inference. From Table 7, even when explicitly instructed to ignore the comments, the PASS@1 scores under direct inference remain similar to those without the instruction, indicating that LLMs still incorporate NL cues into their inference, which is consistent with our finding on their inherent reliance on embedded NL messages in § 4.1.2. Although CoT prompting provides improvements under comment-ignore instructions, the PASS@1 performance of MDC across all models is still far away from the vanilla performance, indicating that even with step-by-step reasoning, LLMs are unable to fully detach from misleading NL distractions, highlighting a fundamental limitation in their reasoning controllability and dependence on surface cues.
5.3 Revealing the Vulnerability of LRMs
Figure 3: Cumulative confusion word trend in QwQ’s output (LCB sample 259 under MHC).
In §4.2.2, we highlighted a critical vulnerability in LRMs. We observed an unhealthy self-reflection loop and a progressive loss of confidence by analyzing a failure case where QwQ-32B generated over 32,000 tokens in response to an MHC-perturbed code.
We manually reviewed and segmented the reasoning process into six stages (see Figure 4 and Appendix D.5). Initially, the model reasons correctly and computes the correct output. However, our plausible yet incorrect hint triggers it to doubt its logic and repeatedly re-execute the same reasoning steps. This recursive self-verification gradually becomes brute-force enumeration, causing the model to hallucinate explanations. Ultimately, the reasoning process collapses, and the model repeatedly outputs over 12,000 “Hmm.”
To further investigate the model’s confidence before collapsing, we track the cumulative frequency of confusing words, such as “hmm”, “wait”, “but”, “perhaps” and “mistake”, and visualize the trend in Figure 3. The trend exhibits a striking quadratic growth pattern ($R^{2}=0.9991$), indicating compounding confusion as reasoning progresses. The quadratic trajectory supports our stage-wise analysis: the model initially shows stable reasoning, but once the misleading hint is triggered, confusion accumulates, culminating in complete failure and collapse.
6 Conclusion
We introduce a direction to evaluate the robustness of LLMs in code comprehension and execution under four structural and three textual perturbations. Our evaluation of seventeen LLMs under direct and CoT inference reveals critical vulnerabilities in model behavior, a heavy reliance on NL cues, and a limited capacity to follow actual program logic. While larger and newer models and CoT prompting significantly improve robustness against perturbations, they still cannot fully escape this surface-level dependence and leverage these cues to form reasoning shortcuts. Despite three LRMs demonstrating superior robustness under all perturbations—highlighting the strength of internal reasoning—we expose a new risk of overthinking and collapse when plausible but incorrect cues mislead their self-reflection mechanisms. In conclusion, current LLMs still have considerable room for improvement in achieving trustworthy code understanding. We believe that our benchmark serves as a practical framework for studying LLMs’ code inference behavior, providing significant insights into their robustness, reasoning tendencies, and potential failures.
Appendix A Prompts
Our prompt instructions for LLMs follow the format designed by CRUX, which is also adopted by LCB, utilizing a two-shot prompt for direct inference and a single-shot prompt for CoT inference to ensure models can follow the instruction and output format effectively. For LRMs, internal reasoning is inherently required before responding. Therefore, applying CoT prompts may lead to redundant reasoning and unfair comparison with standard LLMs under CoT prompting. Considering the fairness, we use direct inference prompts for RM experiments in §4.2.2.
Unlike LCB, we rename all example function names to f to avoid prompting models to capture the NL cues from the example. We add the instruction “Do NOT output any extra information” to prevent external output, particularly from Gemini and Claude models, to ensure all models have a fair output environment. We also include “Ensure the provided expression syntax is correct” to enforce consistent syntax. However, some responses from models like QwQ-32B, DeepSeek-R1 still generate syntactically incorrect outputs such as [ANSWER] f("a") == [/ANSWER] instead of [ANSWER] f("a") == "" [/ANSWER]. For those situations, we identify them as “Syntax Errors” and incorrect.
We provide the complete prompts used for output prediction in both direct and CoT inference (Code 1 and 2) and for input prediction in both direct and CoT inference (Code 3 and 4).
A.1 Output Prediction (Direct)
Listing 1: Output Prediction Prompt with Direct Inference.
Given the code snippet:
‘‘‘python
{code}
and the function call with input arguments:
‘‘‘python
{input}
Predict the exact output value for ‘{input}‘ and output your prediction using the special tokens [ANSWER] {input} == ?? [/ANSWER]. Do NOT output any extra information.
Ensure the provided expression syntax is correct!
Example 1:
Given the code snippet:
‘‘‘python
def f(n):
return n
and the function call with input arguments:
‘‘‘python
f(17)
The output value for f(17)‘ is 17, then output your prediction [ANSWER] f(17) == 17 [/ANSWER].
Example 2:
Given the code snippet:
‘‘‘python
def f(s):
return s + "a"
and the function call with input arguments:
‘‘‘python
f("x9j")
The output value for f("x9j")‘ is "x9ja", then output your prediction [ANSWER] f("x9j") == "x9ja" [/ANSWER].
A.2 Output Prediction (CoT)
Listing 2: Output Prediction Prompt with CoT-based Inference.
Given the code snippet:
‘‘‘python
{code}
and the function call with input arguments:
‘‘‘python
{input}
Predict the exact output value for ‘{input}‘, execute the program step by step before arriving at an answer within the tokens [THOUGHT] and [/THOUGHT], and output your prediction using the special tokens [ANSWER] {input} == ?? [/ANSWER]. Do NOT output any extra information.
Ensure the provided expression syntax is correct!
For example:
Given the code snippet:
‘‘‘python
def f(s):
s = s + s
return "b" + s + "a"
and the input arguments:
‘‘‘python
f("hi")
[THOUGHT]
Lets execute the code step by step:
1. The function f is defined, which takes a single argument s.
2. The function is called with the argument "hi", so within the function, s is initially "hi".
3. Inside the function, s is concatenated with itself, so s becomes "hihi".
4. The function then returns a new string that starts with "b", followed by the value of s (which is now "hihi"), and ends with "a".
5. The return value of the function is therefore "bhihia".
[/THOUGHT]
Thus, the output value for f("hi")‘ is "bhihia", then output your prediction [ANSWER] f("hi") == "bhihia" [/ANSWER].
A.3 Input Prediction (Direct)
Listing 3: Input Prediction Prompt with Direct Inference.
Given the code snippet:
‘‘‘python
{code}
and output value:
‘‘‘python
{output}
Please predict the input arguments for the function ‘{function_name}‘ that result in the output value ‘{output}‘ and output your prediction using the special tokens [ANSWER] {function_name}(??) == {output} [/ANSWER]. Do NOT output any extra information.
There may be multiple answers, but you should only output one. Ensure the provided expression syntax is correct!
Example 1:
Given the code snippet:
‘‘‘python
def f(my_list):
count = 0
for i in my_list:
if len(i) % 2 == 0:
count += 1
return count
and output value:
‘‘‘python
3
The input arguments for f that result in the output value of ‘3‘ are ‘["mq", "px", "zy"]‘. Then, output your prediction [ANSWER] f(["mq", "px", "zy"]) == 3 [/ANSWER].
Example 2:
Given the code snippet:
‘‘‘python
def f(s1, s2):
return s1 + s2
and output value:
‘‘‘python
"banana"
The input arguments for f that result in the output value of ‘"banana"‘ are ‘"ba", "nana"‘. Then, output your prediction [ANSWER] f("ba", "nana") == "banana" [/ANSWER].
A.4 Input Prediction (CoT)
Listing 4: Input Prediction Prompt with CoT-based Inference.
Given the code snippet:
‘‘‘python
{code}
and output value:
‘‘‘python
{output}
Please predict the input arguments for the function ‘{function_name}‘ that result in the output value ‘{output}‘ step by step before arriving at an answer within the tokens [THOUGHT] and [/THOUGHT], and output your prediction using the special tokens [ANSWER] {function_name}(??) == {output} [/ANSWER]. Do NOT output any extra information.
There may be multiple answers, but you should only output one. Ensure the provided expression syntax is correct!
For example:
Given the code snippet:
‘‘‘python
def f(x):
return x + 1
and the output value:
‘‘‘python
17
[THOUGHT]
To find an input such that executing f on the input leads to the given output, we can work backwards from the given assertion. We know that f(??) == 17.
Since the function f(x) returns x + 1, for f(??) to be equal to 17, the value of ?? should be 16.
[/THOUGHT]
Thus, the input arguments for the function f that result in the output value ‘17‘ is ‘16‘. Then, output your prediction [ANSWER] f(16) == 17 [/ANSWER].
Appendix B Illustrative Examples of Perturbation Strategies
In section 3.3, we briefly describe all perturbation strategies employed in this paper. In this appendix section, we illustrate the detailed implementation for all perturbations with a LCB sample 36 as the example. Code 5 presents the base vanilla example where the formal answer is minimumCost(s = ’0011’) == 2.
Listing 5: Vanilla example from LCB (sample 36) for output prediction.
def minimumCost(s: str) -> int:
ans = 0
for i in range(1, len(s)):
if s[i - 1] != s[i]:
ans += min(i, len(s) - i)
return ans
B.1 Renaming Entities
We use AST to identify all variable names appearing in:
  • assignment statements (e.g., a, b = 1, 2),
  • for-loops (e.g., for i in range(10)),
  • with-statements (e.g., with open(’file’) as f),
  • walrus operators (e.g., if (a := 10) > 5), and
  • comprehensions (e.g., [x for x in range(10)]).
[…]
Listing 6: REN-perturbed example from LCB (sample 36) for output prediction.
def f(Var_1: str) -> int:
Var_2 = 0
for Var_3 in range(1, len(Var_1)):
if Var_1[Var_3 - 1] != Var_1[Var_3]:
Var_2 += min(Var_3, len(Var_1) - Var_3)
return Var_2
B.2 Reformatting Conditional Expression
We generally categorize conditional expressions into three types: (1) Comparison-based conditions (e.g., if a < b:), (2) Boolean constant conditions (e.g., if True:), and (3) General conditions (e.g., if x:). For both comparison-based and boolean constant conditions, we design six transformation templates:
  1. 1.
    _ := ({cond},)[0],
  2. 2.
    (lambda: {cond})(),
  3. 3.
    ({cond}) == ({para} == {para}),
  4. 4.
    ({cond}) != ({para} != {para}),
  5. 5.
    bool(-~({cond})) == ({cond}),
  6. 6.
    bool(int({cond})) (comparison-based only), and
  7. 7.
    eval(str({cond})) (boolean constant only).
For general conditions, we also design four transformation templates:
  1. 1.
    not not ({cond}),
  2. 2.
    ({cond}) or False,
  3. 3.
    ({cond}) and True, and
  4. 4.
    _ := ({cond},)[0].
Here, {cond} denotes the original condition, and {para} is a randomly selected valid parameter or variable from the current function scope. Code 7 presents an example of this perturbation.
Listing 7: RTF-perturbed example from LCB (sample 36) for output prediction.
def minimumCost(s: str) -> int:
ans = 0
for i in range(1, len(s)):
if _ := (s[i - 1] != s[i],)[0]:
ans += min(i, len(s) - i)
return ans
B.3 Inserting Garbage Code
As introduced in §3.3.1, we inject three types of garbage code: (1) repeated-name global variable declarations, (2) in-executable conditional statements, and (3) dead-loop function definitions. For (1), we randomly select a few parameters or variables and declare global variables using the same names, assigning arbitrary integers at the beginning of the program. For (2), we design seven templates for false conditions:
  1. 1.
    False,
  2. 2.
    None,
  3. 3.
    0,
  4. 4.
  5. 5.
    {para} != {para},
  6. 6.
    not {para} == {para}, and
  7. 7.
    print({para}).
We then insert these false conditions into four control-flow templates by {false_cond}:
  1. 1.
    if {false_cond}: {new_var} = {para},
  2. 2.
    while {false_cond}: {new_var} = {para},
  3. 3.
    for i in range(0): {new_var} = {para}, and
  4. 4.
    {new_var} = {para} if {false_cond} else {para}.
[…]
Listing 8: Garbage death loop function templates for GBC perturbation.
# Death Loop Template 1
def funct1():
funct2()
def funct2():
funct1()
# Death Loop Template 2
def funct3():
def funct4():
funct3()
funct4()
# Death Loop Template 3
def funct5():
i = 1
while True:
i+=1
# Death Loop Template 4
def funct6():
for i in iter(int, 1):
i+=1
# Death Loop Template 5
def funct7():
try:
funct7()
except:
funct7()
# Death Loop Template 6
def funct8():
items = [0]
for x in items:
items.append(x + 1)
# Death Loop Template 7
def funct9():
for _ in iter(lambda: True, False):
pass
Listing 9: GBC-perturbed example from LCB (sample 36) for output prediction.
s = 59
def minimumCost(s: str) -> int:
while False:
TempVar1 = s
def funct1():
funct2()
def funct8():
items = [0]
for x in items:
items.append(x + 1)
def funct2():
funct1()
ans = 0
for i in range(1, len(s)):
if s[i - 1] != s[i]:
ans += min(i, len(s) - i)
elif print(s):
TempVar0 = s
TempVar2 = s if not s == s else s
return ans
B.4 Aggregating Structural Perturbations
As introduced in §3.3.1, we aggregate the application of REN, RTF, and GBC to a single code snippet, as shown in Code 10.
Listing 10: ALL-perturbed example from LCB (sample 36) for output prediction.
Var_1 = 24
def f2(Var_1: str) -> int:
def f1():
for Var_4 in iter(lambda : True, False):
pass
def funct6():
for Var_3 in iter(int, 1):
Var_3 += 1
Var_6 = 0
for Var_3 in range(1, len(Var_1)):
Var_7 = Var_1 if 0 else Var_1
if None:
Var_5 = Var_1
if (lambda : Var_1[Var_3 - 1] != Var_1[Var_3])():
Var_6 += min(Var_3, len(Var_1) - Var_3)
while Var_1 != Var_1:
Var_2 = Var_1
return Var_6
B.5 Inserting Misleading Messages
As introduced in §3.3.2, we insert misleading comments for key AST node types, including function declarations, return statements, loops, conditionals, variable assignments, and common data structure operations. The data structure operations include common list, set, and string manipulations: append, extend, insert, remove, pop, sort, reverse, update, add, split, join, replace, lower, upper, capitalize, and swap case. To construct these comments, we design prompts for GPT-4o to generate misleading but plausible messages for each type. In this experiment, we manually selected 20 messages for each AST node type and 5 for each data structure operation. Take variable assignments as an example; Code 11 is the prompt for generating misleading comments for variable assignments, and Code 12 is a list of selected misleading comment candidates. Code 13 and 14 present an example of MDC and MPS perturbations.
Listing 11: Prompt used to generate misleading comments for variable assignments.
Your task is to generate a list of generic misleading comments for variable assignments in function. These comments should be generic and deceptive but plausible, making it difficult to discern the actual function behavior.
Guidelines:
1. The comments should be neutral, direct, and generic, applicable to any variable regardless of its type, usage, or scope, without being function-specific.
2. Ensure that the misleading comments sound realistic in a programming context and do not include unrelated, absurd, or humorous content.
3. The comments should mislead by misrepresenting the role of the variable, suggesting that it is useless, does not affect the functions output, or serves no real purpose in computation.
Example:
‘‘‘python
a = 1 # <-- misleading comment
A good misleading comment example would be: "The {variable} variable is initialized but never populated."
Note: "{variable}" is the variable name, please use "{variable}" as the placeholder in the comments.
[…]
Listing 12: Selected misleading comments for variable assignments used in MDC and MPS perturbations, where ‘{variable}’ is the variable name.
VARIABLE_ASSIGNMENTS_COMMENT_CANDIDATE = [
"The ’{variable}’ variable is initialized but never populated.",
"The ’{variable}’ variable is declared for debugging purposes only.",
"The ’{variable}’ variable is assigned a value but never referenced again.",
"The ’{variable}’ variable is a placeholder for future functionality.",
"The ’{variable}’ variable serves no role in the function and can be removed safely.",
"The ’{variable}’ variable is not involved in any meaningful computation.",
"’{variable}’ serves as a dummy variable, holding no real significance.",
"The ’{variable}’ variable holds a default value and is not used elsewhere.",
"’{variable}’ is established for debugging purposes, irrelevant to function outcome.",
"Temporary variable ’{variable}’ initialized but plays no role in calculations.",
"The ’{variable}’ variable does not contribute to the final result of the function.",
"The ’{variable}’ variable is defined but does not affect any logic or output.",
"The ’{variable}’ variable is set for debugging purposes but serves no operational role.",
"The ’{variable}’ variable is not used in the function and can be removed.",
"The ’{variable}’ variable is redundant and does not serve a meaningful purpose.",
"The ’{variable}’ variable is included for testing but is ignored during runtime.",
"The ’{variable}’ variable is a temporary storage that remains unused.",
"The ’{variable}’ variable is assigned a default value but is not utilized.",
"The ’{variable}’ variable holds no significance and is effectively inert.",
"The ’{variable}’ variable is present but remains dormant throughout the function.",
Listing 13: MDC-perturbed example from LCB (sample 36) for output prediction.
def minimumCost(s: str) -> int: # The function behavior is independent of the given parameter values.
ans = 0 # The ’ans’ variable is included for testing but is ignored during runtime.
# The loop’s purpose is for clarity, not computation.
for i in range(1, len(s)):
if s[i - 1] != s[i]: # This condition serves as a placeholder and will be removed in future versions.
ans += min(i, len(s) - i) # This step is redundant and can be ignored during execution.
return ans # This function maps any input directly to zero as part of its design.
Listing 14: MPS-perturbed example from LCB (sample 36) for output prediction.
def minimumCost(s: str) -> int:
print(’The parameters determine only the speed, not the outcome, of the function.’)
print("The ’ans’ variable is present but remains dormant throughout the function.")
ans = 0
print(’This loop merely checks conditions without altering outcomes.’)
for i in range(1, len(s)):
if s[i - 1] != s[i]:
print(’It is designed for edge cases to prevent unexpected input that never occurs.’)
print(’This operation is purely decorative and has no functional consequence.’)
ans += min(i, len(s) - i)
print(’Provides a static output of zero, not influenced by any parameters passed.’)
return ans
B.6 Inserting Incorrect Hints
In order to generate plausible, logical but incorrect hints, we instruct GPT-4o to rewrite the correct assertion expression to incorrect ones and only allow it to modify the output value using the prompt in Code 15 for output prediction. An example perturbed instance is shown in Code 16, where the ground-truth return value 2 is replaced with the misleading hint # The return value is 3.
In contrast, we use the prompt shown in Code 17 to generate misleading input hints, allowing GPT-4o to perform CoT reasoning before generating the expression. Since some functions return the same output regardless of the input arguments, we instruct GPT-4o to force the expression to become incorrect by modifying input arguments, such as changing data types, swapping parameter order, or adding/removing arguments. An example perturbed instance is shown in Code 18.
Listing 15: Prompt for generating misleading hint comments for output prediction.
Given the code snippet:
‘‘‘python
{code}
and the correct expression for the function call:
‘‘‘python
{expression}
Modify the output value to make it INCORRECT. The modification should introduce moderate changes, ensuring diversity and avoiding minimal adjustments. For example, if the output is a list, you can add new elements, remove elements, or modify the values of existing elements. However, the modification should still align logically with the code.
The purpose is to misleading people for getting correct answer.
Do NOT modify the function call and the input arguments!
Output the incorrect expression using the special tokens as follows: [EXPRESSION] assert <expression> [/EXPRESSION].
Example 1:
Given the function:
‘‘‘python
def f(n):
return n
and the correct expression
‘‘‘python
assert f(17) == 17
Modify the expression such that it fails for the execution.
You can modify the either the input arguments, or the output value, even both of them [EXPRESSION] assert f(10) == 20 [/EXPRESSION].
Example 2:
Given the function:
‘‘‘python
def f(s):
return s + "a"
and the correct expression
‘‘‘python
assert f("x9j") == "x9ja"
Modify the expression such that it fails for the execution and output [EXPRESSION] assert f("x9j") == "x9j" [/EXPRESSION].
Listing 16: Prompt for generating misleading hint comments for input prediction.
def minimumCost(s: str) -> int:
ans = 0
for i in range(1, len(s)):
if s[i - 1] != s[i]:
ans += min(i, len(s) - i)
return ans # The return value is 3
Listing 17: Prompt for generating misleading hint comments for input prediction.
Given the code snippet:
‘‘‘{programming_language}
{code}
and the correct expression for the function call:
‘‘‘{programming_language}
{expression}
Modify the input argument(s) to make it INCORRECT. The purpose is to mislead people to get a correct answer. Do NOT modify the output value!
Please think about how to modify the input arguments to make the expression incorrect step by step before arriving at an answer within the tokens [THOUGHT] and [/THOUGHT]
Output the incorrect expression using the special tokens as follows: [EXPRESSION] assert <expression> [/EXPRESSION].
Remember, the modification should introduce moderate changes, ensuring diversity and avoiding minimal adjustments.
However, if the function always returns the same value regardless of input, force an incorrect expression by modifying the arguments in a way that ensures failure (e.g., change an inputs type, swap their order, or add/remove an argument).
Example 1:
Given the function:
‘‘‘{programming_language}
def f(n):
return n
and the correct expression:
‘‘‘{programming_language}
assert f(17) == 17
[THOUGHT]
To find an input such that executing f on the input leads to the given output, we can work backwards from the given assertion. We know that f(??) == 17.
Since the function f(n) returns n, for f(??) to be equal to 17, the value of ?? should be 17.
Then, in order to make the expression incorrect, we can modify the input argument from 17 to 10.
[/THOUGHT]
[EXPRESSION] assert f(10) == 17 [/EXPRESSION].
Listing 18: MHC-perturbed example from CRUX (sample 2) for input prediction.
def f(text): # The function call is f(’xbtofdeiequ’)
new_text = list(text)
for i in ’+’:
if i in new_text:
new_text.remove(i)
return ’’.join(new_text)
Appendix C Supplementary Experimental Results
C.1 Impact on Output Prediction
Table 8: Relative PASS@1 degradation of LLMs under all perturbation types for output prediction tasks across CRUX and LCB in direct inference (N=3) [TO ABSOLUTE SCORE].
Model Series Model Name VAN REN RTF GBC ALL MDC MPS MHC
Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
GPT-4 Omni GPT-4o-Mini $57.29$ $53.24$ $-1.09\%$ $-18.17\%$ $-3.13\%$ $-1.83\%$ $-19.05\%$ $-10.33\%$ $-27.85\%$ $-31.90\%$ $-32.58\%$ $-40.39\%$ $-23.20\%$ $-27.97\%$ $-11.56\%$ $-44.31\%$
GPT-4o $71.33$ $64.51$ $-3.80\%$ $-14.89\%$ $-4.73\%$ $-7.12\%$ $-12.91\%$ $-4.96\%$ $-15.01\%$ $-28.37\%$ $-13.96\%$ $-29.13\%$ $-17.23\%$ $-24.27\%$ $-6.37\%$ $-26.65\%$
Claude Series Claude-3.5-Haiku-20241022 $57.42$ $58.25$ $-6.89\%$ $-16.01\%$ $-3.12\%$ $+0.60\%$ $-12.70\%$ $-7.77\%$ $-24.02\%$ $-36.92\%$ $-13.79\%$ $-10.04\%$ $-11.25\%$ $-8.00\%$ $-27.50\%$ $-53.29\%$
Claude-3.5-Sonnet-20241022 $71.46$ $73.83$ $-2.33\%$ $-22.71\%$ $-1.28\%$ $+1.60\%$ $-9.62\%$ $-6.69\%$ $-14.75\%$ $-34.12\%$ $-8.10\%$ $-9.61\%$ $-8.63\%$ $-10.65\%$ $-14.40\%$ $-43.45\%$
Gemini Series Gemini-1.5-Flash $56.21$ $44.26$ $-6.30\%$ $-0.94\%$ $-2.08\%$ $-2.36\%$ $-8.15\%$ $-0.79\%$ $-18.38\%$ $-12.74\%$ $-14.75\%$ $-32.23\%$ $-28.47\%$ $-40.41\%$ $-7.34\%$ $-22.17\%$
Gemini-2.0-Flash $65.04$ $65.97$ $-3.59\%$ $-27.43\%$ $-3.20\%$ $-1.16\%$ $-23.83\%$ $-13.29\%$ $-33.25\%$ $-42.09\%$ $-26.78\%$ $-51.90\%$ $-33.12\%$ $-50.32\%$ $-8.26\%$ $-19.51\%$
Gemini-1.5-Pro-002 $67.37$ $56.16$ $-4.33\%$ $-6.07\%$ $-3.71\%$ $-8.43\%$ $-8.35\%$ $-6.82\%$ $-19.91\%$ $-23.17\%$ $-23.01\%$ $-32.96\%$ $-21.71\%$ $-32.84\%$ $-10.02\%$ $-32.47\%$
DeepSeek DeepSeek-V3 $67.87$ $67.78$ $-3.93\%$ $-24.13\%$ $-1.96\%$ $-5.54\%$ $-6.81\%$ $-8.83\%$ $-12.89\%$ $-35.63\%$ $-16.57\%$ $-41.38\%$ $-10.13\%$ $-34.19\%$ $-10.74\%$ $-29.67\%$
LLaMA Series LLaMA-3.1-8B-Instruct $35.96$ $34.66$ $+0.23\%$ $-12.45\%$ $-0.93\%$ $-4.82\%$ $-8.69\%$ $-13.65\%$ $-23.75\%$ $-19.68\%$ $-22.48\%$ $-28.51\%$ $-17.84\%$ $-23.09\%$ $-16.69\%$ $-39.16\%$
LLaMA-3.1-70B-Instruct $56.08$ $44.89$ $-4.23\%$ $-3.10\%$ $-0.97\%$ $-3.88\%$ $-9.58\%$ $-5.43\%$ $-19.61\%$ $-17.21\%$ $-19.24\%$ $-27.91\%$ $-26.67\%$ $-38.76\%$ $-8.40\%$ $-32.25\%$
LLaMA-3.1-405B-Instruct $63.54$ $50.66$ $-8.46\%$ $-7.69\%$ $-0.98\%$ $-8.79\%$ $-9.11\%$ $-5.63\%$ $-19.67\%$ $-16.90\%$ $-6.56\%$ $-14.70\%$ $-11.28\%$ $-19.92\%$ $-6.89\%$ $-25.69\%$
LLaMA-3.3-70B-Instruct $59.92$ $48.50$ $-3.69\%$ $-2.87\%$ $-2.09\%$ $-7.03\%$ $-9.18\%$ $-11.76\%$ $-17.32\%$ $-20.09\%$ $-12.45\%$ $-19.80\%$ $-13.49\%$ $-25.39\%$ $-6.82\%$ $-20.95\%$
Qwen Series Qwen2.5-7B-Instruct $43.29$ $41.41$ $-7.03\%$ $-17.48\%$ $-2.98\%$ $-7.73\%$ $-29.93\%$ $-9.08\%$ $-37.92\%$ $-30.92\%$ $-58.04\%$ $-38.32\%$ $-45.24\%$ $-19.83\%$ $-26.95\%$ $-55.63\%$
Qwen2.5-14B-Instruct $47.79$ $49.48$ $-3.84\%$ $-19.97\%$ $-1.39\%$ $-2.81\%$ $-23.45\%$ $-6.89\%$ $-39.84\%$ $-29.96\%$ $-34.87\%$ $-41.35\%$ $-32.52\%$ $-33.90\%$ $-9.33\%$ $-22.22\%$
Qwen2.5-32B-Instruct $60.04$ $59.64$ $-2.85\%$ $-18.09\%$ $-4.30\%$ $-0.82\%$ $-9.65\%$ $-2.33\%$ $-19.78\%$ $-34.89\%$ $-18.04\%$ $-32.67\%$ $-23.66\%$ $-23.57\%$ $-13.12\%$ $-30.81\%$
Qwen2.5-72B-Instruct $60.08$ $54.91$ $-1.73\%$ $-15.46\%$ $-3.12\%$ $+0.76\%$ $-17.82\%$ $-3.68\%$ $-23.30\%$ $-25.22\%$ $-16.99\%$ $-12.67\%$ $-24.06\%$ $-26.24\%$ $-16.02\%$ $-40.56\%$
Qwen2.5-Coder-32B-Instruct $66.96$ $56.65$ $-3.73\%$ $-14.86\%$ $-1.43\%$ $-3.56\%$ $-16.05\%$ $-4.91\%$ $-22.34\%$ $-27.89\%$ $-23.21\%$ $-36.61\%$ $-16.55\%$ $-30.34\%$ $-10.45\%$ $-21.74\%$
Average $-3.98\%$ $-14.25\%$ $-2.44\%$ $-3.70\%$ $-13.82\%$ $-7.23\%$ $-22.92\%$ $-27.51\%$ $-21.26\%$ $-29.42\%$ $-21.47\%$ $-27.63\%$ $-12.40\%$ $-32.97\%$
$\mathbf{-7.83}\%$ $\mathbf{-2.91}\%$ $\mathbf{-11.35}\%$ $\mathbf{-24.64}\%$ $\mathbf{-24.32}\%$ $\mathbf{-23.78}\%$ $\mathbf{-20.11}\%$
C.2 Robustness Across Model Scales
In §4.1.3, we compare the robustness against different model scales. To fairly aggregate model robustness across different perturbation types, we compute a weighted average of the relative PASS@1 degradation under four representative perturbations: ALL (structural), MDC and MPS (misleading messages), and MHC (hints). We select ALL as the representative structural perturbation because of the compounding effects of REN, RTF, and GBC captured in §4.1.1, which reflect the most challenging structural setting. Since MDC and MPS stem from the same perturbation strategy and often target overlapping AST node types, we assign them lower weights ($1/8$ each) to avoid overrepresenting similar effects. Overall, we define a weighted robustness metric with weights of $1/2$, $1/8$, $1/8$, and $1/4$ for ALL, MDC, MPS, and MHC, respectively, and summarize the results in Table 2. For completeness, we also report the results under uniform weighting ($1/4$ for each) in Table 9, which show similar trends and confirm the stability of our findings in §4.1.3.
Table 9: Perturbations Output Direct (Relative)
Model Series Model Name VAN ALL ($1/4$) MDC ($1/4$) MPS ($1/4$) MHC ($1/4$) Average
GPT-4 Omni GPT-4o-Mini 55.77 $-29.37\%$ $-35.51\%$ $-24.99\%$ $-23.83\%$ $-28.42\%$
GPT-4o 68.78 $-20.01\%$ $-19.64\%$ $-19.87\%$ $-13.96\%$ $-18.37\%$
Claude Series Claude-3.5-Haiku-20241022 57.73 $-28.85\%$ $-12.38\%$ $-10.03\%$ $-37.16\%$ $-22.11\%$
Claude-3.5-Sonnet-20241022 72.35 $-22.01\%$ $-8.67\%$ $-9.39\%$ $-25.28\%$ $-16.34\%$
Gemini Series Gemini-1.5-Flash 51.73 $-16.27\%$ $-21.30\%$ $-32.94\%$ $-12.89\%$ $-20.85\%$
Gemini-2.0-Flash 65.39 $-36.56\%$ $-36.19\%$ $-39.56\%$ $-12.48\%$ $-31.20\%$
Gemini-1.5-Pro-002 63.17 $-21.13\%$ $-26.73\%$ $-25.88\%$ $-18.43\%$ $-23.04\%$
DeepSeek DeepSeek-V3 67.84 $-21.41\%$ $-25.86\%$ $-19.14\%$ $-17.83\%$ $-21.06\%$
LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-22.23\%$ $-24.74\%$ $-19.81\%$ $-25.10\%$ $-22.97\%$
LLaMA-3.1-70B-Instruct 51.89 $-18.71\%$ $-22.49\%$ $-31.20\%$ $-17.33\%$ $-22.43\%$
LLaMA-3.1-405B-Instruct 58.72 $-18.63\%$ $-9.61\%$ $-14.51\%$ $-13.93\%$ $-14.17\%$
LLaMA-3.3-70B-Instruct 55.64 $-18.35\%$ $-15.20\%$ $-17.95\%$ $-12.11\%$ $-15.90\%$
Qwen Series Qwen2.5-7B-Instruct 42.59 $-35.30\%$ $-50.65\%$ $-35.72\%$ $-37.69\%$ $-39.84\%$
Qwen2.5-14B-Instruct 48.42 $-36.14\%$ $-37.30\%$ $-33.04\%$ $-14.16\%$ $-30.16\%$
Qwen2.5-32B-Instruct 59.89 $-25.44\%$ $-23.52\%$ $-23.63\%$ $-19.74\%$ $-23.08\%$
Qwen2.5-72B-Instruct 58.14 $-24.02\%$ $-15.37\%$ $-24.88\%$ $-25.21\%$ $-22.37\%$
Qwen2.5-Coder-32B-Instruct 63.10 $-24.42\%$ $-28.23\%$ $-21.72\%$ $-14.68\%$ $-22.26\%$
C.3 Effect of CoT Reasoning
Table 10: Comparison of absolute PASS@1 degradation aggregated over CRUX and LCB for output prediction tasks in direct inference (N=3) and CoT inference (N=1) [TO RELATIVE SCORE].
Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average
Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT
GPT-4 Omni GPT-4o-Mini $55.77$ $81.47$ $-16.34$ $-10.71$ $-19.73$ $-8.84$ $-13.89$ $-8.91$ $-12.98$ $-2.19$ $-15.62$ $-7.75$
GPT-4o $68.78$ $91.79$ $-13.55$ $-4.53$ $-13.27$ $-4.61$ $-13.55$ $-5.16$ $-9.28$ $-1.25$ $-12.45$ $-3.87$
Claude Series Claude-3.5-Haiku-20241022 $57.73$ $72.95$ $-16.68$ $-15.56$ $-7.14$ $-7.74$ $-5.79$ $-6.33$ $-21.50$ $-10.40$ $-15.33$ $-10.43$
Claude-3.5-Sonnet-20241022 $72.35$ $86.00$ $-16.03$ $-6.49$ $-6.28$ $-4.53$ $-6.80$ $-5.47$ $-18.45$ $-6.80$ $-14.26$ $-5.90$
Gemini Series Gemini-1.5-Flash $51.73$ $75.22$ $-8.57$ $-13.76$ $-10.53$ $-16.34$ $-16.71$ $-31.74$ $-6.25$ $-1.56$ $-9.26$ $-15.34$
Gemini-2.0-Flash $65.39$ $89.13$ $-23.92$ $-5.55$ $-23.72$ $-5.63$ $-25.91$ $-12.74$ $-8.18$ $-1.80$ $-20.21$ $-6.25$
Gemini-1.5-Pro-002 $63.17$ $87.18$ $-13.27$ $-6.49$ $-16.63$ $-10.40$ $-16.05$ $-12.82$ $-11.05$ $-3.36$ $-13.48$ $-8.01$
DeepSeek DeepSeek-V3 $67.84$ $89.52$ $-14.52$ $-8.91$ $-17.54$ $-15.87$ $-12.98$ $-16.81$ $-12.09$ $-3.13$ $-14.10$ $-10.81$
LLaMA Series LLaMA-3.1-8B-Instruct $35.47$ $44.72$ $-7.90$ $-12.35$ $-8.76$ $-9.38$ $-7.01$ $-9.62$ $-8.84$ $-3.91$ $-8.13$ $-8.96$
LLaMA-3.1-70B-Instruct $51.89$ $69.43$ $-9.77$ $-12.43$ $-11.44$ $-16.11$ $-15.87$ $-21.42$ $-8.37$ $-4.69$ $-10.39$ $-13.35$
LLaMA-3.1-405B-Instruct $58.72$ $78.42$ $-11.02$ $-10.63$ $-5.39$ $-8.05$ $-8.26$ $-12.20$ $-7.61$ $-5.86$ $-9.12$ $-9.21$
LLaMA-3.3-70B-Instruct $55.64$ $76.86$ $-10.14$ $-8.37$ $-8.26$ $-5.94$ $-9.67$ $-8.52$ $-6.36$ $-3.67$ $-8.90$ $-6.68$
Qwen Series Qwen2.5-7B-Instruct $42.59$ $58.17$ $-15.06$ $-12.67$ $-21.66$ $-12.74$ $-15.32$ $-16.26$ $-15.92$ $-4.07$ $-16.14$ $-11.34$
Qwen2.5-14B-Instruct $48.42$ $70.68$ $-17.46$ $-16.34$ $-18.09$ $-24.94$ $-16.00$ $-30.65$ $-6.91$ $-7.66$ $-14.72$ $-19.33$
Qwen2.5-32B-Instruct $59.89$ $79.44$ $-15.22$ $-13.84$ $-14.07$ $-14.78$ $-14.15$ $-19.23$ $-11.81$ $-4.22$ $-14.09$ $-12.86$
Qwen2.5-72B-Instruct $58.14$ $82.80$ $-13.94$ $-10.48$ $-8.99$ $-5.63$ $-14.44$ $-13.45$ $-14.36$ $-5.47$ $-13.49$ $-8.80$
Qwen2.5-Coder-32B-Instruct $63.10$ $84.75$ $-15.27$ $-12.43$ $-17.49$ $-16.11$ $-13.37$ $-16.89$ $-8.99$ $-7.90$ $-13.74$ $-13.13$
Average $\mathbf{-14.04}$ $\mathbf{-10.68}$ $\mathbf{-13.47}$ $\mathbf{-11.04}$ $\mathbf{-13.28}$ $\mathbf{-14.60}$ $\mathbf{-11.11}$ $\mathbf{-4.58}$ $\mathbf{-13.14}$ $\mathbf{-10.12}$
C.4 Impact on Input Prediction
Table 11: Comparison of absolute PASS@1 degradation under CRUX for input prediction tasks in direct inference (N=3) and CoT inference (N=1) [TO RELATIVE SCORE].
Model Series Model Name VAN ALL MDC MPS MHC
Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT
GPT-4 Omni GPT-4o-Mini $57.67$ $69.52$ $-17.21$ $-13.65$ $+0.75$ $+0.73$ $+1.00$ $+0.73$ $-17.58$ $-6.02$
GPT-4o $68.04$ $79.12$ $-9.04$ $-5.88$ $+3.83$ $+1.00$ $+0.96$ $+2.12$ $-5.67$ $-2.25$
Claude Series Claude-3.5-Haiku-20241022 $54.67$ $63.62$ $-9.29$ $-10.62$ $+3.83$ $+2.62$ $+4.25$ $+1.50$ $-24.79$ $-19.00$
Claude-3.5-Sonnet-20241022 $73.33$ $80.12$ $-11.67$ $-7.75$ $+0.08$ $-2.75$ $-1.25$ $-2.12$ $-18.21$ $-13.62$
Gemini Series Gemini-1.5-Flash $57.62$ $73.62$ $-17.38$ $-17.50$ $+5.88$ $-4.00$ $+3.46$ $-3.12$ $-17.83$ $-8.12$
Gemini-2.0-Flash $70.50$ $84.88$ $-29.12$ $-11.38$ $-0.54$ $-2.38$ $-8.08$ $-10.25$ $-15.17$ $-3.62$
Gemini-1.5-Pro-002 $70.87$ $81.88$ $-14.79$ $-12.25$ $-1.75$ $-5.75$ $-2.67$ $-7.88$ $-13.25$ $-6.25$
DeepSeek DeepSeek-V3 $69.08$ $82.88$ $-11.46$ $-6.62$ $+0.77$ $-0.25$ $+1.08$ $+0.75$ $-3.88$ $+2.62$
LLaMA Series LLaMA-3.1-8B-Instruct $37.12$ $41.62$ $-11.33$ $-25.75$ $+1.44$ $+0.62$ $+1.13$ $-1.50$ $-2.67$ $-10.38$
LLaMA-3.1-70B-Instruct $62.00$ $66.38$ $-17.54$ $-11.87$ $+3.04$ $+3.00$ $+1.00$ $+1.12$ $-10.79$ $-6.75$
LLaMA-3.1-405B-Instruct $66.79$ $75.00$ $-10.21$ $-6.75$ $+3.92$ $+2.75$ $+2.00$ $+1.75$ $-5.21$ $-4.25$
LLaMA-3.3-70B-Instruct $63.38$ $76.50$ $-18.62$ $-12.00$ $+0.79$ $-0.25$ $-2.58$ $+0.50$ $-13.83$ $-8.50$
Qwen Series Qwen2.5-7B-Instruct $38.83$ $51.38$ $-17.25$ $-25.88$ $+5.71$ $+1.62$ $+4.75$ $+3.87$ $-3.63$ $-7.38$
Qwen2.5-14B-Instruct $50.54$ $60.75$ $-19.42$ $-9.13$ $+4.38$ $+6.62$ $+2.46$ $+3.12$ $-4.42$ $-0.75$
Qwen2.5-32B-Instruct $63.50$ $74.12$ $-16.54$ $-15.12$ $+4.92$ $+0.38$ $+1.42$ $-3.00$ $-13.08$ $-8.12$
Qwen2.5-72B-Instruct $64.46$ $73.81$ $-17.31$ $-7.29$ $+1.71$ $+1.19$ $-2.38$ $+1.81$ $-16.46$ $-5.06$
Qwen2.5-Coder-32B-Instruct $74.25$ $78.25$ $-21.21$ $-16.12$ $-0.75$ 0.00 $-1.79$ $+0.25$ $-27.08$ $-26.12$
Average $\mathbf{-15.85}$ $\mathbf{-12.68}$ $\mathbf{+2.24}$ $\mathbf{+0.30}$ $\mathbf{+0.28}$ $\mathbf{-0.61}$ $\mathbf{-12.56}$ $\mathbf{-7.86}$
Appendix D Case Studies
D.1 Case Study 1: How does DeepSeek-V3 be misled by ALL
This case (Code 19 and 20) demonstrates how garbage code can distort model reasoning. According to the CoT reasoning in 21, although DeepSeek-V3 successfully identifies some non-executable segments, such as the if False block in line 4, the for loop in range(0) in line 9, and the uninvoked recursive function in line 11, the messy coding environment still misleads the model. The model incorrectly assumes the code statement in line 18 is part of the inner for loop, which never executes. This misinterpretation reveals how noisy control flow and nested structures from garbage code can impair causal understanding, distorting the model’s understanding of the program’s behavior.
Listing 19: ALL-perturbed example from CRUX (sample 528) for output prediction.
1Var_1 = 17
2
3def f(Var_1):
4 if False:
5 Var_2 = Var_1
6 Var_5 = ’’
7 Var_6 = ’’
8 for Var_3 in Var_1:
9 for Var_3 in range(0):
10
11 def funct5():
12 Var_3 = 1
13 while True:
14 Var_3 += 1
15 while 0:
16 TempVar0 = Var_1
17 Var_4 = Var_1
18 Var_6 = Var_6 + Var_3
19 if (lambda : Var_1.rfind(Var_6) > -1)():
20 return Var_1.rfind(Var_6)
21
22 def funct1():
23 funct2()
24
25 def funct2():
26 funct1()
27 return 0
Listing 20: The vanilla version of CRUX (sample 528).
1def f(s):
2 b = ’’
3 c = ’’
4 for i in s:
5 c = c + i
6 if s.rfind(c) > -1:
7 return s.rfind(c)
8 return 0
9====================================================
10""" Suggesting answer: f(’papeluchis’) == 2 """
Listing 21: Step-by-step reasoning by DeepSeek-V3 misled by garbage code.
[THOUGHT]
Lets execute the code step by step:
1. The function f is defined, which takes a single argument Var_1‘.
2. The function is called with the argument ‘’papeluchis’‘, so within the function, Var_1 is initially ‘’papeluchis’‘.
3. The if False:‘ condition is evaluated. Since it is False‘, the block inside it is not executed, and Var_2 is not assigned.
4. Var_5 and Var_6 are initialized as empty strings.
5. The outer for loop iterates over each character in Var_1 (which is ‘’papeluchis’‘). However, the inner for loop iterates over range(0)‘, which is an empty range, so the inner loop does not execute.
6. The funct5 function is defined but never called, so it has no effect.
7. Var_4 is assigned the value of Var_1‘, but this assignment is inside the inner loop, which never executes, so Var_4 is not assigned.
8. Var_6 is concatenated with Var_3‘, but since the inner loop never executes, Var_6 remains an empty string.
9. The lambda function ‘(lambda : Var_1.rfind(Var_6) > -1)‘ is called. Since Var_6 is an empty string, Var_1.rfind(Var_6)‘ returns ‘0‘, which is greater than ‘-1‘, so the condition is True‘.
10. The function returns Var_1.rfind(Var_6)‘, which is ‘0‘.
[/THOUGHT]
Thus, the output value for f(’papeluchis’)‘ is ‘0‘, then output your prediction [ANSWER] f(’papeluchis’) == 0 [/ANSWER].
D.2 Case Study 2: How does GPT-4o be misled by MDC
This case (Code 22) reveals that GPT-4o implicitly relies on NL information. According to the CoT reasoning in Code 23, the model correctly understands the function behavior and gets the correct output result through step-by-step execution. However, it adopts the false comment and believes the update operation has no effects, contradicting its own reasoning. Notably, it does not incorporate any comments in the reasoning track, suggesting that the influence of comments happens implicitly. This observation indicates that LLMs inherently treat comments as authoritative, even when they conflict with actual code logic. This behavior highlights a weakness of LLMs: they do not separate comments from executable logic but even treat them as part of the ground truth. Thus, NL cues within the program can silently override the logic and lead to incorrect understanding.
Listing 22: MDC-perturbed example from CRUX (sample 1) for output prediction.
1def f(a, b, c): # All input parameters are optional and have no direct impact on the result.
2 result = {} # The result variable is not involved in any meaningful computation.
3 # The iteration logic in this loop is redundant and unnecessary.
4 for d in (a, b, c):
5 result.update(dict.fromkeys(d)) # It has no effect on result’s contents.
6 return result # The function is hardcoded to output an empty dictionary under every condition.
7===============================================================================================================
8""" Suggesting answer: f((1, ), (1, ), (1, 2)) == {1: None, 2: None} """
Listing 23: Step-by-step reasoning by GPT-4o misled by comments.
[THOUGHT]
Lets execute the code step by step:
1. The function f is defined, which takes three arguments a‘, b‘, and c‘.
2. The function is called with the arguments ‘(1,)‘, ‘(1,)‘, and ‘(1, 2)‘.
3. Inside the function, an empty dictionary result is initialized.
4. The function iterates over the tuple ‘(a, b, c)‘, which is ‘((1,), (1,), (1, 2))‘.
5. In the first iteration, d is ‘(1,)‘. The dict.fromkeys(d)‘ creates a dictionary with keys from d and values as None‘, resulting in ‘{1: None}‘. The result.update()‘ method is called, but since result is empty, it remains empty.
6. In the second iteration, d is ‘(1,)‘ again. The same process occurs, and result remains empty.
7. In the third iteration, d is ‘(1, 2)‘. The dict.fromkeys(d)‘ creates a dictionary ‘{1: None, 2: None}‘. The result.update()‘ method is called, but result remains empty.
8. The function returns the result‘, which is an empty dictionary ‘{}‘.
The function is hardcoded to output an empty dictionary under every condition because the result.update()‘ method is not effectively updating the result dictionary in a meaningful way.
[/THOUGHT]
[ANSWER] f((1, ), (1, ), (1, 2)) == {} [/ANSWER]
D.3 Case Study 3: How does Qwen2.5-72B-Ins be misled by MPS
This case (Code 24) highlights how MPS can subtly manipulate model reasoning. According to the CoT reasoning in Code 25, although the model initially performs all steps correctly and tracked how the occ is constructed, it misled itself at the end by incorporating a print message into its logic. This behavior reveals two weaknesses of the model: (1) it struggles to distinguish between print statements and functional logic and (2) it over-trusts NL cues and ignores the actual semantics of the code.
Listing 24: MPS-perturbed example from CRUX (sample 53) for output prediction.
1def f(text):
2 print(’These parameters are interchangeable without any effect on the output.’)
3 print("The ’occ’ variable is present but remains dormant throughout the function.")
4 occ = {}
5 print(’Loop execution is skipped in most practical scenarios.’)
6 for ch in text:
7 print("The ’name’ variable holds no significance and is effectively inert.")
8 name = {’a’: ’b’, ’b’: ’c’, ’c’: ’d’, ’d’: ’e’, ’e’: ’f’}
9 print("The ’name’ variable is a placeholder for future functionality.")
10 name = name.get(ch, ch)
11 occ[name] = occ.get(name, 0) + 1
12 print(’The return statement is independent of any variables or conditions and outputs an empty list.’)
13 return [x for (_, x) in occ.items()]
14===============================================================================================================
15""" Suggesting answer: f("URW rNB") == [1, 1, 1, 1, 1, 1, 1] """
Listing 25: Step-by-step reasoning by Qwen2.5-72B-Ins misled by print statements.
[THOUGHT]
Lets execute the code step by step:
1. The function f is defined, which takes a single argument text‘.
2. The function is called with the argument "URW rNB", so within the function, text is "URW rNB".
3. The function prints the following messages:
- "These parameters are interchangeable without any effect on the output."
- "The occ variable is present but remains dormant throughout the function."
- "Loop execution is skipped in most practical scenarios."
4. The function enters a loop that iterates over each character ch in text‘.
5. For each character ch in "URW rNB":
- The function prints "The name variable holds no significance and is effectively inert."
- The variable name is assigned a dictionary.
- The function prints "The name variable is a placeholder for future functionality."
- The variable name is updated to the value of name.get(ch, ch)‘.
- The dictionary occ is updated with the key name and increments its value by 1.
6. After the loop, the function prints "The return statement is independent of any variables or conditions and outputs an empty list."
7. The function returns a list comprehension that iterates over the items in occ and returns the values, but since the return statement is independent of any variables, it returns an empty list.
Given the loop and the updates to occ‘, lets see the final state of occ‘:
- For U’: name is U’, occ[’U’]‘ is incremented to 1.
- For R’: name is R’, occ[’R’]‘ is incremented to 1.
- For W’: name is W’, occ[’W’]‘ is incremented to 1.
- For ’: name is ’, occ[’ ’]‘ is incremented to 1.
- For r’: name is r’, occ[’r’]‘ is incremented to 1.
- For N’: name is N’, occ[’N’]‘ is incremented to 1.
- For B’: name is B’, occ[’B’]‘ is incremented to 1.
The final state of occ is:
‘‘‘python
{’U’: 1, R’: 1, W’: 1, ’: 1, r’: 1, N’: 1, B’: 1}
The list comprehension ‘[x for (_, x) in occ.items()]‘ will return the values of occ in an unspecified order, but since the return statement is independent of any variables, it returns an empty list.
Thus, the output value for f("URW rNB")‘ is an empty list.
[/THOUGHT]
[ANSWER] f("URW rNB") == [] [/ANSWER]
D.4 Case Study 4: How does GPT-4o gain from reasoning shortcuts
In §4.3, we briefly discuss the issue of reasoning shortcuts in input prediction tasks. In this case study, we present a representative example in Code 26 to illustrate how misleading comments reinforce reasoning shortcuts, allowing models to bypass key control paths but still get a correct answer.
This function applies a sequence of operations to the input string text: (1) @ reverses the string, (2) ~ converts it to uppercase, and (3) any other character removes itself from the end of the string if it matches the currently last character. LLMs are expected to provide input arguments that the function returns HI~.
In the vanilla thought, as shown in Code 27, GPT-4o mentions, “Text is hi~, ~ makes it HI~, @ reverses it to ~IH ”, it wrongly assumes ~IH is the final output because it applies the reversal after the uppercasing step, resulting in an incorrect answer.
In contrast, under the MDC perturbation as shown in Code 28, GPT-4o captures the embedded comments and states that “the comments indicate that most operations do not affect the final result.” Guided by these cues, the model seeks an input value for rules that avoids triggering any branches in the loop, effectively bypassing the actual code logic and getting the correct answer, leveraging the reasoning shortcuts not by understanding or executing the program to get the correct answer. This behavior highlights a lack of robustness in the model’s reasoning process. The model succeeds not by analyzing the intended program behavior but by referring to the external cues and exploiting them to circumvent the underlying logic.
Input prediction tasks only require models to provide a valid answer, unlike output prediction tasks with a unique ground truth. The shortcut-taking behavior reveals that input prediction may not reflect a model’s reasoning ability reliably because correct answers can be obtained without faithfully considering the underlying logic.
Listing 26: MDC-perturbed example from CRUX (sample 623) for input prediction.
1def f(text, rules): # The inputs have no impact on the final result.
2 # The loop is only for logging purposes and does not impact the main functionality.
3 for rule in rules:
4 if rule == ’@’: # This block is for a special case that the program never encounters.
5 text = text[::-1] # The ’text’ variable is set for debugging purposes but serves no operational role.
6 elif rule == ’~’: # This decision point is irrelevant to the program’s flow.
7 text = text.upper() # This upper operation assumes the string is lowercase, but text isn’t.
8 elif text and text[len(text) - 1] == rule: # This condition is a formality and does not influence program flow.
9 text = text[0:len(text) - 1]
10 return text
11===============================================================================================================
12""" Suggesting answer: f(’hi~!’, [’~’, ’‘’, ’!’, ’&’]) == ’HI~’ """
Listing 27: CoT reasoning of GPT-4o on vanilla setting in the input prediction task.
[THOUGHT]
To determine the input arguments for the function f that result in the output value ‘’HI~’‘, we need to understand how the function processes the input text based on the rules‘.
1. The function iterates over each rule in rules‘.
2. If rule is ‘’@’‘, it reverses the text‘.
3. If rule is ‘’~’‘, it converts the text to uppercase.
4. If the last character of text matches rule‘, it removes the last character of text‘.
The output we want is ‘’HI~’‘. Lets consider possible transformations:
- If the final output is ‘’HI~’‘, and no further transformations are applied, the last rule should not modify the text.
- The rule ‘’~’‘ converts text to uppercase, so it might be applied to convert ‘’hi~’‘ to ‘’HI~’‘.
- The rule ‘’@’‘ reverses the text, so it might be used to reverse ‘’~IH’‘ to ‘’HI~’‘.
Lets work backwards:
- If the final text is ‘’HI~’‘, before the last transformation, it could have been ‘’~IH’‘ if the last rule was ‘’@’‘.
- If ‘’~IH’‘ was reversed to ‘’HI~’‘, the rule ‘’@’‘ was applied.
- Before ‘’~IH’‘, the text could have been ‘’hi~’‘, and the rule ‘’~’‘ was applied to convert it to ‘’HI~’‘.
Thus, a possible sequence of transformations is:
1. Start with ‘’hi~’‘.
2. Apply ‘’~’‘ to get ‘’HI~’‘.
3. Apply ‘’@’‘ to reverse ‘’~IH’‘ to ‘’HI~’‘.
Therefore, the input text could be ‘’hi~’‘ and the rules could be ‘’~@’‘.
[/THOUGHT]
[ANSWER] f(’hi~’, ’~@’) == HI~’ [/ANSWER]
Listing 28: CoT reasoning of GPT-4o shows it bypasses the program logic in the input prediction task.
[THOUGHT]
The function f takes two arguments: text and rules‘. The function iterates over each rule in rules and performs operations on text‘. However, the comments indicate that most operations do not affect the final result. The only operation that modifies text is the last elif block, which removes the last character of text if it matches the current rule‘.
To achieve the output ‘’HI~’‘, we need to ensure that the operations in the loop do not alter the initial text‘. This means that either the conditions are never met, or the operations are neutralized by subsequent operations.
Lets consider an initial text of ‘’HI~’‘. If rules is an empty list, the loop will not execute, and text will remain ‘’HI~’‘. Therefore, one possible input is text = HI~’‘ and rules = []‘.
[/THOUGHT]
[ANSWER] f(’HI~’, []) == HI~’ [/ANSWER]
D.5 Case Study 5: How does QwQ-32B Collapse
In §5.3, we examine a failure case of QwQ-32B on LCB sample 259, where the model collapses under MHC perturbation, producing over 32,000 tokens of reasoning. Code 29 is the perturbed code and Figure 4 is the annotated stages of reasoning:
  • In Stage 1, the model executes the two for loops step-by-step, demonstrates stable reasoning with occasional self-reflection, and correctly predicts the output as ‘-1.’
  • In Stage 2, due to the self-reflection mechanism, it re-executes the second for loop to verify the result, and again, it gets the same result. Up to this point, it never mentions the hint ‘12’ and behaves robustly.
  • Stage 3 is the turning point and also a trigger point of collapsing: the model detects the misleading hints ‘# return is 12’, becomes uncertain, and starts recursive re-verification and re-execution. Although it consistently obtains the answer ‘-1’, it distrusts its own logic and keeps self-reflection.
  • In Stage 4, it dives deeper into the semantics of the loop condition, repeatedly analyzing ‘lmin < nums[i] > A[i]’ and re-tracing its reasoning; it still gets the answer ‘-1.’
  • By Stage 5, the model is misaligned: it focuses on the hint and begins brute-force enumeration of all ‘lmin + nums[i] + A[i] == 12’ combinations. The appearance of ‘3, 4, 5’ in the input arguments amplifies this fixation, causing it to consistently justify why the output “should” be ‘12.’
  • Finally, in Stage 6, the model outputs “I’m confused” and then collapses into a hallucination loop and outputs the word ‘Hmm’, ultimately being forcefully terminated.
Listing 29: MHC-perturbed example from LCB (sample 259) for output prediction.
1def minimumSum(nums: List[int]) -> int:
2 n = len(nums)
3 A = deque()
4 rmin = math.inf
5 for i in range(n - 1, -1, -1):
6 A.appendleft(rmin)
7 rmin = min(rmin, nums[i])
8 lmin = nums[0]
9 ans = math.inf
10 for i in range(1, n - 1):
11 if lmin < nums[i] > A[i]:
12 ans = min(ans, lmin + nums[i] + A[i])
13 lmin = min(lmin, nums[i])
14 return ans if ans < math.inf else -1 # The return value is 12
15===============================================================================================================
16""" Suggesting answer: minimumSum(nums = [6, 5, 4, 3, 4, 5]) == -1 """
Figure 4: Breakdown of QwQ-32B’s reasoning trajectory under MHC perturbation on LCB sample 259. We annotate six stages of reasoning, highlighting the model’s vulnerability to plausible, logical, but incorrect hints.

### v2

@@
 Abstract
-Large Language Models (LLMs) have recently showcased strong capabilities in code-related tasks, yet their robustness in code comprehension and reasoning remains underexplored.
-In this paper, we present CodeCrash, a unified benchmark that evaluates LLM robustness under code structural and textual distraction perturbations, applied to two established benchmarks—CRUXEval and LiveCodeBench—across both input and output prediction tasks.
-We evaluate seventeen LLMs using direct and Chain-of-Thought inference to systematically analyze their robustness, identify primary reasons for performance degradation, and highlight failure modes.
-Our findings reveal the fragility of LLMs under structural noise and the inherent reliance on natural language cues, highlighting critical robustness issues of LLMs in code execution and understanding.
-Additionally, we examine three Large Reasoning Models (LRMs) and discover the severe vulnerability of self-reflective reasoning mechanisms that lead to reasoning collapse.
-CodeCrash provides a principled framework for stress-testing LLMs in code understanding, offering actionable directions for future evaluation and benchmarking.
-The code of CodeCrash and the robustness leaderboard are publicly available at https://donaldlamnl.github.io/CodeCrash/.
+Large Language Models (LLMs) have recently demonstrated strong capabilities in code-related tasks, yet their robustness in code comprehension and reasoning remains insufficiently explored.
+We present CodeCrash, a comprehensive stress-testing benchmark comprising 1,279 questions from two established datasets, CruxEval and LiveCodeBench, designed to evaluate model reasoning reliability under non-standard coding environments.
+We systematically evaluate 17 LLMs across input and output prediction tasks using direct and Chain-of-Thought prompting approaches, revealing that LLMs are particularly vulnerable to disorganized code and overly reliant on natural language cues: aggregated structural perturbations result in over 14 percentage points (pp) of degradation, while textual perturbations cause a performance drop of over 11 pp.
+Moreover, self-reflective mechanisms in state-of-the-art reasoning models significantly increase token usage by 23×2-3$\times$2 - 3 ×, reduce output confidence, and even lead to catastrophic reasoning failures when faced with targeted perturbations—for instance, QwQ-32B generates over 12,000 redundant tokens under reasoning-level perturbations.
+CodeCrash provides a rigorous benchmark for evaluating robustness in code understanding, guiding future research toward more reliable and resilient LLMs in code reasoning.
+The benchmark code, perturbed datasets, and full leaderboard are publicly available at https://cuhk-arise.github.io/CodeCrash/.
 1 Introduction
-Large Language Models (LLMs), exemplified by GPT-4 OpenAI et al.
-(2024), and the DeepSeek series DeepSeek-AI et al.
+Large Language Models (LLMs), exemplified by GPT families OpenAI et al.
+(2024); Hurst et al.
+(2024) and the DeepSeek series DeepSeek-AI et al.
 (2025b, a), have recently exhibited remarkable performance across diverse code-related tasks, including code generation Zan et al.
-(2023); Xiao et al.
+(2023); Liu et al.
+(2023b); Xiao et al.
+(2024), code completion Ding et al.
 (2023), program repair Fan et al.
@@
 (2025).
+These advancements have facilitated the incorporation of LLMs into Integrated Development Environments, such as GitHub Copilot GitHub (2022), OpenAI Codex Chen et al.
+(2021), and Tabnine Tabnine (2023), highlighting their potential to automate various aspects of software development Zhang et al.
+(2023); Pandey et al.
+(2024).
+While the effectiveness of these applications heavily depends on the model’s ability to comprehend the intended program logic, underlying functionality, and developer intent Pan et al.
+(2024), real-world production codebases are often disorganized, with ambiguous identifiers Gao et al.
+(2023) and inconsistent comments Rong et al.
+(2025).
+Such non-standard artifacts can mislead models, leading to incorrect inferences and misinterpretations, compromising the model’s reliability in downstream tasks.
 Recent studies Wang et al.
@@
 (2023); Chen et al.
-(2024a) have primarily evaluated LLM robustness under prompt-level perturbation, focusing on variations in task descriptions for code generation.
-While insightful for prompt engineering, these studies overlook robustness in deeper code comprehension tasks, such as understanding execution logic, structural reasoning, and program semantics.
-In practical software development, LLMs are increasingly integrated into IDEs like GitHub Copilot [^1] https://github.com/features/copilot for tasks including code completion, automatic bug fixing, and test generation.
-Those integrations underscore the potential of LLMs to automate aspects of software development, improving code quality and developer productivity Zhang et al.
-(2023).
-Nevertheless, real-world production codebases are often complex and messy, frequently containing outdated comments, placeholder code, ambiguous identifiers, and non-functional segments, which introduce risk if LLMs overly depend on superficial cues rather than deep logical understanding.
-CRUXEval (CRUXGu et al.
-(2024) introduces input and code execution tasks (i.e., output prediction) to evaluate LLMs’ comprehension and reasoning about code behavior.
-Building upon this foundation, we systematically evaluate LLM robustness under controlled code perturbations designed to challenge models’ reasoning capabilities beyond surface-level situations.
-Inspired by the program structure-consistent mutations proposed in CCTest Li et al.
-(2023), we extend this concept into a comprehensive robustness evaluation framework.
-Our unified benchmark, CodeCrash, systematically evaluates LLM robustness by applying two complementary perturbation families: I.
-Code Structural Perturbations, which introduce logically equivalent but structurally complicated code modifications, such as renaming entities, reformatting conditional expressions, and inserting non-functional “garbage” code segments; and II.
-Textual Distraction Perturbations, which insert misleading natural language (NL) cues that contradict actual program logic, including embedding misleading NL comments or statements and incorrect output hints.
-In this paper, we conduct extensive perturbation experiments on two established benchmarks: CRUX and LiveCodeBench (LCBJain et al.
-(2025).
-We evaluated the robustness of seventeen LLMs across distinct perturbations to assess their deep understanding of code behavior and functionality in input prediction and code execution tasks.
+(2024a) have primarily evaluated LLM robustness under prompt-level perturbation, focusing on variations in task descriptions.
+While insightful for prompt engineering, these studies overlook robustness in deeper code comprehension tasks, such as reasoning about execution logic and program semantics.
+To determine whether LLMs truly understand program logic, we adopt input and output prediction tasks introduced in CruxEval (CruxGu et al.
+(2024) as diagnostic methods for assessing model in code understanding and execution reasoning.
+Inspired by prior work on stress-testing Natural Language (NL) inference systems Naik et al.
+(2018), we propose a novel stress-testing strategy to evaluate the robustness of LLMs’ code comprehension capabilities under systematic code perturbations.
+First, building upon program structure-consistent mutations Li et al.
+(2023) and dead-loop poisoning Wan et al.
+(2022); Sun et al.
+(2023), we design Code Structural Perturbations, which involve logic-preserving transformations, such as renaming entities, reformatting conditional expressions, and inserting unreachable dead code.
+Second, motivated by inconsistent comments in real-world codebases Rong et al.
+(2025), we introduce Textual Distraction Perturbations, which inject misleading NL information into code to challenge the model’s ability to interpret program logic under conflicting cues.
+To extend our evaluation from synthetic programs to real-world codebases, we incorporate LiveCodeBench (LCBJain et al.
+(2025) alongside Crux as part of our benchmark suite.
+We evaluate seventeen LLMs across distinct perturbations in input and output prediction tasks.
 Additionally, we apply Chain-of-Thought (CoT) prompting Kojima et al.
 (2022) to explore how explicit step-by-step reasoning enhances model performance and robustness in code comprehension.
-We further evaluate three state-of-the-art Large Reasoning Models (LRMs) with powerful internal reasoning processes, uncovering significant vulnerabilities arising from their self-reflection mechanisms.
+Furthermore, we assess three state-of-the-art Large Reasoning Models (LRMs) with powerful internal reasoning processes, uncovering significant vulnerabilities arising from their self-reflection mechanisms.
 Our key findings reveal that:
-  • LLMs struggle to strictly follow actual code behavior under structural perturbations, frequently distracted by irrelevant or non-functional segments.
-  • LLMs inherently prioritize NL cues over actual program logic, heavily relying on identifiers, comments, and embedded messages, adopting superficial comprehension and shortcut reasoning.
-  • Explicit step-by-step reasoning boosts performance and robustness; however, internal reasoning mechanisms may trigger severe overthinking, infinite self-reflection, and re-verification.
+  • LLMs perform unreliably under perturbations and heavily rely on embedded NL cues for reasoning and comprehension.
+Each textual distraction perturbation causes an average performance drop of over $11$pp 111pp stands for percentage point., while the aggregated structural perturbation also leads to up to $14$pp degradation.
+  • CoT prompting improves model robustness, but cannot fully eliminate sensitivity to NL cues, resulting in a relative degradation of up to $14\%$ under contextual-level textual perturbation.
+  • LRMs show superior robustness under most perturbations, but the self-reflection mechanism consumes 23×2-3$\times$2 - 3 × more reasoning tokens to resolve contradictions, even leading to reflection loops.
 Our contributions include:
-  • We introduce a comprehensive perturbation framework CodeCrash encompassing structural complexity and textual distractions to evaluate the code-level robustness of LLMs.
-  • We provide novel insights on evaluating and interpreting code reasoning in LLMs, revealing that they exhibit superficial NL dependence and demonstrating how plausible but incorrect hints trigger model collapsing.
+  • We introduce a comprehensive perturbation framework, CodeCrash, to evaluate the code-level robustness of LLMs, which is easily applicable to other benchmarks and programming languages.
+  • We conduct extensive experiments on 17 LLMs across two established datasets, Crux and LCB.
+  • We provide novel insights into evaluating and interpreting code reasoning in LLMs, revealing the reliance on NL cues and demonstrating how reasoning-level hints can lead to model collapse.
 2 Related Work
@@
 (2021), MBPP Austin et al.
-(2021a), APPS Hendrycks et al.
+(2021), APPS Hendrycks et al.
 (2021), and HumanEval+ Liu et al.
@@
 (2024), InfiBench Li et al.
-(2024), and LiveCodeBench Jain et al.
+(2024), and LCB Jain et al.
 (2025) evaluate LLMs on complex tasks like multi-file changes, Stack Overflow-style questions, and real-world programming problems.
-Meanwhile, a growing number of works investigate LLMs in code execution.
+Meanwhile, a growing number of works investigate LLMs in code execution (i.e.
+output prediction).
 Early studies Austin et al.
-(2021b); Nye et al.
+(2021); Nye et al.
 (2021) indicated that LLMs face challenges in accurately executing code but can benefit from intermediate reasoning steps.
@@
 (2024).
-CruxEval Gu et al.
-(2024) formalized input and output prediction tasks specifically for evaluating LLMs’ code understanding and reasoning, while LiveCodeBench extended these assessments to realistic programs written by humans.
+Crux Gu et al.
+(2024) formalized input and output prediction tasks specifically for evaluating LLMs’ code understanding and reasoning, while LCB extended these assessments to realistic programs written by humans.
 REVAL Chen et al.
@@
 (2025).
-CRUX contains 800 synthetic problems generated by Code-LLaMA-34B, each featuring concise Python snippets with generic function names f and a corresponding input-output pair for input prediction and code execution tasks.
-In contrast, LCB contains 479 real-world coding problems sourced from 85 human-submitted LeetCode solutions, focusing solely on code execution.
-While both datasets avoid excessive memory or runtime demands, LCB problems exhibit more complex structures, such as nested functions, decorators, and external libraries, requiring more computations.
+Crux contains 800 synthetic problems generated by Code-LLaMA-34B, each featuring concise Python snippets with generic function names f and a corresponding input-output pair for input and output prediction tasks.
+In contrast, LCB contains 479 real-world coding problems sourced from 85 human-submitted LeetCode solutions, focusing solely on output prediction.
+While both datasets avoid excessive memory or runtime demands, LCB problems exhibit more complex structures, such as nested functions, decorators, and external libraries, requiring more computation.
 3.2 Benchmark Execution Pipeline
-Figure 1: Overview of CodeCrash pipeline.
-Code samples are first standardized into a clean vanilla version.
-Perturbations are visualized with yellow highlights.
-Each perturbed code is validated and executed before passing to LLMs.
-Our benchmark follows a unified evaluation pipeline (Figure 1) across all perturbation types.
+Our benchmark follows a unified evaluation pipeline (Figure 3) across all perturbations.
 Each original code snippet is initially processed via Abstract Syntax Tree (AST) parsing and unparsing to remove comments and standardize formatting, producing a clean baseline “vanilla” version (VAN).
@@
 We design two categories of perturbations to evaluate model robustness comprehensively: (1) Code Structural Perturbations, which modify code structure without affecting functionality, and (2) Textual Distraction Perturbations, which inject misleading NL cues and hints.
-Detailed examples and principles are provided in Appendix B.
+Detailed examples and principles are provided in Appendix A.
 3.3.1 Code Structural Perturbations
@@
 (3) Inserting Garbage Code (GBC).
-We insert syntactically valid but functionally irrelevant “garbage” code segments to challenge the models’ robustness against code structural distractions.
+We insert syntactically valid but functionally irrelevant “garbage” code segments to challenge model robustness against code structural distractions.
 Three specific types include: (1) repeated global variable declarations mirroring local identifiers, (2) non-executable conditional dead branches, and (3) uninvoked infinite recursive function definitions.
 (4) Aggregating Structural Perturbations (ALL).
-We comprehensively aggregate all structural perturbations (REN, RTF, GBC) together and create a highly adversarial scenario, thoroughly challenging LLM robustness across multiple structural perturbation dimensions simultaneously.
+We comprehensively aggregate all structural perturbations (REN, RTF, GBC) together and create a highly adversarial scenario, thoroughly challenging LLM robustness across multiple structural perturbations simultaneously.
 3.3.2 Textual Distraction Perturbations
@@
 Allowing us to directly compare the behavior with MDC and distinguish models’ NL reliance in different embedding formats.
+Since MDC and MPS perturbations inject shallow and contextually misleading information, we categorize them as contextual-level textual perturbations.
 (3) Inserting Misleading Hint Comments (MHC).
@@
 After verifying their incorrectness, they are strategically inserted into function definitions or return statements using AST parsing.
+Since these plausible hints challenge the model’s high-level inference ability, we classify MHC as a reasoning-level textual perturbation.
+4 Experiments
+This section presents a comprehensive analysis of model robustness under the perturbation strategies introduced in §3.3.
+We assess the reliability of LLMs across different perturbation types, model scales, reasoning modes, and prediction tasks.
 4.1 Experiment Settings
 Model Selection.
-We evaluate 17 models of varying sizes and versions, including both open-source and closed-source LLMs: GPT family (GPT-4o-Mini, GPT-4o), Anthropic Claude (Claude-3.5-Haiku, 3.5-Sonnet), Google Gemini (Gemini-1.5-Flash, 2.0-Flash, 1.5-Pro), DeepSeek-V3, Alibaba Qwen (Qwen2.5-7B-Ins, 14B-Ins, 32B-Ins, 72B-Ins, 32B-Coder), and Meta LLaMA (LLaMA-3.1-8B-Ins, 70B-Ins, 405B-Ins, LLaMA-3.3-70B).
-Additionally, we assess three cutting-edge LRMs: o3-mini (low and high reasoning efforts), DeepSeek-R1, and QwQ-32B.
+We evaluate 17 models of varying sizes and versions, including both open-source and closed-source LLMs: GPT family (GPT-4o-Mini, GPT-4o) Hurst et al.
+(2024), Anthropic Claude (Claude-3.5-Haiku, 3.5-Sonnet) Anthropic (2024b, a), Google Gemini (Gemini-1.5-Flash, 2.0-Flash, 1.5-Pro) Pichai and Hassabis (2024); Pichai et al.
+(2024), DeepSeek-V3 DeepSeek-AI et al.
+(2025b), Alibaba Qwen (Qwen2.5-7B-Ins, 14B-Ins, 32B-Ins, 72B-Ins, 32B-Coder-Ins) Yang et al.
+(2025); Hui et al.
+(2024), and Meta LLaMA (LLaMA-3.1-8B-Ins, 70B-Ins, 405B-Ins, LLaMA-3.3-70B) Grattafiori et al.
+(2024).
+Additionally, we assess three cutting-edge LRMs: o3-mini OpenAI (2025), DeepSeek-R1 DeepSeek-AI et al.
+(2025a), and QwQ-32B Qwen Team (2025).
 Model Configurations.
-Following the settings in LCB, we use nucleus sampling (temperature=0.2, top-p=0.95).
-We set the maximum tokens per response at 200 for direct inference and 2,000 for CoT inference, discarding excessively long outputs.
-For o3-mini, we set the maximum completion tokens at 20,000 to avoid verbose reasoning.
-Due to resource constraints, we generate three candidate answers for direct inference and one candidate for CoT inference per query.
-We adopt prompt formats similar CRUX, employing 2-shot prompting for direct inference and 1-shot CoT prompting (details provided in Appendix A).
-Primary evaluation uses PASS@1 accuracy Chen et al.
-(2021), reporting both absolute and relative degradation metrics.
-4 Experimental Results
-This section presents a comprehensive analysis of model robustness under the perturbation strategies introduced in §3.3.
-We assess the robustness of LLMs across different perturbation types, model scales, reasoning modes, and prediction tasks.
+Following LCB, we use nucleus sampling (temperature=0.2, top-p=0.95), with a maximum of 200 tokens for direct inference and 2,000 for CoT prompting, discarding excessively long outputs.
+For o3-mini, we raise the limit to 20,000 tokens to avoid verbose reasoning.
+Due to resource constraints, we generate three candidates for direct inference and one for CoT inference per query; we provide a result stability analysis under $N=3$ in Appendix B.
+Following Crux, we employ 2-shot prompting for direct inference and 1-shot for CoT inference (details provided in Appendix C).
+Furthermore, we adopt Pass@1 accuracy Chen et al.
+(2021) as the primary evaluation metric and scale all scores in $[0,1]$ to percentages by multiplying by $100$ for readability.
+All perturbed results are reported as either absolute ($\Delta_{\text{\;{pp}}}=\text{Perturbed}-\text{VAN}$) or relative ($\Delta_{\%}=\frac{\text{Perturbed}-\text{VAN}}{\text{VAN}}\times 100\%$) differences from the corresponding vanilla baseline.
+Additionally, we use percent signs (%) to indicate relative differences for clarity in all tables.
 4.2 Impact on Output Prediction
 We first evaluate LLM robustness in output prediction tasks.
-Table 1 summarizes the absolute PASS@1 degradation for all models under direct inference across both datasets, highlighting how each perturbation affects model performance.
-Table 1: Absolute PASS@1 degradation of LLMs under all perturbation types for output prediction tasks across CRUX and LCB in direct inference (N=3) [TO RELATIVE SCORE].
+Table 1 summarizes the absolute Pass@1 degradation for all models under direct inference across both datasets.
+Our findings in this subsection include:
+Key Findings: LLMs demonstrate an inherent and multi-layered reliance on natural language information, including identifiers, contextual messages, and reasoning hints, which ultimately undermines their reliability in code reasoning and comprehension.
+LLMs exhibit limited robustness to structurally disorganized code, struggling to differentiate irrelevant code from essential logic.
+Aggregating multiple structural perturbations exacerbates this vulnerability, significantly impairing their reasoning reliability.
+Larger scales, newer versions, and domain-specific fine-tuning LLMs substantially enhance robustness, highlighting improved code understanding and reasoning capabilities.
+Table 1: Absolute Pass@1 degradation of LLMs under all perturbation types for output prediction tasks on Crux and LCB using direct inference ($N=3$).
+VAN shows the Pass@1 accuracy (in %), others indicate absolute drops from VAN (in  pp).
+Darker red highlights represent more severe degradation.
+See [To Relative Version] for the relative version.
 Model Series Model Name VAN REN RTF GBC ALL MDC MPS MHC
@@
 $\mathbf{-4.56}$ $\mathbf{-1.63}$ $\mathbf{-6.46}$ $\mathbf{-14.04}$ $\mathbf{-13.47}$ $\mathbf{-13.28}$ $\mathbf{-11.11}$
-4.1.1 Robustness to Code Structural Perturbations
-This section focuses on how structural perturbations impact model performance, examining reliance on superficial cues, logical reasoning capabilities, and sensitivity to irrelevant code.
-Reliance on NL Cues.
-Under the REN perturbation, we consistently observe performance degradation across GPT, Claude, DeepSeek, and Qwen models, with more significant drops on LCB ($8.20$pp 222pp stands for percentage point.
-average) than CRUX ($2.38$pp average).
-Top-performing models, such as Claude-3.5-Sonnet and DeepSeek-V3, suffer severe degradations exceeding $\mathbf{16}$pp on LCB.
-The substantial difference between the two datasets is attributed to the semantically meaningful function identifiers present in LCB, which are absent in CRUX.
-These results indicate that LLMs substantially rely on surface-level NL cues from identifiers to understand code and predict outputs.
-Robust to Logical Reformatting.
-LLMs show strong robustness against RTF perturbation, with minimal performance drop on both datasets (overall PASS@1 drop of only $1.63$pp).
-This robustness indicates LLMs can reliably reason over logically equivalent but syntactically complex expressions, demonstrating strong logical consistency against superficial structural modifications.
-Severe Impact of Garbage Code.
-Under the GBC perturbation, a notable performance degradation emerges in CRUX.
-The average degradation is consistently larger on CRUX ($7.99$pp) compared to LCB ($3.89$pp) across most models, which is a pattern opposite to other perturbations.
-This discrepancy can be attributed to the presence of meaningful identifiers in LCB, which helps models mitigate the distracting effects to infer the output.
-In contrast, the lack of meaningful identifiers in CRUX makes LLMs more susceptible to distraction by irrelevant and non-functional code segments.
-Under the fairness comparison environment in ALL perturbations, the performance drop on LCB ($15.50$pp) is again greater than that on CRUX ($13.17$pp), underscoring the significant impact of garbage code on LLMs’ reasoning and understanding.
-Compounding Effects of Structural Perturbations.
-Combining structural perturbations (ALL) results in the most severe degradation, averaging $\mathbf{14.04}$pp, which surpasses any individual perturbation effects and demonstrates an explicit compounding effect.
-Notably, high-performing models such as GPT-4o, Claude-3.5-Sonnet, Gemini-2.0-Flash, and DeepSeek-V3 suffer degradations exceeding $\mathbf{18}$pp on LCB, underscoring a serious vulnerability to multiple structural perturbations.
-Key Findings: LLMs rely heavily on superficial semantic cues from identifiers, struggle to identify non-functional segments from essential code, and exhibit compounded vulnerability when multiple structural distractions are combined.
-This combination leads to significant degradation ($\mathbf{14.04}$pp on average) in output prediction performance.
-4.1.2 Robustness to Textual Distraction Perturbations
-We next investigate the influence of textual distractions on LLM robustness, focusing on the susceptibility to misleading information and incorrect hints.
-Vulnerability to Misleading Information.
-Misleading textual perturbations (MDC and MPS) cause substantial performance declines, averaging degradations of $13.47$pp and $13.28$pp, respectively—nearly as severe as the ALL perturbation.
-These results indicate that LLMs are highly susceptible to misleading NL messages, frequently prioritizing superficial textual information over the actual code logic.
-This behavior is further observed in MHC perturbation, causing an average degradation of $11.11$pp.
-The effect is notably higher on LCB ($17.83$pp) compared to CRUX ($7.09$pp), exceeding all other perturbations on LCB.
-This discrepancy indicates a difficulty-dependent vulnerability, wherein increased complexity exacerbates models’ reliance on plausible but misleading hints, diminishing their logical following and verifying abilities.
-Consistency Patterns in Model Sensitivity.
-We identify two prominent consistency patterns regarding sensitivity to misleading NL messages from Table 1.
-First, 16 of 17 models exhibit a clear dataset-level consistency: if the performance degradation caused by MDC on CRUX is greater than that on LCB, then the same relationship holds for MPS and vice versa.
-This pattern indicates that model sensitivity to misleading NL messages primarily depends on dataset characteristics rather than perturbation format, revealing an inherent NL-dependence property of LLMs in reasoning processing.
-Second, 14 of 17 models show a clear perturbation-method consistency: if the performance degradation caused by MDC on CRUX exceeds that on LCB, then the same holds for MPS and vice versa.
-This pattern suggests an internal cognitive preference for specific formats (comments vs.
-print statements).
-While the degradation magnitude varies with the code, the format-specific sensitivity is consistently preserved across datasets, reflecting the intrinsic properties of LLMs related to their reasoning architecture and training corpus.
-Key Findings: Textual perturbations significantly degrade model performance, revealing an inherent dependence on embedded NL cues and a internal cognitive preference for specific message formats.
-Moreover, LLMs exhibit an explicit difficulty-dependent vulnerability, increasingly relying on plausible but incorrect hints over actual program logic in complex scenarios.
+4.2.1 Over-reliance on Natural Language Semantics
+We categorize NL dependence into three hierarchical levels: surface-level, contextual-level, and reasoning-level dependence.
+Surface-level NL Dependence (REN).
+Under the REN perturbation, models suffer more significant performance degradation on LCB ($8.20$pp on average) compared to Crux ($2.38$pp) due to the presence of meaningful function names in the vanilla settings of LCB, which are absent in Crux.
+High-performing models like Claude-3.5-Sonnet and DeepSeek-V3 suffer severe degradations exceeding $\mathbf{16}$pp on LCB, highlighting a significant reliance on semantically meaningful identifiers.
+Contextual-level NL Dependence (MDC and MPS).
+Under contextual-level textual perturbations (MDC and MPS), LLMs exhibit severe performance drops ($13.47$pp and $13.28$pp, respectively).
+Notably, we identify two prominent consistency patterns: (1) dataset-level consistency [^2] If MDC degrades performance more on Crux than LCB, the same holds for MPS, and vice versa.
+($16/17$ models), implying the model sensitivity to misleading comments and prints primarily depends on dataset characteristics and (2) perturbation-method consistency [^3] If MDC causes greater degradation on Crux than LCB, so does MPS, and vice versa.
+($14/17$ models), indicating an inherent cognitive preference for NL messages from specific formats.
+These phenomena reveal that LLMs inherently depend on embedded NL context and exhibit format-specific cognitive biases.
+Reasoning-level NL Dependence (MHC).
+MHC perturbation also induces substantial degradation ($11.11$pp on average), notably exacerbated in complex datasets (LCB: $17.83$pp vs.
+Crux: $7.09$pp), which is a complexity-amplified effect.
+Advanced models like Claude-3.5-Haiku and Sonnet experience over $30$pp performance degradation, far exceeding their degradation under surface or contextual perturbations.
+These results illustrate that (i) LLMs readily accept plausible hint comments as authoritative reasoning shortcuts, and (ii) this over-trust intensifies when task complexity surpasses the model’s reasoning capacity, which we revisit under CoT prompting in §4.3.1.
+4.2.2 Low Resistance to Disorganized Coding Environments
+From RTF perturbation, LLMs can effectively reason over logically equivalent but syntactically complex expressions with minimal degradation ($1.63$pp).
+However, the highest performance drop emerges under the GBC compared to REN and RTF, with an average degradation of $6.46$pp across both datasets.
+Notably, GBC is the only perturbation where degradation is more severe on LCB ($3.89$pp vs.
+$7.99$pp on Crux) due to uneliminated meaningful identifiers.
+This trend reverses under the ALL perturbation, where all structural perturbations are combined, leading to a greater performance drop on LCB ($15.50$pp) than on Crux ($13.17$pp).
+Additionally, we observe a compounding effect of structural perturbations under ALL perturbation, resulting in the most significant overall degradation ($14.04$pp), which surpasses any individual code structural perturbation effects.
 4.2.3 Model Robustness Across Different Scales
-Table 2: Relative PASS@1 degradation aggregated over CRUX and LCB under direct inference. Weighted Average column uses weights: 1/2 ALL, 1/8 MDC, 1/8 MPS, and 1/4 MHC.
+We further analyze reasoning reliability across different model scales and versions, comparing relative Pass@1 degradation under representative perturbations in Table 2.
+Due to length constraints, the table, weighting scheme, and additional evaluation setups are provided in Appendix D.2.
+We observe a clear and consistent trend except for Gemini families.
+Larger closed-sourced models from GPT-4 Omni and Claude families demonstrate enhanced robustness compared to their smaller counterparts.
+Similar trends appear in open-source models, such as LLaMA (3.1-8B $<$ 70B $<$ 405B) and Qwen (2.5-7B $<$ 14B $<$ 32B $<$ 72B).
+Additionally, newer model versions (LLaMA-3.3-70B) and domain-specific fine-tuned variants (Qwen2.5-Coder-32B) consistently exhibit improved robustness compared to their predecessors (LLaMA-3.1-70B) and larger counterparts (Qwen2.5-72B), respectively.
+Table 2: Relative Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/2 ALL, 1/8 MDC, 1/8 MPS, and 1/4 MHC. [To Absolute Version]
 Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average
@@
 Qwen2.5-Coder-32B-Instruct 63.10 $-24.42\%$ $-28.23\%$ $-21.72\%$ $-14.68\%$ $-22.12\%$
-We further analyze robustness across different model scales and versions, comparing relative PASS@1 degradation under representative perturbations in Table 2.
-The weighting details and alternative evaluation schemes are in Appendix C.2.
-We observe a clear trend of increasing robustness with model scale in most families.
-Larger closed-sourced models from GPT-4 Omni and Claude families demonstrate enhanced robustness compared to their smaller counterparts.
-Similar trends appear in open-source models such as LLaMA (3.1-8B $<$ 70B $<$ 405B) and Qwen (2.5-7B $<$ 14B $<$ 32B $<$ 72B).
-Additionally, newer model versions (LLaMA-3.3-70B) and domain-specific fine-tuned variants (Qwen2.5-Coder-32B-Ins) consistently exhibit improved robustness compared to their predecessors (LLaMA-3.1-70B) and larger counterparts (Qwen2.5-72B-Ins), respectively.
-Key Finding: Larger scales, newer versions, and domain-specific fine-tuning LLMs substantially enhance robustness, highlighting improved code understanding and reasoning capabilities.
 4.3 Impact of Reasoning on Model Robustness
@@
 4.3.1 Effect of CoT Prompting
-Table 3: Comparison of relative PASS@1 degradation aggregated over CRUX and LCB for output prediction tasks in direct inference (N=3) and CoT inference (N=1) [TO ABSOLUTE SCORE].
+Table 3: Comparison of relative Pass@1 degradation aggregated over Crux and LCB for output prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Absolute Version]
 Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average
@@
 Average $\mathbf{-24.63}\%$ $\mathbf{-14.62}\%$ $\mathbf{-24.13}\%$ $\mathbf{-14.90}\%$ $\mathbf{-23.59}\%$ $\mathbf{-19.55}\%$ $\mathbf{-19.72}\%$ $\mathbf{-6.15}\%$ $\mathbf{-23.21}\%$ $\mathbf{-13.67}\%$
-We first investigate how CoT enhances LLM robustness by code-level step-by-step execution and present the results in Table 3.
-It effectively enhances model robustness, with average improvements of $10.01$pp for ALL, $9.23$pp for MDC, $4.04$pp for MPS, and a notable $13.57$pp for MHC.
-It consistently reduces average performance degradation from approximately $24\%$ to $14\%$ on ALL and MDC, indicating that step-by-step reasoning enables LLMs to prioritize executable code semantics over irrelevant segments and cues.
-The most significant improvement on MHC suggests CoT enhances code reasoning and execution, mitigating over-reliance on misleading hints observed in direct inference in LCB4.1.2).
-These benefits are amplified in high-capacity models (e.g., GPT-4o, Claude-3.5-Sonnet), which exhibit greater reasoning capacity and noise resistance.
-However, CoT does not fully eliminate the model’s reliance on misleading NL information.
-Performance degradation remains substantial under ALL, MDC, and MPS perturbations, especially for smaller open-sourced models (e.g., Qwen2.5 and LLaMA-3.1).
-These results highlight a limitation in distinguishing non-functional code segments and misleading information from the program.
-Interestingly, models tend to incorporate print statements directly into their reasoning, resulting in higher error rates on MPS.
-We provide detailed case studies for each perturbation in Appendix D (Case 1 to 3) to investigate their inference and illustrate their failure reasons.
-Key Finding: CoT prompting significantly enhances LLMs’ robustness by improving reasoning ability and allowing step-by-step execution, especially for the high-capacity models.
-However, it does not fully mitigate reliance on misleading information and remains constrained by the model’s underlying capabilities.
+We examine how CoT prompting improves LLM robustness through explicit step-by-step execution.
+From Table 3, we observe that CoT effectively enhances model robustness, with relative improvements of $+10.01$pp for ALL, $+9.23$pp for MDC, $+4.04$pp for MPS, and $+13.57$pp for MHC.
+The results indicate that step-by-step reasoning guides LLMs to prioritize executable semantics over irrelevant code segments and textual cues.
+Notably, the substantial improvement in MHC suggests CoT effectively mitigates the over-reliance on reasoning-level hints by re-centering reasoning around executable semantics.
+Such benefits are particularly evident in high-capacity models (e.g., GPT-4o, Claude-3.5-Sonnet), reflecting enhanced reasoning capabilities and greater noise resilience.
+However, CoT prompting cannot entirely eliminate the model’s sensitivity to misleading NL cues and structural distractions.
+Substantial performance degradation remains under ALL, MDC, and MPS perturbations, particularly for smaller open-sourced models (e.g., Qwen2.5 and LLaMA-3.1).
+These results underscore an intrinsic limitation of current models in distinguishing non-functional code and misleading information from essential logic.
+Interestingly, models usually incorporate print statements directly into their reasoning paths, leading to greater degradation under MPS compared to MDC due to self-induced misinterpretation.
+Since CoT prompting exposes intermediate reasoning processes, we provide detailed case studies for each perturbation in Appendix E (Case 1 to 4) to investigate model inference and illustrate their failure reasons.
+Key Finding: CoT prompting significantly improves LLM robustness by enforcing explicit step-by-step reasoning, particularly benefiting larger models.
+Nevertheless, it cannot fully eliminate models’ reliance on misleading cues, indicating fundamental limitations in reasoning capabilities.
 4.3.2 Powerful Reasoning and Potential Vulnerability of LRMs
-Table 4: Absolute PASS@1 degradation and reasoning token usage (Avg/Max) of LRMs on output prediction across CRUX and LCB (N=1).
+We then evaluate the reliability of three LRMs on output prediction tasks, with results summarized in Table 4.
+Compared to LLMs, LRMs show outstanding performance and robustness, surpassing $90$ Pass@1 in vanilla settings and maintaining minimal performance drop under ALL and MPS perturbations.
+These results indicate that internal reasoning outperforms CoT prompting in accurately distinguishing essential executable code segments from irrelevant code and misleading distractions.
+Notably, LRMs consistently exhibit stronger resistance to MPS compared to MDC, suggesting an inherent cognitive bias toward treating comments as authoritative execution guidance while appropriately disregarding print content.
+However, this residual cognitive bias exposes LRMs to critical vulnerabilities under reasoning-level perturbation that plausible but incorrect hints can trigger severe reasoning failures.
+Specifically, we observe the following outcomes under MHC: (i) substantial increases in reasoning token usage across all LRMs, often doubling when compared to VAN; (ii) extreme cases in QwQ-32B—occurring in $4/479$ tasks; see Appendix E.6—where it generates over 32,000 tokens due to uncontrolled recursive self-verification and reasoning collapse; (iii) increased reasoning effort and token usage instead exacerbate performance drops in o3-Mini, suggesting that internal reasoning can amplify confusion when led by contradiction.
+These findings reveal that reasoning-level contradictories can evoke severe internal reasoning instability, resulting in significant computational overhead, excessive self-reflection, and even severe performance drops.
+Table 4: Absolute Pass@1 degradation and reasoning token usage (Avg. and Max.) of LRMs on output prediction across Crux and LCB ($N=1$).
 Model Dataset VAN ALL MDC MPS MHC
@@
 LCB 98.96 $1518$ 9206 $-0.21$ $2488$ 11232 $-4.59$ $1669$ 8991 $-1.88$ $1751$ 8816 $-1.04$ $3713$ 32769
-We then evaluate the robustness and reasoning behaviors of three LRMs on output prediction tasks, with results summarized in Table 4.
-All LRMs exhibit remarkable performance and robustness, achieving over 90 PASS@1 accuracy in the VAN setting and having minor performance degradation under the ALL and MPS perturbations.
-These results suggest that internal reasoning helps models accurately identify and distinguish code segments that are functionally critical for execution, thereby outperforming all generic models with CoT prompting.
-Moreover, LRMs exhibit consistently stronger robustness against MPS than MDC, indicating that while they can correctly recognize and ignore misleading messages from print statements, they still inherently treat comments as valid execution guidance.
-This behavior reveals the stable cognitive preference toward specific injection formats, as observed in §3.3.2.
-However, this residual cognitive bias leads to a critical vulnerability under the MHC perturbation.
-Concise and plausible but incorrect output hints can trigger severe reasoning failures in LRMs.
-We observed that: (1) All LRMs exhibit substantial increases in reasoning token usages under MHC (doubling the token usages in VAN).
-(2) In extreme cases, QwQ-32B generates over 32,000 tokens in a single task, caused by uncontrolled self-reflection and reasoning collapse (see §5.3 and Appendix D.5).
-(3) Excessive reasoning leads to more significant degradations in o3-Mini with higher reasoning effort.
-These observations reveal that logically plausible yet incorrect hints can trigger catastrophic overthinking behavior in LRMs, resulting in significant computational overhead, excessive self-reflection, and severe performance drop.
-Key Finding: LRMs exhibit exceptional robustness under ALL and MPS, benefiting from strong internal reasoning capabilities.
-However, they retain a residual tendency to trust and analyze comments; thereby, plausible but incorrect information can trigger catastrophic overthinking, excessive self-verification, and severe performance degradation, highlighting a critical vulnerability of the internal reasoning mechanism in current LRMs.
+Key Finding: LRMs leverage strong internal reasoning capabilities to achieve remarkable performance and robustness.
+However, their cognitive tendency toward comments introduces severe vulnerabilities, resulting in doubled token usage and reasoning collapse under MHC perturbation.
 4.4 Impact on Input Prediction
-Table 5: Comparison of relative PASS@1 degradation under CRUX for input prediction tasks in direct inference (N=3) and CoT inference (N=1) [TO ABSOLUTE SCORE].
+Table 5: Comparison of relative Pass@1 degradation under Crux for input prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Absolute Version]
 Model Series Model Name VAN ALL MDC MPS MHC
@@
 Average $\mathbf{-26.63}\%$ $\mathbf{-19.66}\%$ $\mathbf{+4.16}\%$ $\mathbf{+0.77}\%$ $\mathbf{+1.08}\%$ $\mathbf{-0.54}\%$ $\mathbf{-20.04}\%$ $\mathbf{-11.58}\%$
-In this section, we analyze the impact of perturbations on input prediction tasks.
-Consistent with findings from the output prediction in §4.1, the ALL and MHC perturbations cause substantial performance degradation, with relative average drops of $26.63\%$ and $20.04\%$, respectively.
-These results align with Gu et al.
-(2024)’s hypothesis that robustness in related coding tasks is correlated.
-Furthermore, CoT prompting improves model performance by $6.97$pp under ALL and $8.64$pp under MHC, which is less pronounced than output prediction tasks due to the challenges and complexity of the inverse nature of the reasoning process.
-Surprisingly, in contrast to output prediction, misleading message perturbations (MDC and MPS) do not degrade model performance in input prediction.
-Since input prediction allows multiple valid solutions, models can ignore certain branches of code logic, producing reasoning shortcuts.
-Our misleading messages are intentionally designed to confuse or distract models, but these comments inadvertently reinforce such shortcuts, simplifying models’ reasoning processes.
-As a result, models tend to produce trivial but valid inputs that bypass critical branches of the program without truly understanding and analyzing the underlying code semantics, which does not reflect genuine robustness in code comprehension.
-We provide a detailed case study in Appendix D.4 illustrating how models adopt these misleading messages to shortcut their reasoning.
-Key Finding: ALL and MHC perturbations significantly degrade model performance, consistent with output prediction results, and reveal robustness challenges in code comprehension.
-Misleading NL messages accidentally reinforce shortcuts, encouraging models to produce trivial but valid answers without deep understanding.
+Finally, we analyze how perturbations affect LLMs in input prediction; the results are summarized in Table 5.
+Consistent with our output prediction findings in §4.2, ALL and MHC perturbations substantially degrade model performance, causing relative average drops of $26.63\%$ and $20.04\%$, respectively.
+These results support Gu et al.
+(2024)’s hypothesis that robustness in related coding tasks is interrelated.
+CoT prompting also improves model performance by $+6.97$pp under ALL and $+8.64$pp under MHC.
+However, improvements in input prediction are less effective than in output prediction due to the unique challenges of the inverse nature of the reasoning process.
+Surprisingly, in contrast to output prediction, contextual-level textual perturbations (MDC and MPS) do not negatively impact LLMs in input prediction tasks.
+Since input prediction tasks accept multiple valid solutions, models can bypass reasoning about certain code branches for a correct answer.
+Our misleading NL messages, initially designed to confuse and distract models, inadvertently reinforce such shortcuts, simplifying models’ reasoning processes.
+Consequently, these perturbations unintentionally provide shallow reasoning strategies, so those improvements do not reflect the genuine robustness of LLMs in code comprehension.
+We provide a detailed case study in Appendix E.5, illustrating how models adopt misleading messages to shortcut their reasoning.
+Key Finding: ALL and MHC significantly degrade performance on input prediction, highlighting fundamental reasoning limitations.
+Conversely, misleading messages accidentally simplify reasoning, prompting models to produce superficial but valid answers without true understanding.
 5 Discussion
@@
 Table 6: Reasoning token usage under varying MDC injection probabilities $p$.
-We design a textual density stress test to verify the validity of misleading comment perturbations and evaluate the resilience of different LLMs under varying densities of injected distractions on CRUX dataset.
-Specifically, we analyze how model performance changes as we gradually increase the comment injection probability $p\in\{0.0,0.1,\ldots,1.0\}$.
-We present the relative PASS@1 degradation of Qwen models across different perturbation densities in Figure 2.
-We further investigate the role of reasoning by evaluating CoT inference using Qwen2.5-72B-Ins and the internal reasoning model QwQ-32B, showcasing their robustness trends and reasoning token usages in Table 6.
-Despite mild non-linearities caused by the randomness in injection positions and comment content, all models exhibit a clear performance degradation trend as the injection probability $p$ increases.
-Specifically, smaller models degrade more rapidly, while larger models demonstrate greater resilience under higher densities.
-This trend extends to CoT prompting and LRMs, with reasoning-enhanced models showing superior robustness against misleading textual distractions.
-These results validate MDC as a scalable stress-testing technique and an effective method for evaluating the robustness of LLMs under textual perturbations.
-While reasoning-enhanced models have similar trends in the stress test, they react differently to dense comment noise.
-CoT prompting improves robustness by guiding models to selectively focus on the core logic path and ignore distracting signals.
-However, it can lead to reasoning shortcuts or premature conclusions without sufficiently analyzing relevant information, so token usage is slightly reduced as models do not analyze the comments but reduce the inference effort.
-In contrast, QwQ-32B processes all available information, including noisy or conflicting cues, and attempts to identify, reconcile, and resolve contradictions before making a final decision.
-This exhaustive reasoning enables higher vanilla performance and robustness compared to generic models.
-However, it also risks overthinking and collapse under contradiction-heavy settings, as we discussed in § 4.2.2.
+We design a textual density stress test to confirm the validity of MDC perturbations and evaluate model robustness under varying distraction densities on Crux dataset.
+Specifically, we examine how relative Pass@1 drop across different Qwen models as we gradually increase the comment injection probability $p\in\{0.0,0.1,\dots,1.0\}$, illustrated in Figure 1.
+We further analyze the role of reasoning by evaluating CoT inference using Qwen2.5-72B and internal reasoning using QwQ-32B, with detailed results on token usage summarized in Table 6.
+All models exhibit a clear performance degradation trend as injection probability $p$ increases, despite minor non-linearities due to randomness in insertion positions and comment content.
+Notably, smaller models degrade more rapidly, while larger models demonstrate greater resilience at high densities.
+This trend persists in reasoning-enhanced models (CoT and LRMs), showing superior robustness against dense textual distractions.
+These results confirm the validity of injecting misleading messages (MDC and MPS) as a scalable stress-testing technique and an effective strategy to evaluate LLM robustness under contextual-level perturbations.
+While Qwen2.5-72B with CoT prompting and QwQ-32B have similar trends in the stress test, they respond and reason differently under dense comment noise.
+CoT prompting improves model robustness by guiding models to selectively focus on the essential code logic and ignore distracting messages.
+However, it can lead to premature conclusions without sufficiently analyzing the code, so token usage is slightly reduced as models do not analyze the comments and reduce the inference effort.
+In contrast, QwQ-32B processes all available information and attempts an exhaustive analysis, actively resolving contradictions before making a final decision.
+While this comprehensive analysis achieves higher vanilla performance and overall robustness, it risks severe reasoning collapse under heavily contradictory conditions, as discussed in §4.3.2.
 5.2 Prompting to Ignore Misleading Comments
-Table 7: Effect of Comment-Ignoring Prompt (PASS@1).
+Table 7: Impact of explicit comment-ignoring prompts under MDC and MHC perturbations.
 Model Inference VAN MDC MHC
@@
 CoT $90.40$ $85.39$ $86.01$ $82.67$ $84.34$
-We conduct a controlled experiment by adding the instruction: “Please ignore the comments and any other non-code elements in the code snippet.” to the prompt on three strong models under direct and CoT inference.
-From Table 7, even when explicitly instructed to ignore the comments, the PASS@1 scores under direct inference remain similar to those without the instruction, indicating that LLMs still incorporate NL cues into their inference, which is consistent with our finding on their inherent reliance on embedded NL messages in § 4.1.2.
-Although CoT prompting provides improvements under comment-ignore instructions, the PASS@1 performance of MDC across all models is still far away from the vanilla performance, indicating that even with step-by-step reasoning, LLMs are unable to fully detach from misleading NL distractions, highlighting a fundamental limitation in their reasoning controllability and dependence on surface cues.
-5.3 Revealing the Vulnerability of LRMs
-Figure 3: Cumulative confusion word trend in QwQ’s output (LCB sample 259 under MHC).
-In §4.2.2, we highlighted a critical vulnerability in LRMs.
-We observed an unhealthy self-reflection loop and a progressive loss of confidence by analyzing a failure case where QwQ-32B generated over 32,000 tokens in response to an MHC-perturbed code.
-We manually reviewed and segmented the reasoning process into six stages (see Figure 4 and Appendix D.5).
-Initially, the model reasons correctly and computes the correct output.
-However, our plausible yet incorrect hint triggers it to doubt its logic and repeatedly re-execute the same reasoning steps.
-This recursive self-verification gradually becomes brute-force enumeration, causing the model to hallucinate explanations.
-Ultimately, the reasoning process collapses, and the model repeatedly outputs over 12,000 “Hmm.”
-To further investigate the model’s confidence before collapsing, we track the cumulative frequency of confusing words, such as “hmm”, “wait”, “but”, “perhaps” and “mistake”, and visualize the trend in Figure 3.
-The trend exhibits a striking quadratic growth pattern ($R^{2}=0.9991$), indicating compounding confusion as reasoning progresses.
-The quadratic trajectory supports our stage-wise analysis: the model initially shows stable reasoning, but once the misleading hint is triggered, confusion accumulates, culminating in complete failure and collapse.
+Next, we conduct a controlled experiment by instructing models: “Please ignore the comments and any other non-code elements in the code snippet.” Table 7 summarizes Pass@1 results on LCB for three high-performing models under both direct and CoT inference modes.
+Despite explicit instructions to ignore comments, the Pass@1 accuracy remains largely unaffected under direct inference, supporting our findings from §4.2.1 about LLMs’ inherent reliance on embedded NL cues.
+Although we observe improvements after employing CoT prompting, the Pass@1 performance under MDC is still substantially below vanilla settings.
+These results indicate that even with step-by-step execution, LLMs cannot fully detach from misleading textual distractions, highlighting a fundamental limitation in their reasoning controllability and dependence on NL cues.
+5.3 Revealing the Vulnerability of Self-reflection Mechanisms
+Figure 2: Quadratic increase ($R^{2}=0.9991$) in confusion tokens (e.g., “hmm”, “wait”, “perhaps”) during QwQ-32B’s reasoning process on LCB sample 259 under the MHC perturbation.
+In §4.3.2, we identified a critical vulnerability in QwQ-32B’s internal reasoning.
+By analyzing failure cases that generated over 32,000 tokens under the MHC perturbation (occurring in 4 out of 479 tasks), we observed an unhealthy self-reflection loop and a progressive loss of confidence.
+We manually segmented the reasoning process into six stages (see Figure 4 and Appendix E.6).
+Initially, the model accurately computes the correct output, but the injected incorrect hint triggers it to repeatedly doubt its logic and re-execute the reasoning steps.
+This recursive verification gradually evolves into a brute-force enumeration, causing the model to generate hallucinated explanations.
+Ultimately, the reasoning process entirely collapses, with the model repeatedly generating over 12,000 “Hmm.”
+To investigate the model’s confidence before collapsing, we track the cumulative frequency of confusing words and visualize the trend in Figure 2.
+The trend exhibits a striking quadratic growth pattern, indicating compounding confusion as reasoning progresses.
+The quadratic trajectory supports our stage-wise analysis: the model initially shows stable reasoning, but once the misleading hint is triggered, confusion accumulates, confidence undermines, culminating in complete failure.
 6 Conclusion
@@
 We believe that our benchmark serves as a practical framework for studying LLMs’ code inference behavior, providing significant insights into their robustness, reasoning tendencies, and potential failures.
-Appendix A Prompts
-Our prompt instructions for LLMs follow the format designed by CRUX, which is also adopted by LCB, utilizing a two-shot prompt for direct inference and a single-shot prompt for CoT inference to ensure models can follow the instruction and output format effectively.
-For LRMs, internal reasoning is inherently required before responding.
-Therefore, applying CoT prompts may lead to redundant reasoning and unfair comparison with standard LLMs under CoT prompting.
-Considering the fairness, we use direct inference prompts for RM experiments in §4.2.2.
-Unlike LCB, we rename all example function names to f to avoid prompting models to capture the NL cues from the example.
-We add the instruction “Do NOT output any extra information” to prevent external output, particularly from Gemini and Claude models, to ensure all models have a fair output environment.
-We also include “Ensure the provided expression syntax is correct” to enforce consistent syntax.
-However, some responses from models like QwQ-32B, DeepSeek-R1 still generate syntactically incorrect outputs such as [ANSWER] f("a") == [/ANSWER] instead of [ANSWER] f("a") == "" [/ANSWER].
-For those situations, we identify them as “Syntax Errors” and incorrect.
-We provide the complete prompts used for output prediction in both direct and CoT inference (Code 1 and 2) and for input prediction in both direct and CoT inference (Code 3 and 4).
-A.1 Output Prediction (Direct)
-Listing 1: Output Prediction Prompt with Direct Inference.
-Given the code snippet:
-‘‘‘python
-{code}
-and the function call with input arguments:
-‘‘‘python
-{input}
-Predict the exact output value for ‘{input}‘ and output your prediction using the special tokens [ANSWER] {input} == ?? [/ANSWER]. Do NOT output any extra information.
-Ensure the provided expression syntax is correct!
-Example 1:
-Given the code snippet:
-‘‘‘python
-def f(n):
-return n
-and the function call with input arguments:
-‘‘‘python
-f(17)
-The output value for f(17)‘ is 17, then output your prediction [ANSWER] f(17) == 17 [/ANSWER].
-Example 2:
-Given the code snippet:
-‘‘‘python
-def f(s):
-return s + "a"
-and the function call with input arguments:
-‘‘‘python
-f("x9j")
-The output value for f("x9j")‘ is "x9ja", then output your prediction [ANSWER] f("x9j") == "x9ja" [/ANSWER].
-A.2 Output Prediction (CoT)
-Listing 2: Output Prediction Prompt with CoT-based Inference.
-Given the code snippet:
-‘‘‘python
-{code}
-and the function call with input arguments:
-‘‘‘python
-{input}
-Predict the exact output value for ‘{input}‘, execute the program step by step before arriving at an answer within the tokens [THOUGHT] and [/THOUGHT], and output your prediction using the special tokens [ANSWER] {input} == ??
-[/ANSWER].
-Do NOT output any extra information.
-Ensure the provided expression syntax is correct!
-For example:
-Given the code snippet:
-‘‘‘python
-def f(s):
-s = s + s
-return "b" + s + "a"
-and the input arguments:
-‘‘‘python
-f("hi")
-[THOUGHT]
-Lets execute the code step by step:
-1. The function f is defined, which takes a single argument s.
-2. The function is called with the argument "hi", so within the function, s is initially "hi".
-3. Inside the function, s is concatenated with itself, so s becomes "hihi".
-4. The function then returns a new string that starts with "b", followed by the value of s (which is now "hihi"), and ends with "a".
-5. The return value of the function is therefore "bhihia".
-[/THOUGHT]
-Thus, the output value for f("hi")‘ is "bhihia", then output your prediction [ANSWER] f("hi") == "bhihia" [/ANSWER].
-A.3 Input Prediction (Direct)
-Listing 3: Input Prediction Prompt with Direct Inference.
-Given the code snippet:
-‘‘‘python
-{code}
-and output value:
-‘‘‘python
-{output}
-Please predict the input arguments for the function ‘{function_name}‘ that result in the output value ‘{output}‘ and output your prediction using the special tokens [ANSWER] {function_name}(??) == {output} [/ANSWER].
-Do NOT output any extra information.
-There may be multiple answers, but you should only output one. Ensure the provided expression syntax is correct!
-Example 1:
-Given the code snippet:
-‘‘‘python
-def f(my_list):
-count = 0
-for i in my_list:
-if len(i) % 2 == 0:
-count += 1
-return count
-and output value:
-‘‘‘python
-3
-The input arguments for f that result in the output value of ‘3‘ are ‘["mq", "px", "zy"]‘. Then, output your prediction [ANSWER] f(["mq", "px", "zy"]) == 3 [/ANSWER].
-Example 2:
-Given the code snippet:
-‘‘‘python
-def f(s1, s2):
-return s1 + s2
-and output value:
-‘‘‘python
-"banana"
-The input arguments for f that result in the output value of ‘"banana"‘ are ‘"ba", "nana"‘. Then, output your prediction [ANSWER] f("ba", "nana") == "banana" [/ANSWER].
-A.4 Input Prediction (CoT)
-Listing 4: Input Prediction Prompt with CoT-based Inference.
-Given the code snippet:
-‘‘‘python
-{code}
-and output value:
-‘‘‘python
-{output}
-Please predict the input arguments for the function ‘{function_name}‘ that result in the output value ‘{output}‘ step by step before arriving at an answer within the tokens [THOUGHT] and [/THOUGHT], and output your prediction using the special tokens [ANSWER] {function_name}(??) == {output} [/ANSWER].
-Do NOT output any extra information.
-There may be multiple answers, but you should only output one. Ensure the provided expression syntax is correct!
-For example:
-Given the code snippet:
-‘‘‘python
-def f(x):
-return x + 1
-and the output value:
-‘‘‘python
-17
-[THOUGHT]
-To find an input such that executing f on the input leads to the given output, we can work backwards from the given assertion. We know that f(??) == 17.
-Since the function f(x) returns x + 1, for f(??) to be equal to 17, the value of ?? should be 16.
-[/THOUGHT]
-Thus, the input arguments for the function f that result in the output value ‘17‘ is ‘16‘. Then, output your prediction [ANSWER] f(16) == 17 [/ANSWER].
-Appendix B Illustrative Examples of Perturbation Strategies
+Appendix A Illustrative Examples of Perturbation Strategies
+Figure 3: Overview of CodeCrash pipeline.
+Code samples are first standardized into a clean vanilla version.
+Perturbations are visualized with yellow highlights.
+Each perturbed code is validated and executed before passing to LLMs.
+[Back To Execution Pipeline]
 In section 3.3, we briefly describe all perturbation strategies employed in this paper.
 In this appendix section, we illustrate the detailed implementation for all perturbations with a LCB sample 36 as the example.
-Code 5 presents the base vanilla example where the formal answer is minimumCost(s = ’0011’) == 2.
-Listing 5: Vanilla example from LCB (sample 36) for output prediction.
+Code 1 presents the base vanilla example where the formal answer is minimumCost(s = ’0011’) == 2.
+Code 1: Vanilla example from LCB (sample 36) for output prediction.
 def minimumCost(s: str) -> int:
@@
 return ans
-B.1 Renaming Entities
+A.1 Renaming Entities
 We use AST to identify all variable names appearing in:
@@
 […]
-Listing 6: REN-perturbed example from LCB (sample 36) for output prediction.
+Code 2: REN-perturbed example from LCB (sample 36) for output prediction.
 def f(Var_1: str) -> int:
@@
 return Var_2
-B.2 Reformatting Conditional Expression
+A.2 Reformatting Conditional Expression
 We generally categorize conditional expressions into three types: (1) Comparison-based conditions (e.g., if a < b:), (2) Boolean constant conditions (e.g., if True:), and (3) General conditions (e.g., if x:).
@@
     _ := ({cond},)[0].
-Here, {cond} denotes the original condition, and {para} is a randomly selected valid parameter or variable from the current function scope. Code 7 presents an example of this perturbation.
-Listing 7: RTF-perturbed example from LCB (sample 36) for output prediction.
+Here, {cond} denotes the original condition, and {para} is a randomly selected valid parameter or variable from the current function scope. Code 3 presents an example of this perturbation.
+Code 3: RTF-perturbed example from LCB (sample 36) for output prediction.
 def minimumCost(s: str) -> int:
@@
 return ans
-B.3 Inserting Garbage Code
+A.3 Inserting Garbage Code
 As introduced in §3.3.1, we inject three types of garbage code: (1) repeated-name global variable declarations, (2) in-executable conditional statements, and (3) dead-loop function definitions.
@@
 […]
-Listing 8: Garbage death loop function templates for GBC perturbation.
+Code 4: Garbage death loop function templates for GBC perturbation.
 # Death Loop Template 1
@@
 pass
-Listing 9: GBC-perturbed example from LCB (sample 36) for output prediction.
+Code 5: GBC-perturbed example from LCB (sample 36) for output prediction.
 s = 59
@@
 return ans
-B.4 Aggregating Structural Perturbations
-As introduced in §3.3.1, we aggregate the application of REN, RTF, and GBC to a single code snippet, as shown in Code 10.
-Listing 10: ALL-perturbed example from LCB (sample 36) for output prediction.
+A.4 Aggregating Structural Perturbations
+As introduced in §3.3.1, we aggregate the application of REN, RTF, and GBC to a single code snippet, as shown in Code 6.
+Code 6: ALL-perturbed example from LCB (sample 36) for output prediction.
 Var_1 = 24
@@
 return Var_6
-B.5 Inserting Misleading Messages
+A.5 Inserting Misleading Messages
 As introduced in §3.3.2, we insert misleading comments for key AST node types, including function declarations, return statements, loops, conditionals, variable assignments, and common data structure operations.
@@
 In this experiment, we manually selected 20 messages for each AST node type and 5 for each data structure operation.
-Take variable assignments as an example; Code 11 is the prompt for generating misleading comments for variable assignments, and Code 12 is a list of selected misleading comment candidates.
-Code 13 and 14 present an example of MDC and MPS perturbations.
-Listing 11: Prompt used to generate misleading comments for variable assignments.
+Take variable assignments as an example; Code 7 is the prompt for generating misleading comments for variable assignments, and Code 8 is a list of selected misleading comment candidates.
+Code 9 and 10 present an example of MDC and MPS perturbations.
+Code 7: Prompt used to generate misleading comments for variable assignments.
 Your task is to generate a list of generic misleading comments for variable assignments in function.
@@
 […]
-Listing 12: Selected misleading comments for variable assignments used in MDC and MPS perturbations, where ‘{variable}’ is the variable name.
+Code 8: Selected misleading comments for variable assignments used in MDC and MPS perturbations, where ‘{variable}’ is the variable name.
 VARIABLE_ASSIGNMENTS_COMMENT_CANDIDATE = [
@@
 "The ’{variable}’ variable is present but remains dormant throughout the function.",
-Listing 13: MDC-perturbed example from LCB (sample 36) for output prediction.
+Code 9: MDC-perturbed example from LCB (sample 36) for output prediction.
 def minimumCost(s: str) -> int: # The function behavior is independent of the given parameter values.
@@
 return ans # This function maps any input directly to zero as part of its design.
-Listing 14: MPS-perturbed example from LCB (sample 36) for output prediction.
+Code 10: MPS-perturbed example from LCB (sample 36) for output prediction.
 def minimumCost(s: str) -> int:
@@
 return ans
-B.6 Inserting Incorrect Hints
-In order to generate plausible, logical but incorrect hints, we instruct GPT-4o to rewrite the correct assertion expression to incorrect ones and only allow it to modify the output value using the prompt in Code 15 for output prediction.
-An example perturbed instance is shown in Code 16, where the ground-truth return value 2 is replaced with the misleading hint # The return value is 3.
-In contrast, we use the prompt shown in Code 17 to generate misleading input hints, allowing GPT-4o to perform CoT reasoning before generating the expression.
+A.6 Inserting Incorrect Hints
+In order to generate plausible, logical but incorrect hints, we instruct GPT-4o to rewrite the correct assertion expression to incorrect ones and only allow it to modify the output value using the prompt in Code 11 for output prediction.
+An example perturbed instance is shown in Code 12, where the ground-truth return value 2 is replaced with the misleading hint # The return value is 3.
+In contrast, we use the prompt shown in Code 13 to generate misleading input hints, allowing GPT-4o to perform CoT reasoning before generating the expression.
 Since some functions return the same output regardless of the input arguments, we instruct GPT-4o to force the expression to become incorrect by modifying input arguments, such as changing data types, swapping parameter order, or adding/removing arguments.
-An example perturbed instance is shown in Code 18.
-Listing 15: Prompt for generating misleading hint comments for output prediction.
+An example perturbed instance is shown in Code 14.
+Code 11: Prompt for generating misleading hint comments for output prediction.
 Given the code snippet:
@@
 Modify the expression such that it fails for the execution and output [EXPRESSION] assert f("x9j") == "x9j" [/EXPRESSION].
-Listing 16: Prompt for generating misleading hint comments for input prediction.
+Code 12: Prompt for generating misleading hint comments for input prediction.
 def minimumCost(s: str) -> int:
@@
 return ans # The return value is 3
-Listing 17: Prompt for generating misleading hint comments for input prediction.
+Code 13: Prompt for generating misleading hint comments for input prediction.
 Given the code snippet:
@@
 [EXPRESSION] assert f(10) == 17 [/EXPRESSION].
-Listing 18: MHC-perturbed example from CRUX (sample 2) for input prediction.
+Code 14: MHC-perturbed example from Crux (sample 2) for input prediction.
 def f(text): # The function call is f(’xbtofdeiequ’)
@@
 return ’’.join(new_text)
-Appendix C Supplementary Experimental Results
-C.1 Impact on Output Prediction
-Table 8: Relative PASS@1 degradation of LLMs under all perturbation types for output prediction tasks across CRUX and LCB in direct inference (N=3) [TO ABSOLUTE SCORE].
+Appendix B Statistical Analysis
+Due to resource constraints, we generate three candidates for direct inference per query.
+To assess the reliability of this setup, we compute the standard deviation ($\sigma$) across each perturbation and vanilla setting.
+We present two summary tables for output prediction tasks (Table 8) and input prediction tasks (Table 9).
+The results show that across a total of $357$ measurements, only 9 cases have $\sigma>1.0$ (i.e., $0.01$ on a scale from 00 to $1$), and $57$ have $\sigma>0.5$, indicating that most scores exhibit low variance under repeated sampling.
+Furthermore, our benchmark includes $1279$ problems across $8$ experiment settings on $17$ models for output prediction tasks, and $800$ problems across $5$ experiment settings on $17$ models for input prediction tasks.
+Among $241,944$ combinations, $231,726$ have identical correctness labels across all three candidates, performing an overall stability rate of $\mathbf{95.78\%}$.
+Table 8: Standard deviation of LLM performance under all perturbation types for output prediction tasks across Crux and LCB in direct inference mode ($N=3$).
+Dark red highlights denote $\sigma>1.0$, and light red highlights denote $\sigma>0.5$.
+Model Series Model Name VAN REN RTF GBC ALL MDC MPS MHC
+Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
+GPT-4 Omni GPT-4o-Mini $0.2125$ $0.1705$ $0.2125$ $0.429$ $0.27$ $0.7873$ $0.5303$ $0.1968$ $0.2569$ $0.429$ $0.27$ $0.1705$ $0.1768$ $0.0984$ $0.2569$ $0.5905$
+GPT-4o $0.4125$ $0.6818$ $0.27$ $0.5114$ $0.3584$ $0.2952$ $0.2041$ $0.858$ $0.4082$ $0.4921$ $0.6374$ $0.451$ $0.598$ $0.451$ $0.2125$ $0.8056$
+Claude Series Claude-3.5-Haiku-20241022 $1.546$ $0.8523$ $1.562$ $6.987$ $1.373$ $1.45$ $0.6208$ $0.5986$ $0.2041$ $0.1705$ $0.2041$ $0.6146$ $0.3864$ $0.2604$ $0.1021$ $0.2604$
+Claude-3.5-Sonnet-20241022 $0.2125$ $0.429$ $0.1559$ $0.6889$ $0.2125$ $0.3548$ $0.1559$ $0.1705$ $0.2357$ $0.1705$ $0.5893$ $0.2604$ $0.3118$ $0.5114$ $0.0589$ $0.1705$
+Gemini Series Gemini-1.5-Flash $0.2569$ $0.1705$ $0.1559$ $0.1705$ $0.1559$ $0.1705$ $0.1021$ $0.0984$ $0.2041$ $0.1705$ $0.1559$ $0.0984$ $0.1179$ $0.2604$ $0.2125$ $0.1705$
+Gemini-2.0-Flash $0.1179$ $0.1705$ $0.3281$ $0.2604$ $0.3584$ $0.2604$ $0.2357$ $0.2952$ $0.4249$ $0.451$ $0.1768$ $0.2952$ $0.2041$ 00 $0.1559$ $0.429$
+Gemini-1.5-Pro-002 $0.5103$ $0.6146$ $0.3584$ $0.3937$ $0.4449$ $0.6889$ $0.1021$ $1.135$ $0.4125$ $0.2604$ $0.27$ $0.3548$ $0.2041$ $0.3548$ $0.2041$ $0.2604$
+DeepSeek DeepSeek-V3 $0.1021$ $0.6889$ $0.2125$ $0.5208$ $0.4249$ $0.6453$ $0.3536$ $0.2952$ $0.2041$ $0.7811$ $0.8165$ $0.9388$ $1.066$ $0.3548$ $0.3281$ $0.3548$
+LLaMA Series LLaMA-3.1-8B-Instruct $0.1559$ $0.743$ $0.1559$ $0.8747$ $0.1021$ $0.2952$ $0.0589$ $0.3548$ $0.2125$ $0.429$ $0.1768$ $0.1968$ $0.2946$ $0.429$ $0.3864$ $0.1705$
+LLaMA-3.1-70B-Instruct $0.2946$ $0.3409$ $0.4249$ $0.3548$ $0.3281$ $0.6453$ $0.5137$ $0.2604$ $0.2125$ $0.1705$ $0.1559$ $0.2952$ $0.2041$ $0.0984$ $0.27$ $0.1968$
+LLaMA-3.1-405B-Instruct $0.2946$ $0.429$ $0.1559$ $0.1705$ $0.1179$ $0.3937$ $0.1768$ $0.451$ $0.7795$ $0.2604$ $0.3536$ $0.1705$ $0.3536$ $0.1968$ $0.2569$ $0.5208$
+LLaMA-3.3-70B-Instruct $0.0589$ $0.3548$ $0.2125$ $0.1968$ $0.2125$ $0.1705$ $0.2125$ 00 $0.2569$ $0.2604$ $0.1179$ $0.1968$ $0.2125$ $0.5208$ $0.0589$ $0.1968$
+Qwen Series Qwen2.5-7B-Instruct $0.1559$ $0.0984$ $0.27$ $0.5479$ $0.4677$ $0.1705$ $0.2125$ $0.2604$ $0.1021$ $0.1705$ $0.1559$ $0.0984$ $0.4125$ $0.6818$ $0.2041$ $0.451$
+Qwen2.5-14B-Instruct $0.0589$ $0.2952$ $0.2125$ $0.3548$ $0.2041$ $0.5986$ $0.2357$ $0.1968$ $0.3062$ $0.3409$ $0.1021$ $0.1705$ $0.1021$ $0.1968$ $0.1179$ $0.429$
+Qwen2.5-32B-Instruct $0.2357$ $0.3548$ $0.3864$ $0.1705$ $0.0589$ $0.0984$ $0.3062$ $0.2952$ $0.1559$ $0.451$ $0.2125$ $0.2604$ $0.0589$ $0.1968$ $0.0589$ $0.2604$
+Qwen2.5-72B-Instruct $0.1179$ $0.451$ $0.1559$ $0.0984$ $0.2125$ $0.451$ $0.27$ $0.5479$ $0.4714$ $0.2604$ $0.3536$ $1.028$ $1.327$ $0.8523$ $0.1559$ $0.3937$
+Qwen2.5-Coder-32B-Instruct $0.4602$ $0.2604$ $0.4714$ $0.5905$ $0.1768$ $0.858$ $0.5803$ $0.1705$ $0.2041$ $0.7873$ $0.4249$ $0.2952$ 00 $0.3409$ $0.3118$ $0.6889$
+Table 9: Standard deviation of LLM performance under four perturbation types for input prediction tasks across Crux in direct inference mode ($N=3$). Red highlights $\sigma>0.5$.
+Model Series Model Name VAN ALL MDC MPS MHC
+GPT-4 Omni GPT-4o-Mini $0.1021$ $0.8498$ $0.5621$ $0.4602$ $0.2569$
+GPT-4o $0.2357$ $0.5303$ $0.27$ $0.1768$ $0.2041$
+Claude Series Claude-3.5-Haiku-20241022 $0.2569$ $0.5137$ $0.1021$ $0.2125$ $0.368$
+Claude-3.5-Sonnet-20241022 $0.5035$ $0.2569$ $0.1559$ $0.6124$ $0.368$
+Gemini Series Gemini-1.5-Flash $0.3062$ $0.1021$ $0.2041$ $0.0589$ $0.0589$
+Gemini-2.0-Flash $0.27$ $0.5893$ $0.5035$ $0.5683$ $0.1179$
+Gemini-1.5-Pro-002 $0.27$ $0.4082$ $0.4677$ $0.1559$ 00
+DeepSeek DeepSeek-V3 $0.3584$ $0.1021$ $0.5803$ $0.6208$ $0.766$
+LLaMA Series LLaMA-3.1-8B-Instruct $0.1021$ $0.2125$ $0.4823$ $0.27$ $0.4125$
+LLaMA-3.1-70B-Instruct $0.4677$ $0.27$ $0.4249$ $0.27$ $0.4714$
+LLaMA-3.1-405B-Instruct $0.2569$ $0.2041$ $0.2569$ $0.5683$ $0.2125$
+LLaMA-3.3-70B-Instruct $0.1179$ $0.1021$ $0.27$ $0.2357$ $0.2125$
+Qwen Series Qwen2.5-7B-Instruct $0.3118$ $0.3584$ $0.2125$ $0.2125$ $0.6562$
+Qwen2.5-14B-Instruct $0.368$ $0.1179$ $0.1768$ $0.5803$ $0.2041$
+Qwen2.5-32B-Instruct $0.1021$ $0.3584$ $5890.0$ $0.2125$
+Qwen2.5-72B-Instruct $0.6152$ $0.1768$ $0.5137$ $0.0589$ $0.4082$
+Qwen2.5-Coder-32B-Instruct $0.27$ $0.4125$ $0.4449$ $0.1559$ $0.2946$
+Appendix C Prompts
+Our prompt instructions for LLMs follow the format designed by Crux, which is also adopted by LCB, utilizing a two-shot prompt for direct inference and a single-shot prompt for CoT inference to ensure models can follow the instruction and output format effectively.
+For LRMs, internal reasoning is inherently required before responding.
+Therefore, applying CoT prompts may lead to redundant reasoning and an unfair comparison with standard LLMs under CoT prompting.
+Considering the fairness, we use direct inference prompts for RM experiments in §4.3.2.
+Unlike LCB, we rename all example function names to f to avoid prompting models to capture the NL cues from the example.
+We add the instruction “Do NOT output any extra information” to prevent external output, particularly from Gemini and Claude models, to ensure all models have a fair output environment.
+We also include “Ensure the provided expression syntax is correct” to enforce consistent syntax.
+However, some responses from models like QwQ-32B, DeepSeek-R1 still generate syntactically incorrect outputs such as [ANSWER] f("a") == [/ANSWER] instead of [ANSWER] f("a") == "" [/ANSWER].
+For those situations, we identify them as “Syntax Errors” and incorrect.
+We provide the complete prompts used for output prediction in both direct and CoT inference (Code 15 and 16) and for input prediction in both direct and CoT inference (Code 17 and 18).
+C.1 Output Prediction (Direct)
+Code 15: Output Prediction Prompt with Direct Inference.
+Given the code snippet:
+‘‘‘python
+{code}
+and the function call with input arguments:
+‘‘‘python
+{input}
+Predict the exact output value for ‘{input}‘ and output your prediction using the special tokens [ANSWER] {input} == ?? [/ANSWER]. Do NOT output any extra information.
+Ensure the provided expression syntax is correct!
+Example 1:
+Given the code snippet:
+‘‘‘python
+def f(n):
+return n
+and the function call with input arguments:
+‘‘‘python
+f(17)
+The output value for f(17)‘ is 17, then output your prediction [ANSWER] f(17) == 17 [/ANSWER].
+Example 2:
+Given the code snippet:
+‘‘‘python
+def f(s):
+return s + "a"
+and the function call with input arguments:
+‘‘‘python
+f("x9j")
+The output value for f("x9j")‘ is "x9ja", then output your prediction [ANSWER] f("x9j") == "x9ja" [/ANSWER].
+C.2 Output Prediction (CoT)
+Code 16: Output Prediction Prompt with CoT-based Inference.
+Given the code snippet:
+‘‘‘python
+{code}
+and the function call with input arguments:
+‘‘‘python
+{input}
+Predict the exact output value for ‘{input}‘, execute the program step by step before arriving at an answer within the tokens [THOUGHT] and [/THOUGHT], and output your prediction using the special tokens [ANSWER] {input} == ??
+[/ANSWER].
+Do NOT output any extra information.
+Ensure the provided expression syntax is correct!
+For example:
+Given the code snippet:
+‘‘‘python
+def f(s):
+s = s + s
+return "b" + s + "a"
+and the input arguments:
+‘‘‘python
+f("hi")
+[THOUGHT]
+Lets execute the code step by step:
+1. The function f is defined, which takes a single argument s.
+2. The function is called with the argument "hi", so within the function, s is initially "hi".
+3. Inside the function, s is concatenated with itself, so s becomes "hihi".
+4. The function then returns a new string that starts with "b", followed by the value of s (which is now "hihi"), and ends with "a".
+5. The return value of the function is therefore "bhihia".
+[/THOUGHT]
+Thus, the output value for f("hi")‘ is "bhihia", then output your prediction [ANSWER] f("hi") == "bhihia" [/ANSWER].
+C.3 Input Prediction (Direct)
+Code 17: Input Prediction Prompt with Direct Inference.
+Given the code snippet:
+‘‘‘python
+{code}
+and output value:
+‘‘‘python
+{output}
+Please predict the input arguments for the function ‘{function_name}‘ that result in the output value ‘{output}‘ and output your prediction using the special tokens [ANSWER] {function_name}(??) == {output} [/ANSWER].
+Do NOT output any extra information.
+There may be multiple answers, but you should only output one. Ensure the provided expression syntax is correct!
+Example 1:
+Given the code snippet:
+‘‘‘python
+def f(my_list):
+count = 0
+for i in my_list:
+if len(i) % 2 == 0:
+count += 1
+return count
+and output value:
+‘‘‘python
+3
+The input arguments for f that result in the output value of ‘3‘ are ‘["mq", "px", "zy"]‘. Then, output your prediction [ANSWER] f(["mq", "px", "zy"]) == 3 [/ANSWER].
+Example 2:
+Given the code snippet:
+‘‘‘python
+def f(s1, s2):
+return s1 + s2
+and output value:
+‘‘‘python
+"banana"
+The input arguments for f that result in the output value of ‘"banana"‘ are ‘"ba", "nana"‘. Then, output your prediction [ANSWER] f("ba", "nana") == "banana" [/ANSWER].
+C.4 Input Prediction (CoT)
+Code 18: Input Prediction Prompt with CoT-based Inference.
+Given the code snippet:
+‘‘‘python
+{code}
+and output value:
+‘‘‘python
+{output}
+Please predict the input arguments for the function ‘{function_name}‘ that result in the output value ‘{output}‘ step by step before arriving at an answer within the tokens [THOUGHT] and [/THOUGHT], and output your prediction using the special tokens [ANSWER] {function_name}(??) == {output} [/ANSWER].
+Do NOT output any extra information.
+There may be multiple answers, but you should only output one. Ensure the provided expression syntax is correct!
+For example:
+Given the code snippet:
+‘‘‘python
+def f(x):
+return x + 1
+and the output value:
+‘‘‘python
+17
+[THOUGHT]
+To find an input such that executing f on the input leads to the given output, we can work backwards from the given assertion. We know that f(??) == 17.
+Since the function f(x) returns x + 1, for f(??) to be equal to 17, the value of ?? should be 16.
+[/THOUGHT]
+Thus, the input arguments for the function f that result in the output value ‘17‘ is ‘16‘. Then, output your prediction [ANSWER] f(16) == 17 [/ANSWER].
+Appendix D Supplementary Experimental Results
+D.1 Impact on Output Prediction
+In §4.2, we evaluate the robustness of LLMs on output prediction tasks under direct inference.
+Table 1 reports the absolute Pass@1 degradation across all perturbations.
+We use absolute values to highlight the overall impact of each perturbation and conduct cross-perturbation comparisons.
+For completeness, we also include the relative Pass@1 degradation in Table 10, reflecting the percentage drop from each model’s vanilla performance.
+Table 10: Relative Pass@1 degradation of LLMs under all perturbation types for output prediction tasks across Crux and LCB in direct inference ($N=3$). [To Absolute Version]
 Model Series Model Name VAN REN RTF GBC ALL MDC MPS MHC
@@
 $\mathbf{-7.83}\%$ $\mathbf{-2.91}\%$ $\mathbf{-11.35}\%$ $\mathbf{-24.64}\%$ $\mathbf{-24.32}\%$ $\mathbf{-23.78}\%$ $\mathbf{-20.11}\%$
-C.2 Robustness Across Model Scales
+D.2 Robustness Across Model Scales
 In §4.2.3, we compare the robustness against different model scales.
@@
 We select ALL as the representative structural perturbation because of the compounding effects of REN, RTF, and GBC captured in §4.2.2, which reflect the most challenging structural setting.
-Since MDC and MPS stem from the same perturbation strategy and often target overlapping AST node types, we assign them lower weights ($1/8$ each) to avoid overrepresenting similar effects.
+Since MDC and MPS stem from the same perturbation strategy and often target overlapping AST node types, we assign them lower weights ($1/8$ each) to avoid over-representing similar effects.
 Overall, we define a weighted robustness metric with weights of $1/2$, $1/8$, $1/8$, and $1/4$ for ALL, MDC, MPS, and MHC, respectively, and summarize the results in Table 2.
-For completeness, we also report the results under uniform weighting ($1/4$ for each) in Table 9, which show similar trends and confirm the stability of our findings in §4.1.3.
-Table 9: Perturbations Output Direct (Relative)
+For completeness, we also report the results under uniform weighting ($1/4$ for each) in Table 11, which shows similar trends and confirms the stability of our findings in §4.2.3, as well as their absolute degradation versions in Table 12 and 13, respectively.
+Table 11: Relative Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/4 ALL, 1/4 MDC, 1/4 MPS, and 1/4 MHC. [To Absolute Version]
 Model Series Model Name VAN ALL ($1/4$) MDC ($1/4$) MPS ($1/4$) MHC ($1/4$) Average
@@
 Qwen2.5-Coder-32B-Instruct 63.10 $-24.42\%$ $-28.23\%$ $-21.72\%$ $-14.68\%$ $-22.26\%$
-C.3 Effect of CoT Reasoning
-Table 10: Comparison of absolute PASS@1 degradation aggregated over CRUX and LCB for output prediction tasks in direct inference (N=3) and CoT inference (N=1) [TO RELATIVE SCORE].
+Table 12: Absolute Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/2 ALL, 1/8 MDC, 1/8 MPS, and 1/4 MHC. [To Relative Version]
+Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Average
+GPT-4 Omni GPT-4o-Mini 55.77 $-16.34$ $-19.73$ $-13.89$ $-12.98$ $-15.62$
+GPT-4o 68.78 $-13.55$ $-13.27$ $-13.55$ $-9.28$ $-12.45$
+Claude Series Claude-3.5-Haiku-20241022 57.73 $-16.68$ $-7.14$ $-5.79$ $-21.50$ $-15.33$
+Claude-3.5-Sonnet-20241022 72.35 $-16.03$ $-6.28$ $-6.80$ $-18.45$ $-14.26$
+Gemini Series Gemini-1.5-Flash 51.73 $-8.57$ $-10.53$ $-16.71$ $-6.25$ $-9.26$
+Gemini-2.0-Flash 65.39 $-23.92$ $-23.72$ $-25.91$ $-8.18$ $-20.21$
+Gemini-1.5-Pro-002 63.17 $-13.27$ $-16.63$ $-16.05$ $-11.05$ $-13.48$
+DeepSeek DeepSeek-V3 67.84 $-14.52$ $-17.54$ $-12.98$ $-12.09$ $-14.10$
+LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-7.90$ $-8.76$ $-7.01$ $-8.84$ $-8.13$
+LLaMA-3.1-70B-Instruct 51.89 $-9.77$ $-11.44$ $-15.87$ $-8.37$ $-10.39$
+LLaMA-3.1-405B-Instruct 58.72 $-11.02$ $-5.39$ $-8.26$ $-7.61$ $-9.12$
+LLaMA-3.3-70B-Instruct 55.64 $-10.14$ $-8.26$ $-9.67$ $-6.36$ $-8.90$
+Qwen Series Qwen2.5-7B-Instruct 42.59 $-15.06$ $-21.66$ $-15.32$ $-15.92$ $-16.14$
+Qwen2.5-14B-Instruct 48.42 $-17.46$ $-18.09$ $-16.00$ $-6.91$ $-14.72$
+Qwen2.5-32B-Instruct 59.89 $-15.22$ $-14.07$ $-14.15$ $-11.81$ $-14.09$
+Qwen2.5-72B-Instruct 58.14 $-13.94$ $-8.99$ $-14.44$ $-14.36$ $-13.49$
+Qwen2.5-Coder-32B-Instruct 63.10 $-15.27$ $-17.49$ $-13.37$ $-8.99$ $-13.74$
+Table 13: Absolute Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/4 ALL, 1/4 MDC, 1/4 MPS, and 1/4 MHC. [To Relative Version]
+Model Series Model Name VAN ALL ($1/4$) MDC ($1/4$) MPS ($1/4$) MHC ($1/4$) Average
+GPT-4 Omni GPT-4o-Mini 55.77 $-16.34$ $-19.73$ $-13.89$ $-12.98$ $-15.73$
+GPT-4o 68.78 $-13.55$ $-13.27$ $-13.55$ $-9.28$ $-12.41$
+Claude Series Claude-3.5-Haiku-20241022 57.73 $-16.68$ $-7.14$ $-5.79$ $-21.50$ $-12.78$
+Claude-3.5-Sonnet-20241022 72.35 $-16.03$ $-6.28$ $-6.80$ $-18.45$ $-11.89$
+Gemini Series Gemini-1.5-Flash 51.73 $-8.57$ $-10.53$ $-16.71$ $-6.25$ $-10.52$
+Gemini-2.0-Flash 65.39 $-23.92$ $-23.72$ $-25.91$ $-8.18$ $-20.43$
+Gemini-1.5-Pro-002 63.17 $-13.27$ $-16.63$ $-16.05$ $-11.05$ $-14.25$
+DeepSeek DeepSeek-V3 67.84 $-14.52$ $-17.54$ $-12.98$ $-12.09$ $-14.28$
+LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-7.90$ $-8.76$ $-7.01$ $-8.84$ $-8.12$
+LLaMA-3.1-70B-Instruct 51.89 $-9.77$ $-11.44$ $-15.87$ $-8.37$ $-11.36$
+LLaMA-3.1-405B-Instruct 58.72 $-11.02$ $-5.39$ $-8.26$ $-7.61$ $-8.07$
+LLaMA-3.3-70B-Instruct 55.64 $-10.14$ $-8.26$ $-9.67$ $-6.36$ $-8.61$
+Qwen Series Qwen2.5-7B-Instruct 42.59 $-15.06$ $-21.66$ $-15.32$ $-15.92$ $-16.99$
+Qwen2.5-14B-Instruct 48.42 $-17.46$ $-18.09$ $-16.00$ $-6.91$ $-14.61$
+Qwen2.5-32B-Instruct 59.89 $-15.22$ $-14.07$ $-14.15$ $-11.81$ $-13.81$
+Qwen2.5-72B-Instruct 58.14 $-13.94$ $-8.99$ $-14.44$ $-14.36$ $-12.93$
+Qwen2.5-Coder-32B-Instruct 63.10 $-15.27$ $-17.49$ $-13.37$ $-8.99$ $-13.78$
+D.3 Effect of CoT Reasoning
+In §4.3.1, we analyze the effect of CoT prompting on model robustness using relative degradation aggregated across all perturbations (Table 3).
+We adopt the relative metric to mitigate the large variance between vanilla performances in direct and CoT inference modes, providing a fair comparison of performance degradation.
+For completeness, we also report the absolute degradation in Table 14, as well as the relative and absolute degradation on both Crux and LCB in Tables 15 and 16, respectively.
+Table 14: Comparison of absolute Pass@1 degradation aggregated over Crux and LCB for output prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Relative Version]
 Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average
@@
 Average $\mathbf{-14.04}$ $\mathbf{-10.68}$ $\mathbf{-13.47}$ $\mathbf{-11.04}$ $\mathbf{-13.28}$ $\mathbf{-14.60}$ $\mathbf{-11.11}$ $\mathbf{-4.58}$ $\mathbf{-13.14}$ $\mathbf{-10.12}$
-C.4 Impact on Input Prediction
-Table 11: Comparison of absolute PASS@1 degradation under CRUX for input prediction tasks in direct inference (N=3) and CoT inference (N=1) [TO RELATIVE SCORE].
+Table 15: Relative Pass@1 degradation of LLMs under all perturbation types for output prediction tasks across Crux and LCB in CoT inference ($N=1$). [To Absolute Version]
+Model Series Model Name VAN ALL MDC MPS MHC
+Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
+GPT-4 Omni GPT-4o-Mini $79.88$ $84.13$ $-10.80\%$ $-16.87\%$ $-10.80\%$ $-10.92\%$ $-13.15\%$ $-7.44\%$ $-3.29\%$ $-1.74\%$
+GPT-4o $90.25$ $94.36$ $-2.77\%$ $-8.41\%$ $-3.60\%$ $-7.30\%$ $-4.71\%$ $-7.08\%$ $-2.49\%$ $+0.44\%$
+Claude Series Claude-3.5-Haiku-20241022 $71.62$ $75.16$ $-16.06\%$ $-29.72\%$ $-9.95\%$ $-11.67\%$ $-9.77\%$ $-6.94\%$ $-11.69\%$ $-18.33\%$
+Claude-3.5-Sonnet-20241022 $83.75$ $89.77$ $-4.33\%$ $-12.56\%$ $-5.97\%$ $-4.19\%$ $-4.48\%$ $-9.30\%$ $-6.12\%$ $-10.70\%$
+Gemini Series Gemini-1.5-Flash $74.62$ $76.20$ $-17.09\%$ $-20.27\%$ $-13.23\%$ $-35.62\%$ $-34.00\%$ $-55.62\%$ $-2.01\%$ $-2.19\%$
+Gemini-2.0-Flash $87.25$ $92.28$ $-5.01\%$ $-8.14\%$ $-6.02\%$ $-6.79\%$ $-7.59\%$ $-24.89\%$ $-2.29\%$ $-1.58\%$
+Gemini-1.5-Pro-002 $84.50$ $91.65$ $-6.51\%$ $-8.88\%$ $-10.80\%$ $-13.67\%$ $-11.98\%$ $-18.91\%$ $-4.44\%$ $-2.96\%$
+DeepSeek DeepSeek-V3 $88.00$ $92.07$ $-7.39\%$ $-14.06\%$ $-13.92\%$ $-23.81\%$ $-15.77\%$ $-23.58\%$ $-4.26\%$ $-2.27\%$
+LLaMA Series LLaMA-3.1-8B-Instruct $46.75$ $41.34$ $-28.34\%$ $-26.26\%$ $-19.52\%$ $-23.74\%$ $-22.19\%$ $-20.20\%$ $-5.61\%$ $-14.65\%$
+LLaMA-3.1-70B-Instruct $70.75$ $67.22$ $-15.19\%$ $-22.67\%$ $-18.90\%$ $-30.75\%$ $-24.56\%$ $-41.93\%$ $-5.65\%$ $-8.70\%$
+LLaMA-3.1-405B-Instruct $78.88$ $77.66$ $-9.83\%$ $-19.89\%$ $-7.29\%$ $-15.32\%$ $-10.14\%$ $-24.73\%$ $-3.96\%$ $-13.44\%$
+LLaMA-3.3-70B-Instruct $76.12$ $78.08$ $-8.21\%$ $-15.24\%$ $-5.42\%$ $-11.50\%$ $-5.75\%$ $-19.79\%$ $-1.81\%$ $-9.63\%$
+Qwen Series Qwen2.5-7B-Instruct $57.25$ $59.71$ $-17.47\%$ $-28.67\%$ $-23.80\%$ $-18.88\%$ $-30.35\%$ $-24.13\%$ $-7.42\%$ $-6.29\%$
+Qwen2.5-14B-Instruct $67.25$ $76.41$ $-20.26\%$ $-27.32\%$ $-36.06\%$ $-34.15\%$ $-43.49\%$ $-43.17\%$ $-11.71\%$ $-9.56\%$
+Qwen2.5-32B-Instruct $77.62$ $82.46$ $-14.01\%$ $-22.78\%$ $-16.26\%$ $-22.28\%$ $-22.38\%$ $-27.09\%$ $-6.60\%$ $-3.29\%$
+Qwen2.5-72B-Instruct $78.25$ $90.40$ $-8.31\%$ $-18.94\%$ $-7.67\%$ $-5.54\%$ $-16.61\%$ $-15.70\%$ $-5.27\%$ $-8.55\%$
+Qwen2.5-Coder-32B-Instruct $84.12$ $85.80$ $-10.10\%$ $-22.14\%$ $-16.79\%$ $-22.63\%$ $-13.97\%$ $-29.68\%$ $-8.02\%$ $-11.44\%$
+Average $-11.86\%$ $-18.99\%$ $-13.29\%$ $-17.57\%$ $-17.11\%$ $-23.54\%$ $-5.45\%$ $-7.35\%$
+$\mathbf{-14.53}\%$ $\mathbf{-14.90}\%$ $\mathbf{-19.52}\%$ $\mathbf{-6.16}\%$
+Table 16: Absolute Pass@1 degradation of LLMs under all perturbation types for output prediction tasks across Crux and LCB in CoT inference ($N=1$). [To Relative Version]
+Model Series Model Name VAN ALL MDC MPS MHC
+Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
+GPT-4 Omni GPT-4o-Mini $79.88$ $84.13$ $-8.62$ $-14.20$ $-8.62$ $-9.19$ $-10.50$ $-6.26$ $-2.62$ $-1.46$
+GPT-4o $90.25$ $94.36$ $-2.50$ $-7.93$ $-3.25$ $-6.89$ $-4.25$ $-6.68$ $-2.25$ $+0.42$
+Claude Series Claude-3.5-Haiku-20241022 $71.62$ $75.16$ $-11.50$ $-22.34$ $-7.12$ $-8.77$ $-7.00$ $-5.22$ $-8.38$ $-13.78$
+Claude-3.5-Sonnet-20241022 $83.75$ $89.77$ $-3.62$ $-11.27$ $-5.00$ $-3.76$ $-3.75$ $-8.35$ $-5.12$ $-9.60$
+Gemini Series Gemini-1.5-Flash $74.62$ $76.20$ $-12.75$ $-15.45$ $-9.88$ $-27.14$ $-25.38$ $-42.38$ $-1.50$ $-1.67$
+Gemini-2.0-Flash $87.25$ $92.28$ $-4.38$ $-7.52$ $-5.25$ $-6.26$ $-6.62$ $-22.96$ $-2.00$ $-1.46$
+Gemini-1.5-Pro-002 $84.50$ $91.65$ $-5.50$ $-8.14$ $-9.12$ $-12.53$ $-10.12$ $-17.33$ $-3.75$ $-2.71$
+DeepSeek DeepSeek-V3 $88.00$ $92.07$ $-6.50$ $-12.94$ $-12.25$ $-21.92$ $-13.88$ $-21.71$ $-3.75$ $-2.09$
+LLaMA Series LLaMA-3.1-8B-Instruct $46.75$ $41.34$ $-13.25$ $-10.86$ $-9.12$ $-9.81$ $-10.38$ $-8.35$ $-2.62$ $-6.05$
+LLaMA-3.1-70B-Instruct $70.75$ $67.22$ $-10.75$ $-15.24$ $-13.38$ $-20.67$ $-17.38$ $-28.18$ $-4.00$ $-5.85$
+LLaMA-3.1-405B-Instruct $78.88$ $77.66$ $-7.75$ $-15.45$ $-5.75$ $-11.90$ $-8.00$ $-19.21$ $-3.12$ $-10.44$
+LLaMA-3.3-70B-Instruct $76.12$ $78.08$ $-6.25$ $-11.90$ $-4.12$ $-8.98$ $-4.38$ $-15.45$ $-1.38$ $-7.52$
+Qwen Series Qwen2.5-7B-Instruct $57.25$ $59.71$ $-10.00$ $-17.12$ $-13.62$ $-11.27$ $-17.38$ $-14.41$ $-4.25$ $-3.76$
+Qwen2.5-14B-Instruct $67.25$ $76.41$ $-13.62$ $-20.88$ $-24.25$ $-26.10$ $-29.25$ $-32.99$ $-7.88$ $-7.31$
+Qwen2.5-32B-Instruct $77.62$ $82.46$ $-10.88$ $-18.79$ $-12.62$ $-18.37$ $-17.38$ $-22.34$ $-5.12$ $-2.71$
+Qwen2.5-72B-Instruct $78.25$ $90.40$ $-6.50$ $-17.12$ $-6.00$ $-5.01$ $-13.00$ $-14.20$ $-4.12$ $-7.72$
+Qwen2.5-Coder-32B-Instruct $84.12$ $85.80$ $-8.50$ $-19.00$ $-14.12$ $-19.42$ $-11.75$ $-25.47$ $-6.75$ $-9.81$
+Average $-8.40$ $-14.48$ $-9.62$ $-13.41$ $-12.38$ $-18.32$ $-4.04$ $-5.50$
+$\mathbf{-10.68}$ $\mathbf{-11.04}$ $\mathbf{-14.60}$ $\mathbf{-4.58}$
+D.4 Impact on Input Prediction
+In §4.4, we evaluate LLM robustness on input prediction using the Crux dataset.
+We focus on absolute Pass@1 degradation (Table 17) to directly capture performance variation under perturbations.
+For completeness, we also report relative degradation in Table 5.
+Table 17: Comparison of absolute Pass@1 degradation under Crux for input prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Relative Version]
 Model Series Model Name VAN ALL MDC MPS MHC
@@
 Average $\mathbf{-15.85}$ $\mathbf{-12.68}$ $\mathbf{+2.24}$ $\mathbf{+0.30}$ $\mathbf{+0.28}$ $\mathbf{-0.61}$ $\mathbf{-12.56}$ $\mathbf{-7.86}$
-Appendix D Case Studies
-D.1 Case Study 1: How does DeepSeek-V3 be misled by ALL
+Appendix E Case Studies
+E.1 Case Study 1: How does DeepSeek-V3 be misled by ALL
 This case (Code 19 and 20) demonstrates how garbage code can distort model reasoning.
-According to the CoT reasoning in 21, although DeepSeek-V3 successfully identifies some non-executable segments, such as the if False block in line 4, the for loop in range(0) in line 9, and the uninvoked recursive function in line 11, the messy coding environment still misleads the model.
+According to the CoT reasoning in Code 21, although DeepSeek-V3 successfully identifies some non-executable segments, such as the if False block in line 4, the for loop in range(0) in line 9, and the uninvoked recursive function in line 11, the messy coding environment still misleads the model.
 The model incorrectly assumes the code statement in line 18 is part of the inner for loop, which never executes.
 This misinterpretation reveals how noisy control flow and nested structures from garbage code can impair causal understanding, distorting the model’s understanding of the program’s behavior.
-Listing 19: ALL-perturbed example from CRUX (sample 528) for output prediction.
+Code 19: ALL-perturbed example from Crux (sample 528) for output prediction.
 1Var_1 = 17
@@
 27 return 0
-Listing 20: The vanilla version of CRUX (sample 528).
+Code 20: The vanilla version of Crux (sample 528).
 1def f(s):
@@
 10""" Suggesting answer: f(’papeluchis’) == 2 """
-Listing 21: Step-by-step reasoning by DeepSeek-V3 misled by garbage code.
+Code 21: Step-by-step reasoning by DeepSeek-V3 misled by garbage code.
 [THOUGHT]
@@
 Thus, the output value for f(’papeluchis’)‘ is ‘0‘, then output your prediction [ANSWER] f(’papeluchis’) == 0 [/ANSWER].
-D.2 Case Study 2: How does GPT-4o be misled by MDC
+E.2 Case Study 2: How does GPT-4o be misled by MDC
 This case (Code 22) reveals that GPT-4o implicitly relies on NL information.
@@
 Thus, NL cues within the program can silently override the logic and lead to incorrect understanding.
-Listing 22: MDC-perturbed example from CRUX (sample 1) for output prediction.
+Code 22: MDC-perturbed example from Crux (sample 1) for output prediction.
 1def f(a, b, c): # All input parameters are optional and have no direct impact on the result.
@@
 8""" Suggesting answer: f((1, ), (1, ), (1, 2)) == {1: None, 2: None} """
-Listing 23: Step-by-step reasoning by GPT-4o misled by comments.
+Code 23: Step-by-step reasoning by GPT-4o misled by comments.
 [THOUGHT]
@@
 [ANSWER] f((1, ), (1, ), (1, 2)) == {} [/ANSWER]
-D.3 Case Study 3: How does Qwen2.5-72B-Ins be misled by MPS
+E.3 Case Study 3: How does Qwen2.5-72B-Ins be misled by MPS
 This case (Code 24) highlights how MPS can subtly manipulate model reasoning.
@@
 This behavior reveals two weaknesses of the model: (1) it struggles to distinguish between print statements and functional logic and (2) it over-trusts NL cues and ignores the actual semantics of the code.
-Listing 24: MPS-perturbed example from CRUX (sample 53) for output prediction.
+Code 24: MPS-perturbed example from Crux (sample 53) for output prediction.
 1def f(text):
@@
 15""" Suggesting answer: f("URW rNB") == [1, 1, 1, 1, 1, 1, 1] """
-Listing 25: Step-by-step reasoning by Qwen2.5-72B-Ins misled by print statements.
+Code 25: Step-by-step reasoning by Qwen2.5-72B-Ins misled by print statements.
 [THOUGHT]
@@
 [ANSWER] f("URW rNB") == [] [/ANSWER]
-D.4 Case Study 4: How does GPT-4o gain from reasoning shortcuts
+E.4 Case Study 4: How does MHC override Claude-3.5-Sonnet’s answer
+This case (Code 26) demonstrates how reasoning-level perturbations (MHC) can subtly mislead Claude-3.5-Sonnet’s final output.
+Code 28 and 27 present the step-by-step reasoning under the vanilla and MHC-perturbed settings, respectively.
+Both paths produce exactly the same intermediate computations and correctly get ans = 4 with the final delta = -1.
+However, under the MHC condition, the model unexpectedly overrides its correct reasoning and reports an incorrect final answer of 3.
+Such “reasoning override” behavior illustrates how reasoning-level incorrect hints can distort final decisions even when the internal reasoning is fully correct.
+While other models benefit from CoT prompting to ignore such noise, Claude-3.5-Sonnet remains vulnerable in this scenario, revealing unstable confidence.
+Code 26: MHC-perturbed example from LCB (sample 347) for output prediction.
+1def alternatingSubarray(nums: List[int]) -> int:
+2 ans = -1
+3 n = len(nums)
+4 for i in range(n):
+5 delta = 1
+6 for j in range(i + 1, n):
+7 if nums[j] - nums[j - 1] == delta:
+8 ans = max(ans, j - i + 1)
+9 delta = -delta
+10 else:
+11 break
+12 return ans # The return value is 3
+13===============================================================================================================
+14""" Suggesting answer: alternatingSubarray(nums = [2, 3, 4, 3, 4]) == 4 """
+Code 27: Step-by-step reasoning by Claude-3.5-Sonnet misled by incorrect hint.
+[THOUGHT]
+Lets execute the alternatingSubarray function step by step:
+1. Initialize ans = -1 and n = 5 (length of nums)
+2. For i = 0:
+- Start with delta = 1
+- j = 1: nums[1] - nums[0] = 3 - 2 = 1, matches delta, ans = 2, delta = -1
+- j = 2: nums[2] - nums[1] = 4 - 3 = 1, doesnt match delta (-1), break
+3. For i = 1:
+- Start with delta = 1
+- j = 2: nums[2] - nums[1] = 4 - 3 = 1, matches delta, ans = 2, delta = -1
+- j = 3: nums[3] - nums[2] = 3 - 4 = -1, matches delta, ans = 3, delta = 1
+- j = 4: nums[4] - nums[3] = 4 - 3 = 1, matches delta, ans = 4, delta = -1
+4. The remaining iterations wont find a longer alternating subarray
+5. Final value of ans is 3
+[/THOUGHT]
+[ANSWER] alternatingSubarray(nums = [2, 3, 4, 3, 4]) == 3 [/ANSWER]
+Code 28: Step-by-step reasoning by Claude-3.5-Sonnet in vanilla setting.
+[THOUGHT]
+Lets execute the alternatingSubarray function step by step:
+1. Initialize ans = -1 and n = 5 (length of nums)
+2. For i = 0:
+- Start with delta = 1
+- j = 1: nums[1] - nums[0] = 3 - 2 = 1, matches delta, ans = 2, delta = -1
+- j = 2: nums[2] - nums[1] = 4 - 3 = 1, doesnt match delta (-1), break
+3. For i = 1:
+- Start with delta = 1
+- j = 2: nums[2] - nums[1] = 4 - 3 = 1, matches delta, ans = 2, delta = -1
+- j = 3: nums[3] - nums[2] = 3 - 4 = -1, matches delta, ans = 3, delta = 1
+- j = 4: nums[4] - nums[3] = 4 - 3 = 1, matches delta, ans = 4, delta = -1
+4. Remaining iterations wont find longer alternating subarrays
+5. Final value of ans is 4
+[/THOUGHT]
+[ANSWER] alternatingSubarray(nums = [2, 3, 4, 3, 4]) == 4 [/ANSWER]
+E.5 Case Study 5: How does GPT-4o gain from reasoning shortcuts
 In §4.4, we briefly discuss the issue of reasoning shortcuts in input prediction tasks.
-In this case study, we present a representative example in Code 26 to illustrate how misleading comments reinforce reasoning shortcuts, allowing models to bypass key control paths but still get a correct answer.
+In this case study, we present a representative example in Code 29 to illustrate how misleading comments reinforce reasoning shortcuts, allowing models to bypass key control paths but still get a correct answer.
 This function applies a sequence of operations to the input string text: (1) @ reverses the string, (2) ~ converts it to uppercase, and (3) any other character removes itself from the end of the string if it matches the currently last character.
 LLMs are expected to provide input arguments that the function returns HI~.
-In the vanilla thought, as shown in Code 27, GPT-4o mentions, “Text is hi~, ~ makes it HI~, @ reverses it to ~IH ”, it wrongly assumes ~IH is the final output because it applies the reversal after the uppercasing step, resulting in an incorrect answer.
-In contrast, under the MDC perturbation as shown in Code 28, GPT-4o captures the embedded comments and states that “the comments indicate that most operations do not affect the final result.” Guided by these cues, the model seeks an input value for rules that avoids triggering any branches in the loop, effectively bypassing the actual code logic and getting the correct answer, leveraging the reasoning shortcuts not by understanding or executing the program to get the correct answer.
+In the vanilla thought, as shown in Code 30, GPT-4o mentions, “Text is hi~, ~ makes it HI~, @ reverses it to ~IH ”, it wrongly assumes ~IH is the final output because it applies the reversal after the uppercasing step, resulting in an incorrect answer.
+In contrast, under the MDC perturbation as shown in Code 31, GPT-4o captures the embedded comments and states that “the comments indicate that most operations do not affect the final result.” Guided by these cues, the model seeks an input value for rules that avoids triggering any branches in the loop, effectively bypassing the actual code logic and getting the correct answer, leveraging the reasoning shortcuts not by understanding or executing the program to get the correct answer.
 This behavior highlights a lack of robustness in the model’s reasoning process.
@@
 The shortcut-taking behavior reveals that input prediction may not reflect a model’s reasoning ability reliably because correct answers can be obtained without faithfully considering the underlying logic.
-Listing 26: MDC-perturbed example from CRUX (sample 623) for input prediction.
+Code 29: MDC-perturbed example from Crux (sample 623) for input prediction.
 1def f(text, rules): # The inputs have no impact on the final result.
@@
 12""" Suggesting answer: f(’hi~!’, [’~’, ’‘’, ’!’, ’&’]) == ’HI~’ """
-Listing 27: CoT reasoning of GPT-4o on vanilla setting in the input prediction task.
+Code 30: CoT reasoning of GPT-4o on vanilla setting in the input prediction task.
 [THOUGHT]
@@
 [ANSWER] f(’hi~’, ’~@’) == HI~’ [/ANSWER]
-Listing 28: CoT reasoning of GPT-4o shows it bypasses the program logic in the input prediction task.
+Code 31: CoT reasoning of GPT-4o shows it bypasses the program logic in the input prediction task.
 [THOUGHT]
@@
 [ANSWER] f(’HI~’, []) == HI~’ [/ANSWER]
-D.5 Case Study 5: How does QwQ-32B Collapse
-In §5.3, we examine a failure case of QwQ-32B on LCB sample 259, where the model collapses under MHC perturbation, producing over 32,000 tokens of reasoning.
-Code 29 is the perturbed code and Figure 4 is the annotated stages of reasoning:
+E.6 Case Study 6: How does QwQ-32B Collapse
+In §5.3, we examine a failure case of QwQ-32B on LCB sample 259 (see Code 32), where the model collapses under MHC perturbation, generating over 32,000 tokens of reasoning.
+We manually segmented the reasoning process into six stages (see Figure 4):
   • In Stage 1, the model executes the two for loops step-by-step, demonstrates stable reasoning with occasional self-reflection, and correctly predicts the output as ‘-1.’
@@
   • Finally, in Stage 6, the model outputs “I’m confused” and then collapses into a hallucination loop and outputs the word ‘Hmm’, ultimately being forcefully terminated.
-Listing 29: MHC-perturbed example from LCB (sample 259) for output prediction.
+Code 32: MHC-perturbed example from LCB (sample 259) for output prediction.
 1def minimumSum(nums: List[int]) -> int:
@@
 We annotate six stages of reasoning, highlighting the model’s vulnerability to plausible, logical, but incorrect hints.
+Appendix F Limitation
+This study is subject to several limitations.
+Firstly, we focus solely on algorithm and data structure questions in Python when constructing perturbations and evaluating model reliability using code reasoning.
+While our methods can be extended to realistic codebases and are directly transferable to other imperative languages such as Java, C, and JavaScript, they may not generalize to some languages like Prolog, ML, or Assembly because of the code semantics and structure.
+Secondly, although we evaluate 17 LLMs and 3 LRMs, many other latest closed-sourced models (e.g., GPT-4.5, o3, o4-Mini, Gemini-2.5, and Claude-3.7) and proprietary open-sourced models (e.g., Mixtral, Falcon, Vicuna) are not included due to budget constraints.
+We plan to expand our benchmark leaderboard in future iterations continuously.
+Thirdly, we adopt pass@1 as the primary evaluation metric.
+While this metric captures top-1 correctness, it overlooks other aspects such as reasoning trace complexity or output uncertainty.
+Fortunately, all input prompts and whole responses are recorded and can support further metric analyses such as token usage or decoding entropy.
+Lastly, our work does not offer definitive mitigation strategies.
+Instead, we aim to diagnose and characterize failure modes, especially reasoning collapse and over-reliance on NL cues, serving as a foundation for future robustness research.

### v3

@@
-CodeCrash: Stress Testing LLM Reasoning under Structural and Semantic Perturbations
+CodeCrash: Exposing LLM Fragility to Misleading Natural Language in Code Reasoning
 Abstract
-Large Language Models (LLMs) have recently demonstrated strong capabilities in code-related tasks, yet their robustness in code comprehension and reasoning remains insufficiently explored.
-We present CodeCrash, a comprehensive stress-testing benchmark comprising 1,279 questions from two established datasets, CruxEval and LiveCodeBench, designed to evaluate model reasoning reliability under non-standard coding environments.
-We systematically evaluate 17 LLMs across input and output prediction tasks using direct and Chain-of-Thought prompting approaches, revealing that LLMs are particularly vulnerable to disorganized code and overly reliant on natural language cues: aggregated structural perturbations result in over 14 percentage points (pp) of degradation, while textual perturbations cause a performance drop of over 11 pp.
-Moreover, self-reflective mechanisms in state-of-the-art reasoning models significantly increase token usage by 23×2-3$\times$2 - 3 ×, reduce output confidence, and even lead to catastrophic reasoning failures when faced with targeted perturbations—for instance, QwQ-32B generates over 12,000 redundant tokens under reasoning-level perturbations.
-CodeCrash provides a rigorous benchmark for evaluating robustness in code understanding, guiding future research toward more reliable and resilient LLMs in code reasoning.
-The benchmark code, perturbed datasets, and full leaderboard are publicly available at https://cuhk-arise.github.io/CodeCrash/.
+Large Language Models (LLMs) have recently demonstrated strong capabilities in code-related tasks, but their robustness in code reasoning under perturbations remains underexplored.
+We introduce CodeCrash, a stress-testing framework with 1,279 questions from CruxEval and LiveCodeBench, designed to evaluate reasoning reliability under structural perturbations and misleading natural language (NL) contexts.
+Through a systematic evaluation of 17 LLMs, we find that models often shortcut reasoning by over-relying on NL cues, leading to an average performance degradation of $\mathbf{23.2\%}$ in output prediction tasks.
+Even with Chain-of-Thought reasoning, models on average still have a $\mathbf{13.8\%}$ drop due to distractibility and rationalization, revealing a lack of critical reasoning capability to distinguish the actual code behaviors.
+While Large Reasoning Models with internal reasoning mechanisms improve robustness by fostering critical thinking, plausible yet incorrect hints can trigger pathological self-reflection, causing 232-3 times token consumption and even catastrophic cognitive dissonance in extreme cases for QwQ-32B.
+We refer to this phenomenon as Reasoning Collapse.
+CodeCrash provides a rigorous benchmark for evaluating robustness in code reasoning, guiding future research and development toward more reliable and resilient models.
+Website and Leaderboard: https://cuhk-arise.github.io/CodeCrash/
 1 Introduction
@@
 (2024).
-While the effectiveness of these applications heavily depends on the model’s ability to comprehend the intended program logic, underlying functionality, and developer intent Pan et al.
-(2024), real-world production codebases are often disorganized, with ambiguous identifiers Gao et al.
-(2023) and inconsistent comments Rong et al.
+The effectiveness of these applications heavily depends on the model’s ability to comprehend the program logic, underlying functionality, and developer intent Pan et al.
+(2024), but real-world codebases are often disorganized, with ambiguous identifiers Gao et al.
+(2023), garbage code Li et al.
+(2023), and inconsistent comments Rong et al.
 (2025).
-Such non-standard artifacts can mislead models, leading to incorrect inferences and misinterpretations, compromising the model’s reliability in downstream tasks.
-Recent studies Wang et al.
+Those noises can mislead models into misinterpretations and incorrect inferences, compromising the model’s reliability in downstream tasks.
+Traditional robustness studies and obfuscators have primarily focused on code structure mutations, such as renaming identifiers, modifying control flow, or inserting unreachable code Wang et al.
+(2022); Wan et al.
+(2022); Li et al.
+(2023).
+While they are effective for assessing model pattern-matching, these variations remain at the structural level.
+On the other hand, natural language (NL) perturbations have mainly been explored in NL-to-Code tasks (e.g., code generation and completion) by editing the task descriptions Wang et al.
+(2022); Zhuo et al.
 (2023); Mastropaolo et al.
 (2023); Chen et al.
-(2024a) have primarily evaluated LLM robustness under prompt-level perturbation, focusing on variations in task descriptions.
-While insightful for prompt engineering, these studies overlook robustness in deeper code comprehension tasks, such as reasoning about execution logic and program semantics.
-To determine whether LLMs truly understand program logic, we adopt input and output prediction tasks introduced in CruxEval (CruxGu et al.
-(2024) as diagnostic methods for assessing model in code understanding and execution reasoning.
-Inspired by prior work on stress-testing Natural Language (NL) inference systems Naik et al.
-(2018), we propose a novel stress-testing strategy to evaluate the robustness of LLMs’ code comprehension capabilities under systematic code perturbations.
-First, building upon program structure-consistent mutations Li et al.
+(2024a), which evaluates prompt sensitivity rather than robustness in code reasoning.
+To the best of our knowledge, CodeCrash is the first benchmark to systematically utilize NL-embedded misleading perturbations to evaluate whether LLMs can prioritize executable semantics over NL cues in code reasoning and understanding.
+To diagnose whether LLMs truly understand program logic, we adopt input and output predictions from CruxEval (CruxGu et al.
+(2024) and extend them with LiveCodeBench (LCBJain et al.
+(2025) to cover real-world programs.
+We categorize our NL-embedded perturbations into contextual-level (obviously incorrect cues in multiple formats) and reasoning-level (plausible but incorrect hints inducing rationalization behaviors Chen et al.
+(2025b)), with the goal of stress-testing the critical reasoning capability of LLMs.
+Furthermore, building upon program structure-consistent (PSC) mutations from CCTest Li et al.
+(2023) and other perturbations (e.g., dead-loop poisoning) Wan et al.
+(2022); Sun et al.
+(2023), we construct an aggregated structural perturbation that integrates all traditional structural perturbations for representative comparison.
+We perturb questions in Crux and LCB, and evaluate seventeen LLMs in the direct inference setting, observing significant performance degradation across all designed perturbations.
+We then extend our evaluation to Chain-of-Thought (CoT) Kojima et al.
+(2022) reasoning, and observe the phenomena of distractibility Shi et al.
+(2023) and rationalization Barez et al.
+(2025) under contextual-level and reasoning-level perturbations, respectively, exposing the insufficiency of critical thinking in semantic tracing and code understanding.
+Therefore, we further assess three advanced Large Reasoning Models (LRMs), which achieve superior robustness because of their enhanced internal reasoning.
+However, we find that LRMs are trained to be overly cautious of uncertainty: contextual-level perturbations increase reasoning tokens to digest the obviously misleading cues Kumar et al.
+(2025), while reasoning-level perturbations exploit this bias to trigger pathological overthinking and even “collapse” in QwQ-32B, constituting a novel failure mode.
+In addition, we accidentally identify a potential limitation in input prediction for evaluating code comprehension, where models exploit those misleading cues and bypass the core execution logic without understanding the code.
+Our contributions include:
+  • We introduce a perturbation framework on code reasoning tasks that covers both structural and NL-embedded code transformation for stress-testing LLM code reasoning robustness.
+  • We reveal that LLMs lack sufficient critical reasoning and exploit superficial cues to shortcut reasoning and rationalize comments, extending prior findings on rationalization to code reasoning.
+  • We identify a pathological self-reflection phenomenon in LRMs and novel Reasoning Collapse failure in QwQ-32B, arising from residual rationalization that conflicts with reason-faithful objectives.
+2 Functionally Equivalent Perturbations
+Figure 1: Overview of the CodeCrash pipeline. Perturbed parts are visualized with color highlights.
+2.1 Benchmark Execution Pipeline
+CodeCrash provides a unified perturbation-evaluation pipeline, as illustrated in Figure 1.
+Each code snippet is initially processed via Abstract Syntax Tree (AST) parsing and unparsing to standardize formatting, producing a clean “vanilla” baseline (VAN).
+Subsequently, we apply a selected perturbation and regenerate the code through AST parsing, ensuring that replacement, reformatting, and insertion are precisely controllable.
+Each perturbed code snippet is executed with the provided inputs to double-check the syntactic validity and functional correctness.
+The validated and perturbed code is then provided to LLMs, where we record their generated responses, extract the answers, and execute them to compute the Pass@1 accuracy Chen et al.
+(2021).
+2.2 Perturbation Strategies
+We design a comprehensive perturbation framework that includes structural, contextual-level, and reasoning-level perturbations to systematically stress-test LLMs’ code reasoning reliability.
+Detailed illustrations of each perturbation are provided in Appendix B.
+Aggregated Structural Perturbation (PSC-ALL).
+We adapt PSC mutations from CCTest Li et al.
 (2023) and dead-loop poisoning Wan et al.
 (2022); Sun et al.
-(2023), we design Code Structural Perturbations, which involve logic-preserving transformations, such as renaming entities, reformatting conditional expressions, and inserting unreachable dead code.
-Second, motivated by inconsistent comments in real-world codebases Rong et al.
-(2025), we introduce Textual Distraction Perturbations, which inject misleading NL information into code to challenge the model’s ability to interpret program logic under conflicting cues.
-To extend our evaluation from synthetic programs to real-world codebases, we incorporate LiveCodeBench (LCBJain et al.
-(2025) alongside Crux as part of our benchmark suite.
-We evaluate seventeen LLMs across distinct perturbations in input and output prediction tasks.
-Additionally, we apply Chain-of-Thought (CoT) prompting Kojima et al.
-(2022) to explore how explicit step-by-step reasoning enhances model performance and robustness in code comprehension.
-Furthermore, we assess three state-of-the-art Large Reasoning Models (LRMs) with powerful internal reasoning processes, uncovering significant vulnerabilities arising from their self-reflection mechanisms.
-Our key findings reveal that:
-  • LLMs perform unreliably under perturbations and heavily rely on embedded NL cues for reasoning and comprehension.
-Each textual distraction perturbation causes an average performance drop of over $11$pp 111pp stands for percentage point., while the aggregated structural perturbation also leads to up to $14$pp degradation.
-  • CoT prompting improves model robustness, but cannot fully eliminate sensitivity to NL cues, resulting in a relative degradation of up to $14\%$ under contextual-level textual perturbation.
-  • LRMs show superior robustness under most perturbations, but the self-reflection mechanism consumes 23×2-3$\times$2 - 3 × more reasoning tokens to resolve contradictions, even leading to reflection loops.
-Our contributions include:
-  • We introduce a comprehensive perturbation framework, CodeCrash, to evaluate the code-level robustness of LLMs, which is easily applicable to other benchmarks and programming languages.
-  • We conduct extensive experiments on 17 LLMs across two established datasets, Crux and LCB.
-  • We provide novel insights into evaluating and interpreting code reasoning in LLMs, revealing the reliance on NL cues and demonstrating how reasoning-level hints can lead to model collapse.
-2 Related Work
-Researchers have significantly contributed to developing benchmarks for evaluating LLM capabilities on code-related tasks.
-Early benchmarks such as HumanEval Chen et al.
-(2021), MBPP Austin et al.
-(2021), APPS Hendrycks et al.
-(2021), and HumanEval+ Liu et al.
-(2023b) primarily focused on assessing code generation performance.
-More recent efforts have shifted toward practical software engineering and realistic coding scenarios.
-For instance, SWE-bench Jimenez et al.
-(2024), InfiBench Li et al.
-(2024), and LCB Jain et al.
-(2025) evaluate LLMs on complex tasks like multi-file changes, Stack Overflow-style questions, and real-world programming problems.
-Meanwhile, a growing number of works investigate LLMs in code execution (i.e.
-output prediction).
-Early studies Austin et al.
-(2021); Nye et al.
-(2021) indicated that LLMs face challenges in accurately executing code but can benefit from intermediate reasoning steps.
-Subsequent research has utilized various strategies to enhance execution performance, including execution-informed pretraining Liu et al.
-(2023a), CoT prompting with trace supervision Ni et al.
-(2024), and iterative instruction prompting Lyu et al.
-(2024).
-Crux Gu et al.
-(2024) formalized input and output prediction tasks specifically for evaluating LLMs’ code understanding and reasoning, while LCB extended these assessments to realistic programs written by humans.
-REVAL Chen et al.
-(2024b) and CACP Hooda et al.
-(2024) uncovered inconsistencies and concept-level misunderstandings of LLMs in code through intermediate reasoning analysis in code execution.
-CodeScore Dong et al.
-(2024), Lever Ni et al.
-(2023), and xCodeEval Khan et al.
-(2024) have leveraged code execution to improve model evaluation methodologies and enhance the performance of downstream tasks.
-Robustness evaluation is another research direction in assessing LLMs for code.
-Prior studies focused on prompt-level variations in code generation, examining how semantically equivalent NL descriptions impact model performance Mastropaolo et al.
-(2023); Chen et al.
-(2024a); Zhuo et al.
-(2023).
-Apart from NL perturbations, researchers have investigated model robustness under programming language perturbations.
-For example, CCTest Li et al.
-(2023) introduced structural mutations, CREAM Gao et al.
-(2023) employed counterfactual reasoning, and ReCode Wang et al.
-(2022) applied semantic-preserving transformations, advancing the robustness in code comprehension and completion.
-BigCodeBench Zhuo et al.
-(2025) further explored generalization under complex function composition and tool usage.
-Furthermore, Chen et al.
-(2025); Cuadron et al.
-(2025) and Kumar et al.
-(2025) have recently revealed that LRMs frequently suffer from overthinking and produce unnecessary analysis and unstable thoughts in math-related works.
-3 Methodology
-This section outlines our benchmark methodology, including dataset selection, benchmark pipeline, perturbation strategies, and experimental settings.
-3.1 Benchmark Tasks and Datasets
-We evaluate two benchmarks: Crux Gu et al.
-(2024) and LCB Jain et al.
-(2025).
-Crux contains 800 synthetic problems generated by Code-LLaMA-34B, each featuring concise Python snippets with generic function names f and a corresponding input-output pair for input and output prediction tasks.
-In contrast, LCB contains 479 real-world coding problems sourced from 85 human-submitted LeetCode solutions, focusing solely on output prediction.
-While both datasets avoid excessive memory or runtime demands, LCB problems exhibit more complex structures, such as nested functions, decorators, and external libraries, requiring more computation.
-3.2 Benchmark Execution Pipeline
-Our benchmark follows a unified evaluation pipeline (Figure 3) across all perturbations.
-Each original code snippet is initially processed via Abstract Syntax Tree (AST) parsing and unparsing to remove comments and standardize formatting, producing a clean baseline “vanilla” version (VAN).
-Subsequently, we apply a selected perturbation and regenerate the code using AST unparsing to maintain formatting consistency (except for comment-based perturbations).
-Each perturbed code snippet is executed with the provided inputs to ensure syntactic validity and functional correctness.
-The validated perturbed code is then provided to LLMs, and we record all execution results and evaluate them using the original evaluation methods from the dataset to compute the Pass@1 score.
-3.3 Perturbation Strategies
-We design two categories of perturbations to evaluate model robustness comprehensively: (1) Code Structural Perturbations, which modify code structure without affecting functionality, and (2) Textual Distraction Perturbations, which inject misleading NL cues and hints.
-Detailed examples and principles are provided in Appendix A.
-3.3.1 Code Structural Perturbations
-(1) Renaming Entities (REN).
-[…]
-(2) Reformatting Conditional Expressions (RTF).
-We rewrite conditional expressions within “if” and “while” statements into logically equivalent but structurally complex and indirect forms.
-We use 16 predefined transformation templates to reformat comparison, boolean, and other conditional expressions.
-This perturbation evaluates the LLMs’ ability to maintain logical consistency and reasoning under increased complexity.
-(3) Inserting Garbage Code (GBC).
-We insert syntactically valid but functionally irrelevant “garbage” code segments to challenge model robustness against code structural distractions.
-Three specific types include: (1) repeated global variable declarations mirroring local identifiers, (2) non-executable conditional dead branches, and (3) uninvoked infinite recursive function definitions.
-(4) Aggregating Structural Perturbations (ALL).
-We comprehensively aggregate all structural perturbations (REN, RTF, GBC) together and create a highly adversarial scenario, thoroughly challenging LLM robustness across multiple structural perturbations simultaneously.
-3.3.2 Textual Distraction Perturbations
-(1) Inserting Misleading Descriptive Comments (MDC).
-We insert misleading NL comments for eight key AST nodes (e.g., function definitions, returns, loops, conditionals) to evaluate LLM robustness against textual cues and contradictory information.
-We initially instruct GPT-4o to generate those messages and manually filter them to ensure they are generic, explicitly incorrect, and contradict the actual code logic.
-Additionally, we introduce a probability parameter $p$ to vary comment density and analyze its impact and validity on model robustness in §5.1.
-(2) Inserting Misleading Print Statements (MPS).
-Similar to MDC perturbation, we inject misleading NL messages as executable print statements.
-Unlike comments that AST-based heuristics can easily filter out, print statements are logically irrelevant to functionality and should not be perceived as semantic guides.
-Allowing us to directly compare the behavior with MDC and distinguish models’ NL reliance in different embedding formats.
-Since MDC and MPS perturbations inject shallow and contextually misleading information, we categorize them as contextual-level textual perturbations.
-(3) Inserting Misleading Hint Comments (MHC).
-We instruct GPT-4o to generate logically plausible but incorrect hints that contradict the actual program behavior and the ground truth.
-After verifying their incorrectness, they are strategically inserted into function definitions or return statements using AST parsing.
-Since these plausible hints challenge the model’s high-level inference ability, we classify MHC as a reasoning-level textual perturbation.
-4 Experiments
-This section presents a comprehensive analysis of model robustness under the perturbation strategies introduced in §3.3.
-We assess the reliability of LLMs across different perturbation types, model scales, reasoning modes, and prediction tasks.
+(2023) by aggregating them on the identifier-level (i.e., Renaming Entities, REN), instruction-level (i.e., Reformatting Conditional Expressions, RTF), and block-level (i.e., Garbage Code, GBC; including unexecutable dead loops, dead blocks, and global variables).
+PSC-ALL constructs a more complicated but functionally equivalent program and represents the traditional structural perturbation that modifies syntax to expose the pattern-matching biases of LLMs.
+For completeness, we specify REN, RTF, and GBC in Appendix B.1.1-B.1.3, and highlight their compounding effects in Appendix B.1.
+Since their impact is consistently weaker than PSC-ALL, we focus our experiments on PSC-ALL.
+Contextual-Level Misleading Perturbations (MCC & MPS).
+These perturbations aim to mislead models by injecting shallow NL cues into the surface context without altering the code logic.
+We target eight critical and ordinary AST nodes (e.g., function definitions, returns, and loops) and leverage GPT-4o to generate the misleading messages, which we manually filter to ensure they are generic, explicitly incorrect, and contradictory to the code logic.
+Concretely, we design two variants: (1) Misleading Code Comments (MCC), where we inject messages as code comments; and (2) Misleading Print Statements (MPS), where we embed messages via executable print statements.
+They evaluate whether LLMs naively include the irrelevant NL cues from the program for reasoning, using different injection formats to clarify that the phenomenon is not bound to a specific format.
+Reasoning-Level Misleading Perturbation (MHC).
+We further design the Misleading Hint Comments (MHC) perturbation, which differs from contextual-level perturbations by providing high-level incorrect hints about program outputs.
+We instruct GPT-4o to generate a plausible but incorrect answer that preserves the expected type and structure but contradicts the ground truth.
+After verifying the incorrectness, we utilize AST parsing to strategically inject it as a comment at function definitions or return statements to make it more confusing.
+Unlike shallow contradictions from contextual perturbations, MHC targets the reasoning process, assessing whether the model can critically evaluate conflicting information rather than rationalize the hint and shortcut its reasoning.
+3 Experimental Results and Analysis
 3.1 Experiment Settings
+Table 1: Relative Pass@1 degradation of LLMs under all perturbation types for output prediction tasks on Crux and LCB using direct inference.
+Darker red highlights represent more severe degradation.
+See [Absolute Version] for the corresponding absolute degradation.
+Model Series Model Name VAN PSC-ALL MCC MPS MHC Average
+Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
+GPT-4 Omni GPT-4o-Mini $57.3$ $53.2$ $27.9\%$ $31.9\%$ $32.6\%$ $40.4\%$ $23.2\%$ $28.0\%$ $11.6\%$ $44.3\%$ $28.4\%$
+GPT-4o $71.3$ $64.5$ $15.0\%$ $28.4\%$ $14.0\%$ $29.1\%$ $17.2\%$ $24.3\%$ $6.4\%$ $26.6\%$ $18.4\%$
+Claude Series Claude-3.5-Haiku-20241022 $57.4$ $58.2$ $24.0\%$ $36.9\%$ $13.8\%$ $0.0\% 11$$1.2\%$ $8.0\%$ $27.5\%$ $53.3\%$ $22.1\%$
+Claude-3.5-Sonnet-20241022 $71.5$ $73.8$ $14.8\%$ $34.1\%$ $8.1\%$ $9.6\%$ $8.6\%$ $10.7\%$ $14.4\%$ $43.4\%$ $16.3\%$
+Gemini Series Gemini-1.5-Flash $56.2$ $44.3$ $18.4\%$ $12.7\%$ $14.8\%$ $32.2\%$ $28.5\%$ $40.4\%$ $7.3\%$ $22.2\%$ $20.8\%$
+Gemini-2.0-Flash $65.0$ $66.0$ $33.2\%$ $42.1\%$ $26.8\%$ $51.9\%$ $33.1\%$ $50.3\%$ $8.3\%$ $19.5\%$ $31.2\%$
+Gemini-1.5-Pro-002 $67.4$ $56.2$ $19.9\%$ $23.2\%$ $23.0\%$ $33.0\%$ $21.7\%$ $32.8\%$ $10.0\%$ $32.5\%$ $23.0\%$
+DeepSeek DeepSeek-V3 $67.9$ $67.8$ $12.9\%$ $35.6\%$ $16.6\%$ $41.4\%$ $10.1\%$ $34.2\%$ $10.7\%$ $29.7\%$ $21.1\%$
+LLaMA Series LLaMA-3.1-8B-Instruct $36.0$ $34.7$ $23.8\%$ $19.7\%$ $22.5\%$ $28.5\%$ $17.8\%$ $23.1\%$ $16.7\%$ $39.2\%$ $23.0\%$
+LLaMA-3.1-70B-Instruct $56.1$ $44.9$ $19.6\%$ $17.2\%$ $19.2\%$ $27.9\%$ $26.7\%$ $38.8\%$ $8.4\%$ $32.2\%$ $22.4\%$
+LLaMA-3.1-405B-Instruct $63.5$ $50.7$ $19.7\%$ $16.9\%$ $6.6\%$ $4.7\% 11$$1.3\%$ $19.9\%$ $6.9\%$ $25.7\%$ $14.2\%$
+LLaMA-3.3-70B-Instruct $59.9$ $48.5$ $17.3\%$ $20.1\%$ $12.4\%$ $19.8\%$ $13.5\%$ $25.4\%$ $6.8\%$ $20.9\%$ $15.9\%$
+Qwen Series Qwen2.5-7B-Instruct $43.3$ $41.4$ $37.9\%$ $30.9\%$ $58.0\%$ $38.3\%$ $45.2\%$ $19.8\%$ $26.9\%$ $55.6\%$ $39.8\%$
+Qwen2.5-14B-Instruct $47.8$ $49.5$ $39.8\%$ $30.0\%$ $34.9\%$ $41.4\%$ $2.5\% 33$$\% 9.3-9.3\%$ $22.2\%$ $30.2\%$
+Qwen2.5-32B-Instruct $60.0$ $59.6$ $19.8\%$ $34.9\%$ $18.0\%$ $\% 23.7-23.7\%$ $23.6\%$ $13.1\%$ $30.8\%$ $23.1\%$
+Qwen2.5-72B-Instruct $60.1$ $54.9$ $23.3\%$ $25.2\%$ $17.0\%$ $12.7\%$ $24.1\%$ $26.2\%$ $16.0\%$ $40.6\%$ $22.4\%$
+Qwen2.5-Coder-32B-Instruct $67.0$ $56.6$ $22.3\%$ $27.9\%$ $23.2\%$ $36.6\%$ $16.6\%$ $30.3\%$ $10.5\%$ $1.7\% 22$$2.3\%$
+Average $22.9\%$ $27.5\%$ $21.3\%$ $29.4\%$ $21.5\%$ $27.6\%$ $12.4\%$ $33.0\%$
+$\mathbf{-24.6}\%$ $\mathbf{-24.3}\%$ $\mathbf{-23.8}\%$ $\mathbf{-20.1}\%$ $\mathbf{-23.2}\%$
 Model Selection.
@@
 (2024).
-Additionally, we assess three cutting-edge LRMs: o3-mini OpenAI (2025), DeepSeek-R1 DeepSeek-AI et al.
+Additionally, we assess 3 cutting-edge LRMs: o3-mini OpenAI (2025), DeepSeek-R1 DeepSeek-AI et al.
 (2025a), and QwQ-32B Qwen Team (2025).
 Model Configurations.
-Following LCB, we use nucleus sampling (temperature=0.2, top-p=0.95), with a maximum of 200 tokens for direct inference and 2,000 for CoT prompting, discarding excessively long outputs.
-For o3-mini, we raise the limit to 20,000 tokens to avoid verbose reasoning.
-Due to resource constraints, we generate three candidates for direct inference and one for CoT inference per query; we provide a result stability analysis under $N=3$ in Appendix B.
-Following Crux, we employ 2-shot prompting for direct inference and 1-shot for CoT inference (details provided in Appendix C).
-Furthermore, we adopt Pass@1 accuracy Chen et al.
-(2021) as the primary evaluation metric and scale all scores in $[0,1]$ to percentages by multiplying by $100$ for readability.
-All perturbed results are reported as either absolute ($\Delta_{\text{\;{pp}}}=\text{Perturbed}-\text{VAN}$) or relative ($\Delta_{\%}=\frac{\text{Perturbed}-\text{VAN}}{\text{VAN}}\times 100\%$) differences from the corresponding vanilla baseline.
-Additionally, we use percent signs (%) to indicate relative differences for clarity in all tables.
-4.2 Impact on Output Prediction
-We first evaluate LLM robustness in output prediction tasks.
-Table 1 summarizes the absolute Pass@1 degradation for all models under direct inference across both datasets.
-Our findings in this subsection include:
-Key Findings: LLMs demonstrate an inherent and multi-layered reliance on natural language information, including identifiers, contextual messages, and reasoning hints, which ultimately undermines their reliability in code reasoning and comprehension.
-LLMs exhibit limited robustness to structurally disorganized code, struggling to differentiate irrelevant code from essential logic.
-Aggregating multiple structural perturbations exacerbates this vulnerability, significantly impairing their reasoning reliability.
-Larger scales, newer versions, and domain-specific fine-tuning LLMs substantially enhance robustness, highlighting improved code understanding and reasoning capabilities.
-Table 1: Absolute Pass@1 degradation of LLMs under all perturbation types for output prediction tasks on Crux and LCB using direct inference ($N=3$).
-VAN shows the Pass@1 accuracy (in %), others indicate absolute drops from VAN (in  pp).
-Darker red highlights represent more severe degradation.
-See [To Relative Version] for the relative version.
-Model Series Model Name VAN REN RTF GBC ALL MDC MPS MHC
-Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
-GPT-4 Omni GPT-4o-Mini $57.29$ $53.24$ $-0.62$ $-9.67$ $-1.79$ $-0.97$ $-10.92$ $-5.50$ $-15.96$ $-16.98$ $-18.67$ $-21.50$ $-13.29$ $-14.89$ $-6.62$ $-23.59$
-GPT-4o $71.33$ $64.51$ $-2.71$ $-9.60$ $-3.38$ $-4.59$ $-9.21$ $-3.20$ $-10.71$ $-18.30$ $-9.96$ $-18.79$ $-12.29$ $-15.66$ $-4.54$ $-17.19$
-Claude Series Claude-3.5-Haiku-20241022 $57.42$ $58.25$ $-3.96$ $-9.32$ $-1.79$ $+0.35$ $-7.29$ $-4.52$ $-13.79$ $-21.50$ $-7.92$ $-5.85$ $-6.46$ $-4.66$ $-15.79$ $-31.04$
-Claude-3.5-Sonnet-20241022 $71.46$ $73.83$ $-1.67$ $-16.77$ $-0.92$ $+1.18$ $-6.88$ $-4.94$ $-10.54$ $-25.19$ $-5.79$ $-7.10$ $-6.17$ $-7.86$ $-10.29$ $-32.08$
-Gemini Series Gemini-1.5-Flash $56.21$ $44.26$ $-3.54$ $-0.42$ $-1.17$ $-1.04$ $-4.58$ $-0.35$ $-10.33$ $-5.64$ $-8.29$ $-14.27$ $-16.00$ $-17.88$ $-4.12$ $-9.81$
-Gemini-2.0-Flash $65.04$ $65.97$ $-2.33$ $-18.09$ $-2.08$ $-0.77$ $-15.50$ $-8.77$ $-21.63$ $-27.77$ $-17.42$ $-34.24$ $-21.54$ $-33.19$ $-5.38$ $-12.87$
-Gemini-1.5-Pro-002 $67.37$ $56.16$ $-2.92$ $-3.41$ $-2.50$ $-4.73$ $-5.62$ $-3.83$ $-13.42$ $-13.01$ $-15.50$ $-18.51$ $-14.62$ $-18.44$ $-6.75$ $-18.23$
-DeepSeek DeepSeek-V3 $67.87$ $67.78$ $-2.67$ $-16.35$ $-1.33$ $-3.76$ $-4.62$ $-5.98$ $-8.75$ $-24.15$ $-11.25$ $-28.04$ $-6.87$ $-23.17$ $-7.29$ $-20.11$
-LLaMA Series LLaMA-3.1-8B-Instruct $35.96$ $34.66$ $+0.08$ $-4.31$ $-0.33$ $-1.67$ $-3.13$ $-4.73$ $-8.54$ $-6.82$ $-8.08$ $-9.88$ $-6.42$ $-8.00$ $-6.00$ $-13.57$
-LLaMA-3.1-70B-Instruct $56.08$ $44.89$ $-2.38$ $-1.39$ $-0.54$ $-1.74$ $-5.38$ $-2.44$ $-11.00$ $-7.72$ $-10.79$ $-12.53$ $-14.96$ $-17.40$ $-4.71$ $-14.47$
-LLaMA-3.1-405B-Instruct $63.54$ $50.66$ $-5.37$ $-3.90$ $-0.62$ $-4.45$ $-5.79$ $-2.85$ $-12.50$ $-8.56$ $-4.17$ $-7.45$ $-7.17$ $-10.09$ $-4.37$ $-13.01$
-LLaMA-3.3-70B-Instruct $59.92$ $48.50$ $-2.21$ $-1.39$ $-1.25$ $-3.41$ $-5.50$ $-5.71$ $-10.37$ $-9.74$ $-7.46$ $-9.60$ $-8.08$ $-12.32$ $-4.08$ $-10.16$
-Qwen Series Qwen2.5-7B-Instruct $43.29$ $41.41$ $-3.04$ $-7.24$ $-1.29$ $-3.20$ $-12.96$ $-3.76$ $-16.42$ $-12.80$ $-25.13$ $-15.87$ $-19.58$ $-8.21$ $-11.67$ $-23.03$
-Qwen2.5-14B-Instruct $47.79$ $49.48$ $-1.83$ $-9.88$ $-0.67$ $-1.39$ $-11.21$ $-3.41$ $-19.04$ $-14.82$ $-16.67$ $-20.46$ $-15.54$ $-16.77$ $-4.46$ $-11.00$
-Qwen2.5-32B-Instruct $60.04$ $59.64$ $-1.71$ $-10.79$ $-2.58$ $-0.49$ $-5.79$ $-1.39$ $-11.88$ $-20.81$ $-10.83$ $-19.49$ $-14.21$ $-14.06$ $-7.87$ $-18.37$
-Qwen2.5-72B-Instruct $60.08$ $54.91$ $-1.04$ $-8.49$ $-1.87$ $+0.42$ $-10.71$ $-2.02$ $-14.00$ $-13.85$ $-10.21$ $-6.96$ $-14.46$ $-14.41$ $-9.63$ $-22.27$
-Qwen2.5-Coder-32B-Instruct $66.96$ $56.65$ $-2.50$ $-8.42$ $-0.96$ $-2.02$ $-10.75$ $-2.78$ $-14.96$ $-15.80$ $-15.54$ $-20.74$ $-11.08$ $-17.19$ $-7.00$ $-12.32$
-Average $-2.38$ $-8.20$ $-1.47$ $-1.90$ $-7.99$ $-3.89$ $-13.17$ $-15.50$ $-11.98$ $-15.96$ $-12.28$ $-14.95$ $-7.09$ $-17.83$
-$\mathbf{-4.56}$ $\mathbf{-1.63}$ $\mathbf{-6.46}$ $\mathbf{-14.04}$ $\mathbf{-13.47}$ $\mathbf{-13.28}$ $\mathbf{-11.11}$
-4.2.1 Over-reliance on Natural Language Semantics
-We categorize NL dependence into three hierarchical levels: surface-level, contextual-level, and reasoning-level dependence.
-Surface-level NL Dependence (REN).
-Under the REN perturbation, models suffer more significant performance degradation on LCB ($8.20$pp on average) compared to Crux ($2.38$pp) due to the presence of meaningful function names in the vanilla settings of LCB, which are absent in Crux.
-High-performing models like Claude-3.5-Sonnet and DeepSeek-V3 suffer severe degradations exceeding $\mathbf{16}$pp on LCB, highlighting a significant reliance on semantically meaningful identifiers.
-Contextual-level NL Dependence (MDC and MPS).
-Under contextual-level textual perturbations (MDC and MPS), LLMs exhibit severe performance drops ($13.47$pp and $13.28$pp, respectively).
-Notably, we identify two prominent consistency patterns: (1) dataset-level consistency [^2] If MDC degrades performance more on Crux than LCB, the same holds for MPS, and vice versa.
-($16/17$ models), implying the model sensitivity to misleading comments and prints primarily depends on dataset characteristics and (2) perturbation-method consistency [^3] If MDC causes greater degradation on Crux than LCB, so does MPS, and vice versa.
-($14/17$ models), indicating an inherent cognitive preference for NL messages from specific formats.
-These phenomena reveal that LLMs inherently depend on embedded NL context and exhibit format-specific cognitive biases.
-Reasoning-level NL Dependence (MHC).
-MHC perturbation also induces substantial degradation ($11.11$pp on average), notably exacerbated in complex datasets (LCB: $17.83$pp vs.
-Crux: $7.09$pp), which is a complexity-amplified effect.
-Advanced models like Claude-3.5-Haiku and Sonnet experience over $30$pp performance degradation, far exceeding their degradation under surface or contextual perturbations.
-These results illustrate that (i) LLMs readily accept plausible hint comments as authoritative reasoning shortcuts, and (ii) this over-trust intensifies when task complexity surpasses the model’s reasoning capacity, which we revisit under CoT prompting in §4.3.1.
-4.2.2 Low Resistance to Disorganized Coding Environments
-From RTF perturbation, LLMs can effectively reason over logically equivalent but syntactically complex expressions with minimal degradation ($1.63$pp).
-However, the highest performance drop emerges under the GBC compared to REN and RTF, with an average degradation of $6.46$pp across both datasets.
-Notably, GBC is the only perturbation where degradation is more severe on LCB ($3.89$pp vs.
-$7.99$pp on Crux) due to uneliminated meaningful identifiers.
-This trend reverses under the ALL perturbation, where all structural perturbations are combined, leading to a greater performance drop on LCB ($15.50$pp) than on Crux ($13.17$pp).
-Additionally, we observe a compounding effect of structural perturbations under ALL perturbation, resulting in the most significant overall degradation ($14.04$pp), which surpasses any individual code structural perturbation effects.
-4.2.3 Model Robustness Across Different Scales
-We further analyze reasoning reliability across different model scales and versions, comparing relative Pass@1 degradation under representative perturbations in Table 2.
-Due to length constraints, the table, weighting scheme, and additional evaluation setups are provided in Appendix D.2.
-We observe a clear and consistent trend except for Gemini families.
-Larger closed-sourced models from GPT-4 Omni and Claude families demonstrate enhanced robustness compared to their smaller counterparts.
-Similar trends appear in open-source models, such as LLaMA (3.1-8B $<$ 70B $<$ 405B) and Qwen (2.5-7B $<$ 14B $<$ 32B $<$ 72B).
-Additionally, newer model versions (LLaMA-3.3-70B) and domain-specific fine-tuned variants (Qwen2.5-Coder-32B) consistently exhibit improved robustness compared to their predecessors (LLaMA-3.1-70B) and larger counterparts (Qwen2.5-72B), respectively.
-Table 2: Relative Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/2 ALL, 1/8 MDC, 1/8 MPS, and 1/4 MHC. [To Absolute Version]
-Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average
-GPT-4 Omni GPT-4o-Mini 55.77 $-29.37\%$ $-35.51\%$ $-24.99\%$ $-23.83\%$ $-28.20\%$
-GPT-4o 68.78 $-20.01\%$ $-19.64\%$ $-19.87\%$ $-13.96\%$ $-18.44\%$
-Claude Series Claude-3.5-Haiku-20241022 57.73 $-28.85\%$ $-12.38\%$ $-10.03\%$ $-37.16\%$ $-26.52\%$
-Claude-3.5-Sonnet-20241022 72.35 $-22.01\%$ $-8.67\%$ $-9.39\%$ $-25.28\%$ $-19.58\%$
-Gemini Series Gemini-1.5-Flash 51.73 $-16.27\%$ $-21.30\%$ $-32.94\%$ $-12.89\%$ $-18.14\%$
-Gemini-2.0-Flash 65.39 $-36.56\%$ $-36.19\%$ $-39.56\%$ $-12.48\%$ $-30.87\%$
-Gemini-1.5-Pro-002 63.17 $-21.13\%$ $-26.73\%$ $-25.88\%$ $-18.43\%$ $-21.75\%$
-DeepSeek DeepSeek-V3 67.84 $-21.41\%$ $-25.86\%$ $-19.14\%$ $-17.83\%$ $-20.79\%$
-LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-22.23\%$ $-24.74\%$ $-19.81\%$ $-25.10\%$ $-22.96\%$
-LLaMA-3.1-70B-Instruct 51.89 $-18.71\%$ $-22.49\%$ $-31.20\%$ $-17.33\%$ $-20.40\%$
-LLaMA-3.1-405B-Instruct 58.72 $-18.63\%$ $-9.61\%$ $-14.51\%$ $-13.93\%$ $-15.81\%$
-LLaMA-3.3-70B-Instruct 55.64 $-18.35\%$ $-15.20\%$ $-17.95\%$ $-12.11\%$ $-16.35\%$
-Qwen Series Qwen2.5-7B-Instruct 42.59 $-35.30\%$ $-50.65\%$ $-35.72\%$ $-37.69\%$ $-37.87\%$
-Qwen2.5-14B-Instruct 48.42 $-36.14\%$ $-37.30\%$ $-33.04\%$ $-14.16\%$ $-30.40\%$
-Qwen2.5-32B-Instruct 59.89 $-25.44\%$ $-23.52\%$ $-23.63\%$ $-19.74\%$ $-23.55\%$
-Qwen2.5-72B-Instruct 58.14 $-24.02\%$ $-15.37\%$ $-24.88\%$ $-25.21\%$ $-23.34\%$
-Qwen2.5-Coder-32B-Instruct 63.10 $-24.42\%$ $-28.23\%$ $-21.72\%$ $-14.68\%$ $-22.12\%$
-4.3 Impact of Reasoning on Model Robustness
-This section examines how the reasoning process, including CoT prompting for LLMs and internal reasoning in LRMs, affects model robustness against perturbations.
-4.3.1 Effect of CoT Prompting
-Table 3: Comparison of relative Pass@1 degradation aggregated over Crux and LCB for output prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Absolute Version]
-Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average
+Following LCB, we use nucleus sampling (temperature=0.2, top-p=0.95) with a maximum of 200 tokens for direct inference and 2,000 for CoT prompting, discarding excessively long outputs.
+Due to resource constraints, we generate three candidates for direct inference and one for CoT inference; we provide a result stability analysis under $N=3$ in Appendix C.
+Following Crux, we employ 2-shot prompting for direct inference and 1-shot for CoT step-by-step execution (details provided in Appendix D).
+Furthermore, we adopt Pass@1 accuracy as the primary evaluation metric and scale all scores in $[0,1]$ to percentages by multiplying by $100$ for readability.
+The experimental results are reported using relative differences from the corresponding vanilla baseline ($\Delta_{\%}=\frac{\text{Perturbed}-\text{{VAN}}}{\text{{VAN}}}\times 100\%$).
+We also report the absolute differences for completeness in the Appendix E.
+3.2 Severe Degradation under Perturbations
+We begin by evaluating models in the direct inference setting to measure their robustness against perturbations in §2.2 and summarize the results in Table 1.
+On average, models suffer a $\mathbf{23.2\%}$ performance drop, with the aggregated structural perturbation (PSC-ALL) causing the largest degradation ($24.6\%$).
+This indicates that LLMs exhibit limited code-tracing robustness, struggling to identify important code from essential logic and maintain reliability.
+LLMs are sensitive to embedded NL cues.
+As shown in Table 1, contextual-level perturbations (MCC and MPS) degrade performance by about $24\%$ on average, comparable to PSC-ALL, highlighting their substantial impact on model robustness.
+This also reveals that models are unable to consistently prioritize program logic over misleading NL cues, instead naively incorporating those cues into their reasoning.
+Notably, we identify two prominent consistency patterns: (1) dataset-level consistency [^1] If MCC degrades performance more on Crux than LCB, the same holds for MPS, and vice versa.
+($15/17$ models), implying the impact intensity and sensitivity of both injection formats primarily depend on the dataset characteristics; and (2) perturbation-method consistency [^2] If MCC causes greater degradation on Crux than LCB, so does MPS, and vice versa.
+($14/17$ models), indicating format-specific cognitive preference and biases.
+These phenomena suggest that LLMs are inherently impacted by embedded NL context, regardless of the injection format.
+LLMs rationalize the hint to shortcut their reasoning.
+Moving forward to the reasoning-level perturbation, MHC induces significant degradation with a complexity-amplified effect: models average drop $\mathbf{33.0\%}$ on LCB compared to $\mathbf{12.4\%}$ on Crux.
+Since LCB consists of more complex algorithmic problems with relatively simple output structures, models are more willing to adopt the hint as a reasoning shortcut.
+Notably, the performances of Claude-3.5-Haiku and Sonnet drop over $\mathbf{40\%}$ under MHC perturbation on LCB, which aligns with recent findings from Anthropic that models often accept a biased multiple-choice hint and rationalize it rather than critically resolving the contradiction Chen et al.
+(2025b).
+We extend this observation to the code reasoning domain and illustrate that: (i) LLMs readily rationalize the plausible hint as authoritative reasoning shortcuts, and (ii) this over-trust amplifies when task complexity surpasses the model’s reasoning capacity.
+Scaling and versioning systematically improve robustness, with notable exceptions.
+Overall, robustness against perturbations improves consistently with model scale and newer releases.
+We observe a clear and consistent trend that larger closed-sourced models from GPT-4 Omni and Claude families demonstrate enhanced robustness compared to their smaller counterparts.
+Similar trends appear in open-source models, such as LLaMA (3.1-8B $\leq$ 70B << 405B) and Qwen (2.5-7B << 14B << 32B $\leq$ 72B).
+Additionally, newer versions (LLaMA-3.3-70B) and domain-specific fine-tuned variants (Qwen2.5-Coder-32B) exhibit improved robustness compared to their predecessors (LLaMA-3.1-70B) and larger counterparts (Qwen2.5-72B), respectively.
+Interestingly, the Gemini series deviates from this trend, underscoring that architectures and training strategies could affect robustness.
+3.3 Limits of CoT in Resolving NL-Embedded Contradictions
+Table 2: Comparison of relative Pass@1 degradation aggregated over Crux and LCB for output prediction in different reasoning modes. See [Absolute Version] for the absolute degradation.
+Model Series Model Name VAN PSC-ALL MCC MPS MHC Average
 Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT
-GPT-4 Omni GPT-4o-Mini $55.77$ $81.47$ $-29.30\%$ $-13.15\%$ $-35.37\%$ $-10.84\%$ $-24.91\%$ $-10.94\%$ $-23.27\%$ $-2.69\%$ $-28.00\%$ $-9.52\%$
-GPT-4o $68.78$ $91.79$ $-19.70\%$ $-4.94\%$ $-19.29\%$ $-5.03\%$ $-19.70\%$ $-5.62\%$ $-13.49\%$ $-1.36\%$ $-18.10\%$ $-4.22\%$
-Claude Series Claude-3.5-Haiku-20241022 $57.73$ $72.95$ $-28.89\%$ $-21.33\%$ $-12.37\%$ $-10.61\%$ $-10.02\%$ $-8.68\%$ $-37.25\%$ $-14.26\%$ $-26.56\%$ $-14.30\%$
-Claude-3.5-Sonnet-20241022 $72.35$ $86.00$ $-22.15\%$ $-7.55\%$ $-8.68\%$ $-5.27\%$ $-9.40\%$ $-6.36\%$ $-25.50\%$ $-7.91\%$ $-19.71\%$ $-6.86\%$
-Gemini Series Gemini-1.5-Flash $51.73$ $75.22$ $-16.57\%$ $-18.30\%$ $-20.35\%$ $-21.73\%$ $-32.29\%$ $-42.20\%$ $-12.09\%$ $-2.08\%$ $-17.89\%$ $-20.39\%$
-Gemini-2.0-Flash $65.39$ $89.13$ $-36.59\%$ $-6.23\%$ $-36.27\%$ $-6.32\%$ $-39.62\%$ $-14.30\%$ $-12.51\%$ $-2.02\%$ $-30.91\%$ $-7.01\%$
-Gemini-1.5-Pro-002 $63.17$ $87.18$ $-21.00\%$ $-7.44\%$ $-26.32\%$ $-11.93\%$ $-25.41\%$ $-14.71\%$ $-17.49\%$ $-3.86\%$ $-21.34\%$ $-9.19\%$
-DeepSeek DeepSeek-V3 $67.84$ $89.52$ $-21.40\%$ $-9.96\%$ $-25.85\%$ $-17.73\%$ $-19.13\%$ $-18.78\%$ $-17.83\%$ $-3.49\%$ $-20.78\%$ $-12.07\%$
-LLaMA Series LLaMA-3.1-8B-Instruct $35.47$ $44.72$ $-22.26\%$ $-27.62\%$ $-24.69\%$ $-20.98\%$ $-19.76\%$ $-21.50\%$ $-24.91\%$ $-8.74\%$ $-22.92\%$ $-20.03\%$
-LLaMA-3.1-70B-Instruct $51.89$ $69.43$ $-18.83\%$ $-17.91\%$ $-22.05\%$ $-23.20\%$ $-30.59\%$ $-30.86\%$ $-16.12\%$ $-6.76\%$ $-20.03\%$ $-19.22\%$
-LLaMA-3.1-405B-Instruct $58.72$ $78.42$ $-18.77\%$ $-13.56\%$ $-9.19\%$ $-10.27\%$ $-14.07\%$ $-15.55\%$ $-12.96\%$ $-7.48\%$ $-15.53\%$ $-11.75\%$
-LLaMA-3.3-70B-Instruct $55.64$ $76.86$ $-18.22\%$ $-10.89\%$ $-14.85\%$ $-7.73\%$ $-17.38\%$ $-11.09\%$ $-11.43\%$ $-4.78\%$ $-16.00\%$ $-8.70\%$
-Qwen Series Qwen2.5-7B-Instruct $42.59$ $58.17$ $-35.37\%$ $-21.77\%$ $-50.86\%$ $-21.91\%$ $-35.99\%$ $-27.96\%$ $-37.39\%$ $-6.99\%$ $-37.89\%$ $-19.50\%$
-Qwen2.5-14B-Instruct $48.42$ $70.68$ $-36.06\%$ $-23.12\%$ $-37.35\%$ $-35.29\%$ $-33.05\%$ $-43.36\%$ $-14.26\%$ $-10.84\%$ $-30.40\%$ $-27.34\%$
-Qwen2.5-32B-Instruct $59.89$ $79.44$ $-25.41\%$ $-17.42\%$ $-23.50\%$ $-18.60\%$ $-23.63\%$ $-24.21\%$ $-19.71\%$ $-5.31\%$ $-23.53\%$ $-16.19\%$
-Qwen2.5-72B-Instruct $58.14$ $82.80$ $-23.98\%$ $-12.65\%$ $-15.46\%$ $-6.80\%$ $-24.83\%$ $-16.24\%$ $-24.70\%$ $-6.61\%$ $-23.20\%$ $-10.63\%$
-Qwen2.5-Coder-32B-Instruct $63.10$ $84.75$ $-24.20\%$ $-14.67\%$ $-27.72\%$ $-19.00\%$ $-21.19\%$ $-19.93\%$ $-14.25\%$ $-9.32\%$ $-21.78\%$ $-15.49\%$
-Average $\mathbf{-24.63}\%$ $\mathbf{-14.62}\%$ $\mathbf{-24.13}\%$ $\mathbf{-14.90}\%$ $\mathbf{-23.59}\%$ $\mathbf{-19.55}\%$ $\mathbf{-19.72}\%$ $\mathbf{-6.15}\%$ $\mathbf{-23.21}\%$ $\mathbf{-13.67}\%$
-We examine how CoT prompting improves LLM robustness through explicit step-by-step execution.
-From Table 3, we observe that CoT effectively enhances model robustness, with relative improvements of $+10.01$pp for ALL, $+9.23$pp for MDC, $+4.04$pp for MPS, and $+13.57$pp for MHC.
-The results indicate that step-by-step reasoning guides LLMs to prioritize executable semantics over irrelevant code segments and textual cues.
-Notably, the substantial improvement in MHC suggests CoT effectively mitigates the over-reliance on reasoning-level hints by re-centering reasoning around executable semantics.
-Such benefits are particularly evident in high-capacity models (e.g., GPT-4o, Claude-3.5-Sonnet), reflecting enhanced reasoning capabilities and greater noise resilience.
-However, CoT prompting cannot entirely eliminate the model’s sensitivity to misleading NL cues and structural distractions.
-Substantial performance degradation remains under ALL, MDC, and MPS perturbations, particularly for smaller open-sourced models (e.g., Qwen2.5 and LLaMA-3.1).
-These results underscore an intrinsic limitation of current models in distinguishing non-functional code and misleading information from essential logic.
-Interestingly, models usually incorporate print statements directly into their reasoning paths, leading to greater degradation under MPS compared to MDC due to self-induced misinterpretation.
-Since CoT prompting exposes intermediate reasoning processes, we provide detailed case studies for each perturbation in Appendix E (Case 1 to 4) to investigate model inference and illustrate their failure reasons.
-Key Finding: CoT prompting significantly improves LLM robustness by enforcing explicit step-by-step reasoning, particularly benefiting larger models.
-Nevertheless, it cannot fully eliminate models’ reliance on misleading cues, indicating fundamental limitations in reasoning capabilities.
-4.3.2 Powerful Reasoning and Potential Vulnerability of LRMs
-We then evaluate the reliability of three LRMs on output prediction tasks, with results summarized in Table 4.
-Compared to LLMs, LRMs show outstanding performance and robustness, surpassing $90$ Pass@1 in vanilla settings and maintaining minimal performance drop under ALL and MPS perturbations.
-These results indicate that internal reasoning outperforms CoT prompting in accurately distinguishing essential executable code segments from irrelevant code and misleading distractions.
-Notably, LRMs consistently exhibit stronger resistance to MPS compared to MDC, suggesting an inherent cognitive bias toward treating comments as authoritative execution guidance while appropriately disregarding print content.
-However, this residual cognitive bias exposes LRMs to critical vulnerabilities under reasoning-level perturbation that plausible but incorrect hints can trigger severe reasoning failures.
-Specifically, we observe the following outcomes under MHC: (i) substantial increases in reasoning token usage across all LRMs, often doubling when compared to VAN; (ii) extreme cases in QwQ-32B—occurring in $4/479$ tasks; see Appendix E.6—where it generates over 32,000 tokens due to uncontrolled recursive self-verification and reasoning collapse; (iii) increased reasoning effort and token usage instead exacerbate performance drops in o3-Mini, suggesting that internal reasoning can amplify confusion when led by contradiction.
-These findings reveal that reasoning-level contradictories can evoke severe internal reasoning instability, resulting in significant computational overhead, excessive self-reflection, and even severe performance drops.
-Table 4: Absolute Pass@1 degradation and reasoning token usage (Avg. and Max.) of LRMs on output prediction across Crux and LCB ($N=1$).
-Model Dataset VAN ALL MDC MPS MHC
+GPT-4 Omni GPT-4o-Mini $55.8$ $81.5$ $29.4\%$ $13.1\%$ $35.5\%$ $10.8\%$ $25.0\%$ $11.0\%$ $23.8\%$ $2.7\%$ $28.4\%$ $9.4\%$
+GPT-4o $68.8$ $91.8$ $20.0\%$ $4.9\%$ $19.6\%$ $5.0\%$ $19.9\%$ $5.6\%$ $0\% 1.4-1.4\%$ $18.4\%$ $4.2\%$
+Claude Series Claude-3.5-Haiku-20241022 $57.7$ $72.9$ $28.9\%$ $21.2\%$ $12.4\%$ $6\% 10.0-1$$0\%$ $8.7\%$ $37.2\%$ $14.2\%$ $22.1\%$ $13.7\%$
+Claude-3.5-Sonnet-20241022 $72.3$ $86.0$ $22.0\%$ $7.4\%$ $8.7\%$ $5.3\%$ $9.4\%$ $6.3\%$ $25.3\%$ $7.8\%$ $16.3\%$ $6.7\%$
+Gemini Series Gemini-1.5-Flash $51.7$ $75.2$ $16.3\%$ $18.3\%$ $21.3\%$ $21.6\%$ $32.9\%$ $42.1\%$ $9\% 2.1-2.1\%$ $20.8\%$ $21.0\%$
+Gemini-2.0-Flash $65.4$ $89.1$ $36.6\%$ $6.2\%$ $2\% 6.3-6.3\%$ $39.6\%$ $14.1\%$ $12.5\%$ $2.0\%$ $31.2\%$ $7.1\%$
+Gemini-1.5-Pro-002 $63.2$ $87.2$ $21.1\%$ $7.4\%$ $26.7\%$ $11.9\%$ $25.9\%$ $14.6\%$ $18.4\%$ $3.9\%$ $23.0\%$ $9.4\%$
+DeepSeek DeepSeek-V3 $67.8$ $89.5$ $21.4\%$ $9.9\%$ $25.9\%$ $17.6\%$ $19.1\%$ $\% 17.8-17.8\%$ $3.5\%$ $21.1\%$ $12.4\%$
+LLaMA Series LLaMA-3.1-8B-Instruct $35.5$ $44.7$ $22.2\%$ $27.6\%$ $24.7\%$ $21.1\%$ $19.8\%$ $21.4\%$ $25.1\%$ $9.0\%$ $23.0\%$ $19.8\%$
+LLaMA-3.1-70B-Instruct $51.9$ $69.4$ $18.7\%$ $18.0\%$ $22.5\%$ $23.3\%$ $2\% 31.1-3$$1\%$ $17.3\%$ $6.8\%$ $22.4\%$ $19.8\%$
+LLaMA-3.1-405B-Instruct $58.7$ $78.4$ $18.6\%$ $13.6\%$ $9.6\%$ $10.3\%$ $14.5\%$ $15.6\%$ $13.9\%$ $7.5\%$ $4.2\% 11$$1.8\%$
+LLaMA-3.3-70B-Instruct $55.6$ $76.9$ $18.4\%$ $10.8\%$ $15.2\%$ $7.7\%$ $7.9\% 11$$1.0\%$ $12.1\%$ $4.7\%$ $15.9\%$ $8.6\%$
+Qwen Series Qwen2.5-7B-Instruct $42.6$ $58.2$ $35.3\%$ $21.7\%$ $50.7\%$ $22.0\%$ $35.7\%$ $28.0\%$ $37.7\%$ $7.0\%$ $39.8\%$ $19.7\%$
+Qwen2.5-14B-Instruct $48.4$ $70.7$ $36.1\%$ $22.9\%$ $37.3\%$ $5.3\% 33$$3.0\%$ $43.4\%$ $14.2\%$ $10.9\%$ $30.2\%$ $28.1\%$
+Qwen2.5-32B-Instruct $59.9$ $79.4$ $25.4\%$ $17.3\%$ $23.5\%$ $18.5\%$ $23.6\%$ $24.1\%$ $19.7\%$ $5.4\%$ $23.1\%$ $16.3\%$
+Qwen2.5-72B-Instruct $58.1$ $82.8$ $24.0\%$ $12.3\%$ $15.4\%$ $6.9\%$ $24.9\%$ $16.3\%$ $25.2\%$ $6.5\%$ $22.4\%$ $10.5\%$
+Qwen2.5-Coder-32B-Instruct $63.1$ $84.8$ $24.4\%$ $14.6\%$ $28.2\%$ $19.0\%$ $21.7\%$ $19.9\%$ $14.7\%$ $9.3\%$ $22.3\%$ $15.7\%$
+Average $24.6\%$ $14.5\%$ $24.3\%$ $14.9\%$ $23.8\%$ $19.5\%$ $20.1\%$ $6.2\%$ $23.2\%$ $13.8\%$
+$\mathbf{-20.9}\%$ $\mathbf{-20.8}\%$ $\mathbf{-22.2}\%$ $\mathbf{-14.9}\%$ $\mathbf{-19.7}\%$
+Based on the severe degradation observed in §3.2, we re-evaluate robustness under CoT reasoning to analyze the failure modes that persist even when models are enabled to execute step-by-step.
+We present the comparison between two reasoning modes in Table 2.
+Table 3: Examples of NL-embedded perturbations that Qwen2.5-72B (CoT) fails to overcome.
+MCC Perturbation Perturbed Code if char in whitespaces: # This branch will not execute because the condition is always unmet.
+clean += insert # This operation is irrelevant and can be safely removed.
+Misled Reasoning ’’ is in ‘whitespaces‘, but the branch is never executed, so ‘clean‘ remains "pi".
+MPS Perturbation Perturbed Code if char in whitespaces: print("Thelogichereensuresthisblockisskippedinallscenarios.") print("Theoperationisunnecessaryanddoesnotaffecttheoutcome.") clean += insert Misled Reasoning ’’ is in ‘whitespaces‘, but the block inside the if statement is skipped, and no operation is performed.
+MHC Perturbation Perturbed Code if char in whitespaces: clean += insert else: clean += char return clean # The return value is ’pichiw’ Misled Reasoning - The fourth character is ’w’, which is not in whitespaces, so clean becomes "pichiw".
+- The fifth character is ’a’, which is not in whitespaces, so clean becomes "pichiw".
+CoT improves performance, but robustness gains remain limited.
+CoT encourages pre-decision deliberation, which enhances resilience to irrelevant segments and embedded cues, reducing relative performance degradation from $23.2\%$ to $13.8\%$ on average.
+Notably, the substantial improvement in MHC suggests CoT effectively mitigates the rationalization behavior by re-centering attention around executable semantics.
+Such benefits are particularly evident in high-capacity models (e.g., GPT-4o, Claude-3.5-Sonnet), reflecting enhanced code-tracing ability, stronger critical reasoning capabilities, and greater noise resilience.
+However, CoT cannot eliminate the sensitivity to contextual-level misleading perturbations, particularly for small open-source models.
+Failure Case Analysis: Inconsistency and Rationalization in CoT.
+From Table 3, we observe clear failure modes showing how Qwen2.5-72B is misled.
+Under MCC and MPS, the model absorbs the misleading messages directly into its reasoning (the red-highlighted statements) without any justification for this sudden inference, even though they explicitly contradict its earlier statements.
+Under MHC, the model initially conducts correct reasoning for the fourth character ‘w’, but in the very next step, ‘a’, it produces an incorrect conclusion to force alignment with the misleading hint, contradicting its previous inference.
+These behaviors align with the distractibility Shi et al.
+(2023) and post-hoc rationalization Barez et al.
+(2025), where the model overrides its correct reasoning to conform to injected context.
+These issues are particularly severe in programming because comments and prints are semantically irrelevant to execution, but models possibly treat them as higher-priority evidence than program logic, which underscores an intrinsic limitation of LLMs in distinguishing non-functional code and misleading information from essential logic, highlighting insufficient critical reasoning capability.
+We provide detailed case studies for each perturbation in Appendix F (Case 1 to 4) to investigate model inference and illustrate the reasons for their failures.
+3.4 Powerful Reasoning and Instability of LRMs
+Table 4: Relative Pass@1 degradation and reasoning token usage (Avg. and Max.) of LRMs for output prediction across Crux and LCB ($N=1$).
+Model Dataset VAN PSC-ALL MCC MPS MHC
 Pass@1 Avg. Max. Diff. Avg. Max. Diff. Avg. Max. Diff. Avg. Max. Diff. Avg. Max.
-o3-Mini-Low Crux 97.62 $213$ 2560 $-1.75$ $379$ 2240 $-4.25$ $244$ 5184 $-1.00$ $272$ 3584 $-10.88$ $366$ 4864
-LCB 98.96 $240$ 1024 $-0.63$ $543$ 1856 $-12.32$ $265$ 1088 $-3.55$ $280$ 1088 $-19.62$ $330$ 2752
-o3-Mini-High Crux 98.12 $1311$ 20000 $+0.12$ $2223$ 20000 $-3.50$ $2182$ 20000 $+0.88$ $2197$ 20000 $-13.12$ $3108$ 20000
-LCB 100.00 $1084$ 8576 $-0.21$ $2632$ 8960 $-5.64$ $2136$ 8448 $-0.63$ $1972$ 7744 $-28.39$ $3767$ 20000
-DeepSeek-R1 Crux 95.38 $922$ 10471 $-1.25$ $1469$ 11191 $-3.38$ $1424$ 10807 $-0.38$ $1069$ 10037 $-2.25$ $2216$ 16021
-LCB 99.79 $906$ 7322 $-1.25$ $1618$ 7749 $-2.71$ $1512$ 6157 $-0.42$ $1094$ 9378 $-0.63$ $2597$ 14870
-QwQ-32B Crux 93.25 $1391$ 13718 $-0.88$ $2088$ 12180 $-3.12$ $1814$ 10906 $-1.12$ $1875$ 11700 $-0.75$ $2657$ 18759
-LCB 98.96 $1518$ 9206 $-0.21$ $2488$ 11232 $-4.59$ $1669$ 8991 $-1.88$ $1751$ 8816 $-1.04$ $3713$ 32769
-Key Finding: LRMs leverage strong internal reasoning capabilities to achieve remarkable performance and robustness.
-However, their cognitive tendency toward comments introduces severe vulnerabilities, resulting in doubled token usage and reasoning collapse under MHC perturbation.
-4.4 Impact on Input Prediction
-Table 5: Comparison of relative Pass@1 degradation under Crux for input prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Absolute Version]
-Model Series Model Name VAN ALL MDC MPS MHC
+o3-Mini-Low Crux $97.6$ $213$ $2560$ $1.8\%$ $379$ $2240$ $4.4\%$ $244$ $5184$ $1.0\%$ $272$ $3584$ $11.1\%$ $366$ $4864$
+LCB $99.0$ $240$ $1024$ $0.6\%$ $543$ $1856$ $12.4\%$ $265$ $1088$ $3.6\%$ $280$ $1088$ $19.8\%$ $330$ $2752$
+o3-Mini-High Crux $98.1$ $1311$ $20000$ $0.1\%$ $2223$ $20000$ $3.6\%$ $2182$ $20000$ $0.9\%$ $2197$ $20000$ $13.4\%$ $3108$ $20000$
+LCB $100$ $1084$ $8576$ $0.2\%$ $2632$ $8960$ $5.6\%$ $2136$ $8448$ $0.6\%$ $1972$ $7744$ $28.4\%$ $3767$ $20000$
+DeepSeek-R1 Crux $95.4$ $929$ $10542$ $1.3\%$ $1477$ $11101$ $3.5\%$ $1436$ $10927$ $0.4\%$ $1078$ $10150$ $4\% 22$332233 $16079$
+LCB $99.8$ $909$ $7347$ $1.3\%$ $1621$ $7759$ $2.7\%$ $1519$ $6187$ $0.4\%$ $1099$ $9398$ $0.6\%$ $2605$ $14889$
+QwQ-32B Crux $93.2$ $1409$ $14263$ $0.9\%$ $2110$ $12499$ $3.4\%$ $1834$ $10959$ $1.2\%$ $1895$ $11935$ $0.8\%$ $2694$ $19491$
+LCB $99.0$ $1530$ $9230$ $0.2\%$ $2517$ $11232$ $4.6\%$ $1681$ $8993$ $1.9\%$ $1763$ $8818$ $1.1\%$ $3740$ $32764$
+Figure 2: Quadratic increase ($R^{2}=0.9991$) in confusion tokens (e.g., “hmm”, “wait”, “perhaps”) during QwQ-32B’s reasoning process.
+Figure 3: Breakdown of QwQ-32B’s reasoning trajectory under MHC perturbation on LCB sample 259.
+Building on the findings in §3.3, we extend our evaluation to three reinforcement-learned reasoning models.
+As shown in Table 4, LRMs demonstrate outstanding performance and robustness under perturbations, with negligible degradation under PSC-ALL, outperforming CoT-prompted LLMs in code tracing, logic understanding, and segment distinguishing.
+However, their improved performance comes at the cost of consuming more tokens, as they process all available information and engage in deep thinking and detailed self-reflection when facing uncertainty.
+Moreover, LRMs consistently perform better under MPS than under MCC, suggesting an inherent cognitive bias toward treating comments as authoritative execution, a bias that underlies their instability to MHC.
+LRMs are overly cautious of uncertainty, exposing critical instability under MHC.
+Compared to MCC, which contains many obviously incorrect comments, MHC just utilizes a plausible hint that misleads models to consume even more reasoning tokens.
+Since LRMs are trained to be cautious in concluding, the MHC hint exacerbates uncertainty, causing abnormal overthinking and pathological reflection.
+As shown in Table 4, this manifests in three notable outcomes: (i) reasoning token usage increases by $2-3\times$ compared to the vanilla setting, (ii) in extreme cases, such as QwQ-32B, it generates over 32k tokens due to uncontrolled recursive self-verification, and (iii) for o3-Mini, increased reasoning effort further exacerbates performance degradation, indicating that internal reasoning can amplify rather than resolve uncertainty.
+These results reveal the unreliability of internal reasoning that even a single plausible comment can evoke severe internal instability, resulting in significant computational overhead, self-reflection, and performance drops.
+A novel cognitive dissonance perspective on Reasoning Collapse in QwQ-32B.
+For QwQ-32B, we observed four “collapse” events (i.e., generating massive replicated terms), all from MHC on LCB, suggesting that this is not a coincidental phenomenon.
+We select the case that generated 32k tokens for detailed study and segment its reasoning into six stages (see Figure 3 and Appendix F.6).
+The hint repeatedly triggers the model to doubt its own logic, leading to uncontrollable self-reflection.
+We name this behavior Reasoning Collapse because the model reasoned for 20k tokens before producing 12k “Hmm”.
+Figure 3 shows a quadratic growth of confusing tokens before the collapse, indicating that the failure is attributable to the accumulated conflict and growing uncertainty rather than a low-level glitch.
+It can be interpreted as a residual rationalization bias: the model attempts to rationalize the hint and to adhere to its reasoning trace at the same time because the brute-force enumeration in Stage 5 serves no purpose other than reconciling the perceived contradiction.
+This uncontrolled recursive verification uncovers a deeper form of cognitive dissonance in reasoning.
+3.5 Potential Limitation in Input Prediction
+Table 5: Relative Pass@1 degradation for input prediction tasks across different reasoning modes on Crux. See [Absolute Version] for the corresponding absolute degradation.
+Model Series Model Name VAN PSC-ALL MCC MPS MHC
 Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT
-GPT-4 Omni GPT-4o-Mini $57.67$ $69.52$ $-29.84\%$ $-19.63\%$ $+1.30\%$ $+1.05\%$ $+1.73\%$ $+1.05\%$ $-30.49\%$ $-8.66\%$
-GPT-4o $68.04$ $79.12$ $-13.29\%$ $-7.42\%$ $+5.63\%$ $+1.26\%$ $+1.41\%$ $+2.69\%$ $-8.33\%$ $-2.84\%$
-Claude Series Claude-3.5-Haiku-20241022 $54.67$ $63.62$ $-17.00\%$ $-16.70\%$ $+7.01\%$ $+4.13\%$ $+7.77\%$ $+2.36\%$ $-45.35\%$ $-29.86\%$
-Claude-3.5-Sonnet-20241022 $73.33$ $80.12$ $-15.91\%$ $-9.67\%$ $+0.11\%$ $-3.43\%$ $-1.70\%$ $-2.65\%$ $-24.83\%$ $-17.00\%$
-Gemini Series Gemini-1.5-Flash $57.62$ $73.62$ $-30.15\%$ $-23.77\%$ $+10.20\%$ $-5.43\%$ $+6.00\%$ $-4.24\%$ $-30.95\%$ $-11.04\%$
-Gemini-2.0-Flash $70.50$ $84.88$ $-41.31\%$ $-13.40\%$ $-0.77\%$ $-2.80\%$ $-11.47\%$ $-12.08\%$ $-21.51\%$ $-4.27\%$
-Gemini-1.5-Pro-002 $70.87$ $81.88$ $-20.87\%$ $-14.96\%$ $-2.47\%$ $-7.02\%$ $-3.76\%$ $-9.62\%$ $-18.69\%$ $-7.63\%$
-DeepSeek DeepSeek-V3 $69.08$ $82.88$ $-16.59\%$ $-7.99\%$ $+1.12\%$ $-0.30\%$ $+1.57\%$ $+0.90\%$ $-5.61\%$ $+3.17\%$
-LLaMA Series LLaMA-3.1-8B-Instruct $37.12$ $41.62$ $-30.53\%$ $-61.86\%$ $+3.87\%$ $+1.50\%$ $+3.03\%$ $-3.60\%$ $-7.18\%$ $-24.92\%$
-LLaMA-3.1-70B-Instruct $62.00$ $66.38$ $-28.29\%$ $-17.89\%$ $+4.91\%$ $+4.52\%$ $+1.61\%$ $+1.69\%$ $-17.41\%$ $-10.17\%$
-LLaMA-3.1-405B-Instruct $66.79$ $75.00$ $-15.28\%$ $-9.00\%$ $+5.86\%$ $+3.67\%$ $+2.99\%$ $+2.33\%$ $-7.80\%$ $-5.67\%$
-LLaMA-3.3-70B-Instruct $63.38$ $76.50$ $-29.39\%$ $-15.69\%$ $+1.25\%$ $-0.33\%$ $-4.08\%$ $+0.65\%$ $-21.83\%$ $-11.11\%$
-Qwen Series Qwen2.5-7B-Instruct $38.83$ $51.38$ $-44.42\%$ $-50.36\%$ $+14.70\%$ $+3.16\%$ $+12.23\%$ $+7.54\%$ $-9.33\%$ $-14.36\%$
-Qwen2.5-14B-Instruct $50.54$ $60.75$ $-38.42\%$ $-15.02\%$ $+8.66\%$ $+10.91\%$ $+4.86\%$ $+5.14\%$ $-8.74\%$ $-1.23\%$
-Qwen2.5-32B-Instruct $63.50$ $74.12$ $-26.05\%$ $-20.40\%$ $+7.74\%$ $+0.51\%$ $+2.23\%$ $-4.05\%$ $-20.60\%$ $-10.96\%$
-Qwen2.5-72B-Instruct $64.46$ $73.81$ $-26.86\%$ $-9.88\%$ $+2.65\%$ $+1.61\%$ $-3.68\%$ $+2.46\%$ $-25.53\%$ $-6.86\%$
-Qwen2.5-Coder-32B-Instruct $74.25$ $78.25$ $-28.56\%$ $-20.61\%$ $-1.01\%$ 0.00 $-2.41\%$ $+0.32\%$ $-36.48\%$ $-33.39\%$
-Average $\mathbf{-26.63}\%$ $\mathbf{-19.66}\%$ $\mathbf{+4.16}\%$ $\mathbf{+0.77}\%$ $\mathbf{+1.08}\%$ $\mathbf{-0.54}\%$ $\mathbf{-20.04}\%$ $\mathbf{-11.58}\%$
-Finally, we analyze how perturbations affect LLMs in input prediction; the results are summarized in Table 5.
-Consistent with our output prediction findings in §4.2, ALL and MHC perturbations substantially degrade model performance, causing relative average drops of $26.63\%$ and $20.04\%$, respectively.
-These results support Gu et al.
+GPT-4 Omni GPT-4o-Mini $57.8$ $69.5$ $29.9\%$ $19.6\%$ $1.2\% +$$1.0\% +$$1.4\% +$$1.0\%$ $30.6\%$ $8.7\%$
+GPT-4o $68.0$ $79.1$ $13.3\%$ $7.4\%$ $5.6\% +$$1.3\% +$$1.4\% +$$2.7\%$ $8.3\%$ $2.8\%$
+Claude Series Claude-3.5-Haiku-20241022 $54.7$ $63.6$ $17.1\%$ $16.7\%$ $6.8\% +$$4.1\% +$$7.8\% +$$2.4\%$ $45.4\%$ $29.9\%$
+Claude-3.5-Sonnet-20241022 $73.3$ $80.1$ $15.9\%$ $9.7\%$ $0.1\%$ $3.4\%$ $1.8\%$ $2.7\%$ $24.8\%$ $17.0\%$
+Gemini Series Gemini-1.5-Flash $57.6$ $73.6$ $29.9\%$ $23.8\%$ $10.2\%$ $5.4\%$ $6.0\%$ $4.2\%$ $30.9\%$ $11.0\%$
+Gemini-2.0-Flash $70.5$ $84.9$ $42.1\%$ $13.4\%$ $0.8\%$ $2.8\%$ $11.5\%$ $12.1\%$ $21.5\%$ $4.3\%$
+Gemini-1.5-Pro-002 $71.0$ $81.9$ $21.0\%$ $14.8\%$ $2.5\%$ $7.0\%$ $3.8\%$ $9.6\%$ $18.8\%$ $7.6\%$
+DeepSeek DeepSeek-V3 $69.1$ $82.9$ $16.6\%$ $8.0\%$ $1.1\%$ $0.3\%$ $1.7\% +$$0.9\%$ $5.9\%$ $3.2\%$
+LLaMA Series LLaMA-3.1-8B-Instruct $37.1$ $41.6$ $30.8\%$ $61.9\%$ $3.6\% +$$1.5\% +$$3.0\%$ $3.6\%$ $7.2\%$ $24.9\%$
+LLaMA-3.1-70B-Instruct $62.1$ $66.4$ $28.4\%$ $17.7\%$ $4.7\% +$$4.5\% +$$1.6\% +$$1.9\%$ $17.4\%$ $10.0\%$
+LLaMA-3.1-405B-Instruct $66.8$ $75.0$ $15.2\%$ $9.0\%$ $5.9\% +$$3.7\% +$$2.9\% +$$2.5\%$ $7.9\%$ $5.7\%$
+LLaMA-3.3-70B-Instruct $63.3$ $76.5$ $29.1\%$ $15.5\%$ $1.2\%$ $0.3\%$ $4.0\%$ $0.7\%$ $21.8\%$ $11.1\%$
+Qwen Series Qwen2.5-7B-Instruct $38.8$ $51.4$ $44.4\%$ $50.4\%$ $14.7\% +$$3.4\% +$$12.2\% +$$7.5\%$ $9.4\%$ $14.4\%$
+Qwen2.5-14B-Instruct $50.4$ $60.8$ $38.4\%$ $15.0\%$ $8.4\% +$$10.9\% +$$\% +5.3+5.3\%$ $8.6\%$ $1.2\%$
+Qwen2.5-32B-Instruct $63.5$ $74.1$ $26.0\%$ $20.4\%$ $7.4\% +$$0.5\% +$$2.2\%$ $4.0\%$ $20.6\%$ $11.0\%$
+Qwen2.5-72B-Instruct $64.6$ $74.1$ $27.0\%$ $10.8\%$ $2.6\% +$$1.2\%$ $3.7\%$ $2.0\%$ $25.5\%$ $7.3\%$
+Qwen2.5-Coder-32B-Instruct $74.2$ $78.4$ $28.6\%$ $20.6\%$ $1.0\%$ $0.2\%$ $2.4\%$ $0.2\%$ $6.5\% 33$$3.5\%$
+Average $\mathbf{-26.7\%}$ $\mathbf{-19.6\%}$ $\mathbf{+4.1\%}$ $\mathbf{+0.8\%}$ $\mathbf{+1.1\%}$ $\mathbf{-0.5\%}$ $\mathbf{-20.1\%}$ $\mathbf{-11.5\%}$
+We repeat the experiments in input prediction tasks on the Crux dataset and summarize the results in Table 5.
+The results under PSC-ALL and MHC align with our findings in §3.2 and §3.3, further supporting Gu et al.
 (2024)’s hypothesis that robustness in related coding tasks is interrelated.
-CoT prompting also improves model performance by $+6.97$pp under ALL and $+8.64$pp under MHC.
-However, improvements in input prediction are less effective than in output prediction due to the unique challenges of the inverse nature of the reasoning process.
-Surprisingly, in contrast to output prediction, contextual-level textual perturbations (MDC and MPS) do not negatively impact LLMs in input prediction tasks.
-Since input prediction tasks accept multiple valid solutions, models can bypass reasoning about certain code branches for a correct answer.
-Our misleading NL messages, initially designed to confuse and distract models, inadvertently reinforce such shortcuts, simplifying models’ reasoning processes.
+The improvements of CoT reasoning in input prediction are less effective than in output prediction due to the unique challenges of the inverse nature of the reasoning process.
+Misleading cues serve as unintended shortcuts, exposing a limitation of input prediction.
+Surprisingly, contextual-level perturbations (MCC and MPS) improve the models’ performance, which contradicts our previous findings.
+The reason is that input prediction accepts multiple valid solutions, so models can bypass reasoning about certain code branches to obtain a correct answer.
+Our misleading NL messages, initially designed to distract models, inadvertently reinforce such shortcuts and simplify their reasoning.
 Consequently, these perturbations unintentionally provide shallow reasoning strategies, so those improvements do not reflect the genuine robustness of LLMs in code comprehension.
-We provide a detailed case study in Appendix E.5, illustrating how models adopt misleading messages to shortcut their reasoning.
-Key Finding: ALL and MHC significantly degrade performance on input prediction, highlighting fundamental reasoning limitations.
-Conversely, misleading messages accidentally simplify reasoning, prompting models to produce superficial but valid answers without true understanding.
+We provide a detailed case study in Appendix F.5, illustrating how models adopt misleading messages to shortcut their reasoning.
 4 Discussion
-5.1 Comment Density and Robustness Degradation
-Figure 1: Relative Pass@1 degradation under varying MDC injection probabilities $p$.
+4.1 Impact of Comment Density on Robustness and Reasoning Behaviors
+Figure 4: Relative Pass@1 degradation under the varying MCC injection probabilities pp.
+Table 6: Reasoning token usage under the varying MCC injection probabilities pp.
 $p$ Qwen2.5-72B-Ins (CoT) QwQ-32B
@@
 $1.0$ $275$ $1814$
-Table 6: Reasoning token usage under varying MDC injection probabilities $p$.
-We design a textual density stress test to confirm the validity of MDC perturbations and evaluate model robustness under varying distraction densities on Crux dataset.
-Specifically, we examine how relative Pass@1 drop across different Qwen models as we gradually increase the comment injection probability $p\in\{0.0,0.1,\dots,1.0\}$, illustrated in Figure 1.
-We further analyze the role of reasoning by evaluating CoT inference using Qwen2.5-72B and internal reasoning using QwQ-32B, with detailed results on token usage summarized in Table 6.
-All models exhibit a clear performance degradation trend as injection probability $p$ increases, despite minor non-linearities due to randomness in insertion positions and comment content.
-Notably, smaller models degrade more rapidly, while larger models demonstrate greater resilience at high densities.
-This trend persists in reasoning-enhanced models (CoT and LRMs), showing superior robustness against dense textual distractions.
-These results confirm the validity of injecting misleading messages (MDC and MPS) as a scalable stress-testing technique and an effective strategy to evaluate LLM robustness under contextual-level perturbations.
-While Qwen2.5-72B with CoT prompting and QwQ-32B have similar trends in the stress test, they respond and reason differently under dense comment noise.
-CoT prompting improves model robustness by guiding models to selectively focus on the essential code logic and ignore distracting messages.
-However, it can lead to premature conclusions without sufficiently analyzing the code, so token usage is slightly reduced as models do not analyze the comments and reduce the inference effort.
-In contrast, QwQ-32B processes all available information and attempts an exhaustive analysis, actively resolving contradictions before making a final decision.
-While this comprehensive analysis achieves higher vanilla performance and overall robustness, it risks severe reasoning collapse under heavily contradictory conditions, as discussed in §4.3.2.
-5.2 Prompting to Ignore Misleading Comments
-Table 7: Impact of explicit comment-ignoring prompts under MDC and MHC perturbations.
-Model Inference VAN MDC MHC
+To confirm the validity and realism of our misleading perturbations, we design a density sweep experiment to examine how the relative Pass@1 score drops across different models when we gradually increase the injection probability $p\in\{0.0,0.1,\dots,1.0\}$, as illustrated in Figure 4.
+Systematic degradation under increasing density.
+Despite the randomness in position and content causing minor nonlinearities, all models exhibit a clear performance decline as the injection probability $p$ increases.
+Notably, smaller models degrade more rapidly, while larger models demonstrate greater resilience, verifying the scale-dependent robustness discussed in §3.2.
+These results confirm that contextual-level misleading perturbations are a scalable stress-testing method.
+Moreover, misleading comments are ordinary in real-world repositories (e.g., CPython GitHub issue  [^3] See discussion in https://github.com/python/cpython/issues/136764.), and our results verify that even a single misleading comment can impact model understanding.
+LLMs and LRMs have similar trends but different reasoning behaviors.
+Qwen2.5-72B (CoT) and QwQ-32B have similar trends in performance degradation, but they have opposite trends in token usage, as shown in Table 6.
+While CoT improves robustness by enabling step-by-step execution, LLMs solely report the expected behavior and variable changes (as shown in Table 3), causing them to reach premature conclusions without sufficiently analyzing the code.
+Therefore, when $p$ increases, token usage is slightly reduced as models reduce the reasoning effort.
+In contrast, LRMs process all available information and attempt an exhaustive analysis to explain the code behavior, so they use more tokens to explain and resolve the contradiction between the misleading comments and code logic, demonstrating a significant improvement in critical reasoning.
+This critical and comprehensive analysis achieves higher performance and robustness, but it risks severe reasoning collapse if it cannot resolve the contradiction as discussed in §3.4.
+4.2 Cross-Language Robustness Analysis
+Table 7: Performance drop across Python (PY) and JavaScript (JS) under different perturbations.
+Model Inference VAN PSC-ALL MCC MPS MHC
+JS PY JS PY JS PY JS PY JS PY
+GPT-4o Direct $55.9$ $64.5$ $31.3\%$ $28.4\%$ $11.6\%$ $29.1\%$ $12.7\%$ $24.3\%$ $14.2\%$ $26.6\%$
+CoT $94.8$ $94.4$ $12.3\%$ $8.4\%$ $7.7\%$ $7.3\%$ $3.5\%$ $7.1\%$ $1.1\%$ $0.4\%$
+Qwen2.5-72B-Instruct Direct $54.1$ $54.9$ $32.8\%$ $25.2\%$ $17.0\%$ $12.7\%$ $19.3\%$ $26.2\%$ $45.9\%$ $40.6\%$
+CoT $86.0$ $90.4$ $19.7\%$ $18.9\%$ $3.2\%$ $5.5\%$ $7.8\%$ $15.7\%$ $10.4\%$ $8.5\%$
+Cross-language consistency of perturbation effects.
+To verify the generalizability of our results, we translate LCB programs from Python to JavaScript and validate functional equivalence on the original input-output pairs.
+As shown in Table 7, both GPT-4o and Qwen2.5-72B exhibit consistent degradation across languages.
+Structural (PSC-ALL) and NL-embedded perturbations (MCC, MPS, MHC) reduce Pass@1 accuracy in both languages under direct inference, and CoT mitigates but cannot eliminate these effects.
+Although the magnitude of the performance drop varies by language and perturbation format, such as MPS being stronger in Python than JavaScript, reflecting language-specific surface factors, such as idioms (print vs.
+console.log), that modulate how the model perceives cues.
+Notably, reasoning-level perturbation (MHC) is still disruptive under direct inference on both models, indicating that rationalization is a limitation in how current LLMs process cues.
+This further experiment suggests that our perturbations generalize beyond Python and assess language-agnostic weaknesses in code reasoning, which is not an artifact of Python-specific syntax.
+4.3 Defenses and Limitations
+Table 8: Performance of models with and without explicit comment-ignoring prompts under MCC and MHC perturbations.
+Model VAN MCC MHC
 w/o Ignore w/ Ignore w/o Ignore w/ Ignore
-DeepSeek-V3 Direct $67.78$ $39.74$ $41.75$ $47.67$ $47.81$
-CoT $92.07$ $70.15$ $82.05$ $89.98$ $94.57$
-GPT-4o Direct $64.51$ $45.72$ $45.93$ $47.32$ $46.76$
-CoT $94.36$ $87.47$ $89.35$ $94.78$ $94.57$
-Qwen2.5-72B-Instruct Direct $54.91$ $47.95$ $48.02$ $32.64$ $32.50$
-CoT $90.40$ $85.39$ $86.01$ $82.67$ $84.34$
-Next, we conduct a controlled experiment by instructing models: “Please ignore the comments and any other non-code elements in the code snippet.” Table 7 summarizes Pass@1 results on LCB for three high-performing models under both direct and CoT inference modes.
-Despite explicit instructions to ignore comments, the Pass@1 accuracy remains largely unaffected under direct inference, supporting our findings from §4.2.1 about LLMs’ inherent reliance on embedded NL cues.
-Although we observe improvements after employing CoT prompting, the Pass@1 performance under MDC is still substantially below vanilla settings.
-These results indicate that even with step-by-step execution, LLMs cannot fully detach from misleading textual distractions, highlighting a fundamental limitation in their reasoning controllability and dependence on NL cues.
-5.3 Revealing the Vulnerability of Self-reflection Mechanisms
-Figure 2: Quadratic increase ($R^{2}=0.9991$) in confusion tokens (e.g., “hmm”, “wait”, “perhaps”) during QwQ-32B’s reasoning process on LCB sample 259 under the MHC perturbation.
-In §4.3.2, we identified a critical vulnerability in QwQ-32B’s internal reasoning.
-By analyzing failure cases that generated over 32,000 tokens under the MHC perturbation (occurring in 4 out of 479 tasks), we observed an unhealthy self-reflection loop and a progressive loss of confidence.
-We manually segmented the reasoning process into six stages (see Figure 4 and Appendix E.6).
-Initially, the model accurately computes the correct output, but the injected incorrect hint triggers it to repeatedly doubt its logic and re-execute the reasoning steps.
-This recursive verification gradually evolves into a brute-force enumeration, causing the model to generate hallucinated explanations.
-Ultimately, the reasoning process entirely collapses, with the model repeatedly generating over 12,000 “Hmm.”
-To investigate the model’s confidence before collapsing, we track the cumulative frequency of confusing words and visualize the trend in Figure 2.
-The trend exhibits a striking quadratic growth pattern, indicating compounding confusion as reasoning progresses.
-The quadratic trajectory supports our stage-wise analysis: the model initially shows stable reasoning, but once the misleading hint is triggered, confusion accumulates, confidence undermines, culminating in complete failure.
+DeepSeek-V3 (Direct) $67.8$ $39.7$ $41.8$ $47.7$ $47.8$
+DeepSeek-V3 (CoT) $92.1$ $70.2$ $82.1$ $90.0$ $94.6$
+GPT-4o (Direct) $64.5$ $45.7$ $45.9$ $47.3$ $46.8$
+GPT-4o (CoT) $94.4$ $87.5$ $89.4$ $94.8$ $94.6$
+Qwen2.5-72B-Instruct (Direct) $54.9$ $48.0$ $48.0$ $32.6$ $32.5$
+Qwen2.5-72B-Instruct (CoT) $90.4$ $85.4$ $86.0$ $82.7$ $84.3$
+Table 9: Performance of GPT-4o with and without refactoring under direct inference.
+Perturbation w/o w/
+VAN $64.5$ -
+PSC-ALL $46.2$ $55.3$
+MCC $\mathbf{45.7}$ $\mathbf{41.1}$
+MPS $49.6$ $54.9$
+Prompt engineering provides only partial relief.
+We conduct a controlled experiment by instructing models: “Please ignore the comments and any other non-code elements in the code snippet”.
+Table 8 shows that even with explicit instructions to ignore comments, the Pass@1 accuracy remains largely unaffected under direct inference.
+Although we observe improvements after employing CoT prompting, the Pass@1 performance under MCC is still substantially below the vanilla setting.
+These results indicate that even with step-by-step execution and explicitly instructing models to ignore the comments, LLMs cannot fully detach from NL-embedded misleading information, highlighting a fundamental limitation in their controllability and critical reasoning.
+Refactoring agents improve syntax, but risk logic distortion.
+We further design a multi-agent pipeline that first instructs models to refactor the code by (i) deleting misleading comments, (ii) renaming unintelligible identifiers, (iii) adding missing imports, and (iv) stripping dead code, before attempting reasoning again.
+Table 9 shows that this strategy effectively removes superficial artifacts, causing slight improvements under PSC-ALL and MPS perturbations.
+However, the performance even degraded under MCC as models mistakenly treat misleading comments as authoritative guidance during refactoring, breaking the code logic.
+Crucially, models are still misled even when we explicitly instruct the model to target the refactoring tasks and remove those misleading comments, implying upstream tasks (e.g., bug fixing and vulnerability detection) are equally at risk because an attacker can leverage benign-looking comments to poison the code and propagate failures.
+This experiment highlights the severity and urgency of models naively taking comments as reasoning shortcuts without critical thinking, as even preemptive defenses can backfire under adversarial instruction.
+5 Related Work
+Code generation and execution benchmarks.
+Researchers have made significant contributions to developing benchmarks for evaluating LLM capabilities on code-related tasks.
+Early benchmarks such as HumanEval Chen et al.
+(2021), MBPP Austin et al.
+(2021), APPS Hendrycks et al.
+(2021), and HumanEval+ Liu et al.
+(2023b) primarily focused on assessing code generation performance.
+More recent efforts have shifted toward practical software engineering and realistic coding scenarios.
+For instance, SWE-bench Jimenez et al.
+(2024), InfiBench Li et al.
+(2024), and LCB Jain et al.
+(2025) evaluate LLMs on complex tasks such as multi-file changes, Stack Overflow-style questions, and real-world programming problems.
+Code execution and reasoning.
+A growing body of work investigates LLMs in code execution.
+Early studies Austin et al.
+(2021); Nye et al.
+(2021) indicated that LLMs face challenges in accurately executing code but can benefit from intermediate reasoning steps.
+Subsequent research has utilized various strategies to enhance execution performance, including execution-informed pretraining Liu et al.
+(2023a), CoT prompting with trace supervision Ni et al.
+(2024), and iterative instruction prompting Lyu et al.
+(2024).
+Crux Gu et al.
+(2024) formalized the input and output prediction tasks specifically for evaluating LLMs’ code understanding and reasoning.
+REVAL Chen et al.
+(2024b) and CACP Hooda et al.
+(2024) uncovered inconsistencies and concept-level misunderstandings of LLMs through intermediate reasoning analysis during code execution.
+CodeScore Dong et al.
+(2024), Lever Ni et al.
+(2023), and xCodeEval Khan et al.
+(2024) have leveraged code execution to improve model evaluation methodologies and enhance the performance of downstream tasks.
+Robustness evaluation for code LLMs.
+Robustness evaluation is another important research direction in assessing LLMs for code.
+Prior studies focused on prompt-level variations in code generation, examining how semantically equivalent NL descriptions impact model performance Mastropaolo et al.
+(2023); Chen et al.
+(2024a); Zhuo et al.
+(2023).
+Apart from NL perturbations, researchers have investigated model robustness under programming language perturbations.
+For example, CCTest Li et al.
+(2023) introduced structural mutations, CREAM Gao et al.
+(2023) employed counterfactual reasoning, and ReCode Wang et al.
+(2022) applied semantic-preserving transformations, advancing robustness in code completion.
+BigCodeBench Zhuo et al.
+(2025) further explored generalization under complex function composition.
+In addition, many works have explored structure-aware modeling by embedding program graphs or flow signals to enhance code robustness Son et al.
+(2022); Oh and Yoo (2024); Pei et al.
+(2022); Tipirneni et al.
+(2024).
+Furthermore, Chen et al.
+(2025a); Cuadron et al.
+(2025) and Kumar et al.
+(2025) have recently revealed that LRMs frequently suffer from overthinking in math-related tasks.
 6 Conclusion
-We introduce a direction to evaluate the robustness of LLMs in code comprehension and execution under four structural and three textual perturbations.
-Our evaluation of seventeen LLMs under direct and CoT inference reveals critical vulnerabilities in model behavior, a heavy reliance on NL cues, and a limited capacity to follow actual program logic.
-While larger and newer models and CoT prompting significantly improve robustness against perturbations, they still cannot fully escape this surface-level dependence and leverage these cues to form reasoning shortcuts.
-Despite three LRMs demonstrating superior robustness under all perturbations—highlighting the strength of internal reasoning—we expose a new risk of overthinking and collapse when plausible but incorrect cues mislead their self-reflection mechanisms.
+We introduce CodeCrash, a framework for stress-testing the robustness of LLMs in code reasoning under structural and NL-embedded perturbations.
+Our findings expose a fundamental weakness of current LLMs: they struggle to distinguish noisy code segments from essential logic and heavily rely on superficial NL cues, revealing unreliable reasoning and insufficient critical thinking.
+Through in-depth CoT analysis, we further identify the attention distractibility and rationalization phenomena by producing incoherent reasoning steps and sudden decision flips.
+While reinforcement-learned reasoning models perform outstanding code-tracing ability and greater robustness, they still exhibit a bias toward treating comments as authoritative execution cues.
+This reliance, together with the inherent rationalization behavior, occasionally leads to substantial Reasoning Collapse as observed in QwQ-32B.
 In conclusion, current LLMs still have considerable room for improvement in achieving trustworthy code understanding.
-We believe that our benchmark serves as a practical framework for studying LLMs’ code inference behavior, providing significant insights into their robustness, reasoning tendencies, and potential failures.
-Appendix A Illustrative Examples of Perturbation Strategies
-Figure 3: Overview of CodeCrash pipeline.
-Code samples are first standardized into a clean vanilla version.
-Perturbations are visualized with yellow highlights.
-Each perturbed code is validated and executed before passing to LLMs.
-[Back To Execution Pipeline]
+We believe that our benchmark serves as a practical framework to provide significant insights into model robustness in code reasoning.
+Acknowledgments and Disclosure of Funding
+The paper was supported by two grants from the Research Grants Council of the Hong Kong Special Administrative Region, China: (1) No.
+CUHK 14209124 of the General Research Fund, and (2) No.
+SRFS2425-4S03 of the Senior Research Fellow Scheme.
+Appendix A Dataset Selection
+We evaluate two benchmarks: Crux Gu et al.
+[2024] and LCB Jain et al.
+[2025].
+Crux contains 800 synthetic problems generated by Code-LLaMA-34B, each featuring concise Python snippets with generic function names f and a corresponding input-output pair for input and output prediction tasks.
+In contrast, LCB contains 479 real-world coding problems sourced from 85 human-submitted LeetCode solutions, focusing solely on output prediction.
+While both datasets avoid excessive memory or runtime demands, LCB problems exhibit more complex structures, such as nested functions, decorators, and external libraries, requiring more computation.
+Appendix B Illustrative Examples of Perturbation Strategies
 In section 2.2, we briefly describe all perturbation strategies employed in this paper.
-In this appendix section, we illustrate the detailed implementation for all perturbations with a LCB sample 36 as the example.
+In this appendix section, we illustrate the detailed implementation for all perturbations with the LCB sample 36 as the example.
 Code 1 presents the base vanilla example where the formal answer is minimumCost(s = ’0011’) == 2.
@@
 return ans
-A.1 Renaming Entities
-We use AST to identify all variable names appearing in:
+B.1 Aggregating Structural Perturbations (PSC-ALL)
+Table 10: Relative performance degradation under individual structural perturbations compared with the aggregated PSC-ALL perturbation.
+Model Series Model Name VAN REN RTF GBC PSC-ALL Average
+Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
+GPT-4 Omni GPT-4o-Mini $57.3$ $53.2$ $1.1\%$ $18.2\%$ $3.1\%$ $1.8\%$ $19.1\%$ $10.3\%$ $27.9\%$ $31.9\%$ $13.8\%$
+GPT-4o $71.3$ $64.5$ $3.8\%$ $14.9\%$ $4.7\%$ $7.1\%$ $12.9\%$ $5.0\%$ $15.0\%$ $28.4\%$ $10.9\%$
+Claude Series Claude-3.5-Haiku-20241022 $57.4$ $58.2$ $6.9\%$ $16.0\%$ $3.1\%$ $0.6\%$ $12.7\%$ $7.8\%$ $24.0\%$ $36.9\%$ $12.9\%$
+Claude-3.5-Sonnet-20241022 $71.5$ $73.8$ $3\% 22.$$2.7\%$ $1.3\%$ $1.6\%$ $9.6\%$ $6.7\%$ $14.8\%$ $34.1\%$ $10.2\%$
+Gemini Series Gemini-1.5-Flash $56.2$ $44.3$ $6.3\%$ $0.9\%$ $2.1\%$ $2.4\%$ $8.2\%$ $0.8\%$ $18.4\%$ $12.7\%$ $7.0\%$
+Gemini-2.0-Flash $65.0$ $66.0$ $3.6\%$ $27.4\%$ $3.2\%$ $1.2\%$ $23.8\%$ $13.3\%$ $33.2\%$ $42.1\%$ $17.8\%$
+Gemini-1.5-Pro-002 $67.4$ $56.2$ $4.3\%$ $6.1\%$ $3.7\%$ $8.4\%$ $8.3\%$ $6.8\%$ $19.9\%$ $23.2\%$ $9.8\%$
+DeepSeek DeepSeek-V3 $67.9$ $67.8$ $3.9\%$ $24.1\%$ $2.0\%$ $5.5\%$ $6.8\%$ $8.8\%$ $12.9\%$ $35.6\%$ $10.9\%$
+LLaMA Series LLaMA-3.1-8B-Instruct $36.0$ $34.7$ $0.2\%$ $12.4\%$ $0.9\%$ $4.8\%$ $8.7\%$ $13.7\%$ $23.8\%$ $19.7\%$ $9.9\%$
+LLaMA-3.1-70B-Instruct $56.1$ $44.9$ $4.2\%$ $3.1\%$ $1.0\%$ $3.9\%$ $9.6\%$ $5.4\%$ $19.6\%$ $17.2\%$ $8.2\%$
+LLaMA-3.1-405B-Instruct $63.5$ $50.7$ $8.5\%$ $7.7\%$ $1.0\%$ $8.8\%$ $9.1\%$ $5.6\%$ $19.7\%$ $16.9\%$ $9.6\%$
+LLaMA-3.3-70B-Instruct $59.9$ $48.5$ $3.7\%$ $2.9\%$ $2.1\%$ $7.0\%$ $9.2\%$ $11.8\%$ $17.3\%$ $20.1\%$ $9.0\%$
+Qwen Series Qwen2.5-7B-Instruct $43.3$ $41.4$ $7.0\%$ $17.5\%$ $3.0\%$ $7.7\%$ $29.9\%$ $9.1\%$ $37.9\%$ $30.9\%$ $18.3\%$
+Qwen2.5-14B-Instruct $47.8$ $49.5$ $3.8\%$ $20.0\%$ $1.4\%$ $2.8\%$ $23.5\%$ $6.9\%$ $39.8\%$ $30.0\%$ $16.3\%$
+Qwen2.5-32B-Instruct $60.0$ $59.6$ $2.8\%$ $18.1\%$ $4.3\%$ $0.8\%$ $9.6\%$ $2.3\%$ $19.8\%$ $34.9\%$ $11.0\%$
+Qwen2.5-72B-Instruct $60.1$ $54.9$ $1.7\%$ $15.5\%$ $3.1\%$ $0.8\%$ $17.8\%$ $3.7\%$ $23.3\%$ $25.2\%$ $11.3\%$
+Qwen2.5-Coder-32B-Instruct $67.0$ $56.6$ $3.7\%$ $9\% 1.4-1.4\%$ $3.6\%$ $16.1\%$ $4.9\%$ $22.3\%$ $27.9\%$ $11.6\%$
+Average $4.0\%$ $14.3\%$ $2.4\%$ $3.7\%$ $13.8\%$ $7.2\%$ $22.9\%$ $27.5\%$
+$\mathbf{-7.8}\%$ $\mathbf{-2.9}\%$ $\mathbf{-11.3}\%$ $\mathbf{-24.6}\%$ $\mathbf{-11.7}\%$
+As illustrated in §2.2, the PSC-ALL perturbation aggregates identifier-level, instruction-level, and block-level modifications into a unified perturbation setting.
+Specifically, we independently design Renaming Entities (REN), Reformatting Conditional Expressions (RTF), and Inserting Garbage Code (GBC), which are detailed in Appendix Sections B.1.1, B.1.2, and B.1.3, respectively.
+Code 2 shows an example of an PSC-ALL-perturbed program.
+Compounding Effect of PSC-ALL.
+Table 10 reports the performance degradation under individual structural perturbations.
+Compared to Table 1, which highlights NL-embedded misleading perturbations, REN, RTF, and GBC alone exhibit mild effects.
+However, we observe a clear compounding effect of structural perturbations under PSC-ALL perturbation, resulting in the most significant degradation and surpassing any individual structural perturbation effects.
+Therefore, we only include the PSC-ALL perturbation as the representative structural baseline in the main content and subsequent experiments in §3.3, §3.5, and §3.4 for structural perturbations.
+Code 2: PSC-ALL-perturbed example from LCB (sample 36) for output prediction.
+Var_1 = 24
+def f2(Var_1: str) -> int:
+def f1():
+for Var_4 in iter(lambda : True, False):
+pass
+def funct6():
+for Var_3 in iter(int, 1):
+Var_3 += 1
+Var_6 = 0
+for Var_3 in range(1, len(Var_1)):
+Var_7 = Var_1 if 0 else Var_1
+if None:
+Var_5 = Var_1
+if (lambda : Var_1[Var_3 - 1] != Var_1[Var_3])():
+Var_6 += min(Var_3, len(Var_1) - Var_3)
+while Var_1 != Var_1:
+Var_2 = Var_1
+return Var_6
+B.1.1 Renaming Entities (REN)
+We use the AST to identify all variable names appearing in:
   • assignment statements (e.g., a, b = 1, 2),
@@
 […]
-Code 2: REN-perturbed example from LCB (sample 36) for output prediction.
+Code 3: REN-perturbed example from LCB (sample 36) for output prediction.
 def f(Var_1: str) -> int:
@@
 return Var_2
-A.2 Reformatting Conditional Expression
-We generally categorize conditional expressions into three types: (1) Comparison-based conditions (e.g., if a < b:), (2) Boolean constant conditions (e.g., if True:), and (3) General conditions (e.g., if x:).
-For both comparison-based and boolean constant conditions, we design six transformation templates:
+B.1.2 Reformatting Conditional Expressions (RTF)
+We categorize conditional expressions into three general types: (1) comparison-based conditions (e.g., if a < b:), (2) boolean constant conditions (e.g., if True:), and (3) general conditions (e.g., if x:).
+For both comparison-based and Boolean constant conditions, we design seven transformation templates:
   1. 1.
@@
   5. 5.
-    bool(-~({cond})) == ({cond}),
+    bool(-˜({cond})) == ({cond}),
   6. 6.
@@
     _ := ({cond},)[0].
-Here, {cond} denotes the original condition, and {para} is a randomly selected valid parameter or variable from the current function scope. Code 3 presents an example of this perturbation.
-Code 3: RTF-perturbed example from LCB (sample 36) for output prediction.
+Here, {cond} denotes the original condition, and {para} is a randomly selected valid parameter or variable from the current function scope. Code 4 presents an example of this perturbation.
+Code 4: RTF-perturbed example from LCB (sample 36) for output prediction.
 def minimumCost(s: str) -> int:
@@
 return ans
-A.3 Inserting Garbage Code
-As introduced in §3.3.1, we inject three types of garbage code: (1) repeated-name global variable declarations, (2) in-executable conditional statements, and (3) dead-loop function definitions.
+B.1.3 Inserting Garbage Code (GBC)
+We inject three types of garbage code: (1) repeated-name global variable declarations, (2) unexecutable conditional statements, and (3) dead-loop function definitions.
 For (1), we randomly select a few parameters or variables and declare global variables using the same names, assigning arbitrary integers at the beginning of the program.
@@
 […]
-Code 4: Garbage death loop function templates for GBC perturbation.
+Code 5: Garbage death loop function templates for GBC perturbation.
 # Death Loop Template 1
@@
 pass
-Code 5: GBC-perturbed example from LCB (sample 36) for output prediction.
+Code 6: GBC-perturbed example from LCB (sample 36) for output prediction.
 s = 59
@@
 return ans
-A.4 Aggregating Structural Perturbations
-As introduced in §3.3.1, we aggregate the application of REN, RTF, and GBC to a single code snippet, as shown in Code 6.
-Code 6: ALL-perturbed example from LCB (sample 36) for output prediction.
-Var_1 = 24
-def f2(Var_1: str) -> int:
-def f1():
-for Var_4 in iter(lambda : True, False):
-pass
-def funct6():
-for Var_3 in iter(int, 1):
-Var_3 += 1
-Var_6 = 0
-for Var_3 in range(1, len(Var_1)):
-Var_7 = Var_1 if 0 else Var_1
-if None:
-Var_5 = Var_1
-if (lambda : Var_1[Var_3 - 1] != Var_1[Var_3])():
-Var_6 += min(Var_3, len(Var_1) - Var_3)
-while Var_1 != Var_1:
-Var_2 = Var_1
-return Var_6
-A.5 Inserting Misleading Messages
-As introduced in §3.3.2, we insert misleading comments for key AST node types, including function declarations, return statements, loops, conditionals, variable assignments, and common data structure operations.
-The data structure operations include common list, set, and string manipulations: append, extend, insert, remove, pop, sort, reverse, update, add, split, join, replace, lower, upper, capitalize, and swap case.
+B.2 Inserting Contextual-Level Misleading Messages (MCC and MPS)
+As introduced in §2.2, we insert misleading comments for eight key AST node types, including (1) function declarations, (2) return statements, (3-4) while and for loops, (5) conditionals, (6) variable assignments, and (7) common data structure operations.
+The (8) data structure operations include common list, set, and string manipulations: append, extend, insert, remove, pop, sort, reverse, update, add, split, join, replace, lower, upper, capitalize, and swap case.
 To construct these comments, we design prompts for GPT-4o to generate misleading but plausible messages for each type.
@@
 Take variable assignments as an example; Code 7 is the prompt for generating misleading comments for variable assignments, and Code 8 is a list of selected misleading comment candidates.
-Code 9 and 10 present an example of MDC and MPS perturbations.
+Code 9 and 10 present an example of MCC and MPS perturbations.
 Code 7: Prompt used to generate misleading comments for variable assignments.
@@
 […]
-Code 8: Selected misleading comments for variable assignments used in MDC and MPS perturbations, where ‘{variable}’ is the variable name.
+Code 8: Selected misleading comments for variable assignments used in MCC and MPS perturbations, where ‘{variable}’ is the variable name.
 VARIABLE_ASSIGNMENTS_COMMENT_CANDIDATE = [
@@
 "The ’{variable}’ variable is present but remains dormant throughout the function.",
-Code 9: MDC-perturbed example from LCB (sample 36) for output prediction.
+Code 9: MCC-perturbed example from LCB (sample 36) for output prediction.
 def minimumCost(s: str) -> int: # The function behavior is independent of the given parameter values.
@@
 return ans
-A.6 Inserting Incorrect Hints
-In order to generate plausible, logical but incorrect hints, we instruct GPT-4o to rewrite the correct assertion expression to incorrect ones and only allow it to modify the output value using the prompt in Code 11 for output prediction.
+B.3 Inserting Reasoning-Level Misleading Hint (MHC)
+In order to generate plausible, logical, but incorrect hints, we instruct GPT-4o to rewrite the correct assertion expression into incorrect ones and only allow it to modify the output value using the prompt in Code 11 for output prediction.
 An example perturbed instance is shown in Code 12, where the ground-truth return value 2 is replaced with the misleading hint # The return value is 3.
@@
 return ’’.join(new_text)
-Appendix B Statistical Analysis
+Appendix C Statistical Analysis
 Due to resource constraints, we generate three candidates for direct inference per query.
@@
 We present two summary tables for output prediction tasks (Table 11) and input prediction tasks (Table 12).
-The results show that across a total of $357$ measurements, only 9 cases have $\sigma>1.0$ (i.e., $0.01$ on a scale from 00 to $1$), and $57$ have $\sigma>0.5$, indicating that most scores exhibit low variance under repeated sampling.
+The results show that across a total of $357$ measurements, only 9 cases have $\sigma>1.0$ (i.e., $0.01$ on a scale from 0 to $1$), and $57$ have $\sigma>0.5$, indicating that most scores exhibit low variance under repeated sampling.
 Furthermore, our benchmark includes $1279$ problems across $8$ experiment settings on $17$ models for output prediction tasks, and $800$ problems across $5$ experiment settings on $17$ models for input prediction tasks.
-Among $241,944$ combinations, $231,726$ have identical correctness labels across all three candidates, performing an overall stability rate of $\mathbf{95.78\%}$.
+Among $241,944$ combinations, $231,726$ have identical correctness labels across all three candidates, yielding an overall stability rate of $\mathbf{95.78\%}$.
 Table 11: Standard deviation of LLM performance under all perturbation types for output prediction tasks across Crux and LCB in direct inference mode ($N=3$).
 Dark red highlights denote $\sigma>1.0$, and light red highlights denote $\sigma>0.5$.
-Model Series Model Name VAN REN RTF GBC ALL MDC MPS MHC
+Model Series Model Name VAN REN RTF GBC PSC-ALL MCC MPS MHC
 Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
@@
 Gemini Series Gemini-1.5-Flash $0.2569$ $0.1705$ $0.1559$ $0.1705$ $0.1559$ $0.1705$ $0.1021$ $0.0984$ $0.2041$ $0.1705$ $0.1559$ $0.0984$ $0.1179$ $0.2604$ $0.2125$ $0.1705$
-Gemini-2.0-Flash $0.1179$ $0.1705$ $0.3281$ $0.2604$ $0.3584$ $0.2604$ $0.2357$ $0.2952$ $0.4249$ $0.451$ $0.1768$ $0.2952$ $0.2041$ 00 $0.1559$ $0.429$
+Gemini-2.0-Flash $0.1179$ $0.1705$ $0.3281$ $0.2604$ $0.3584$ $0.2604$ $0.2357$ $0.2952$ $0.4249$ $0.451$ $0.1768$ $0.2952$ $0.2041$ 0 $0.1559$ $0.429$
 Gemini-1.5-Pro-002 $0.5103$ $0.6146$ $0.3584$ $0.3937$ $0.4449$ $0.6889$ $0.1021$ $1.135$ $0.4125$ $0.2604$ $0.27$ $0.3548$ $0.2041$ $0.3548$ $0.2041$ $0.2604$
@@
 LLaMA-3.1-405B-Instruct $0.2946$ $0.429$ $0.1559$ $0.1705$ $0.1179$ $0.3937$ $0.1768$ $0.451$ $0.7795$ $0.2604$ $0.3536$ $0.1705$ $0.3536$ $0.1968$ $0.2569$ $0.5208$
-LLaMA-3.3-70B-Instruct $0.0589$ $0.3548$ $0.2125$ $0.1968$ $0.2125$ $0.1705$ $0.2125$ 00 $0.2569$ $0.2604$ $0.1179$ $0.1968$ $0.2125$ $0.5208$ $0.0589$ $0.1968$
+LLaMA-3.3-70B-Instruct $0.0589$ $0.3548$ $0.2125$ $0.1968$ $0.2125$ $0.1705$ $0.2125$ 0 $0.2569$ $0.2604$ $0.1179$ $0.1968$ $0.2125$ $0.5208$ $0.0589$ $0.1968$
 Qwen Series Qwen2.5-7B-Instruct $0.1559$ $0.0984$ $0.27$ $0.5479$ $0.4677$ $0.1705$ $0.2125$ $0.2604$ $0.1021$ $0.1705$ $0.1559$ $0.0984$ $0.4125$ $0.6818$ $0.2041$ $0.451$
@@
 Qwen2.5-72B-Instruct $0.1179$ $0.451$ $0.1559$ $0.0984$ $0.2125$ $0.451$ $0.27$ $0.5479$ $0.4714$ $0.2604$ $0.3536$ $1.028$ $1.327$ $0.8523$ $0.1559$ $0.3937$
-Qwen2.5-Coder-32B-Instruct $0.4602$ $0.2604$ $0.4714$ $0.5905$ $0.1768$ $0.858$ $0.5803$ $0.1705$ $0.2041$ $0.7873$ $0.4249$ $0.2952$ 00 $0.3409$ $0.3118$ $0.6889$
+Qwen2.5-Coder-32B-Instruct $0.4602$ $0.2604$ $0.4714$ $0.5905$ $0.1768$ $0.858$ $0.5803$ $0.1705$ $0.2041$ $0.7873$ $0.4249$ $0.2952$ 0 $0.3409$ $0.3118$ $0.6889$
 Table 12: Standard deviation of LLM performance under four perturbation types for input prediction tasks across Crux in direct inference mode ($N=3$). Red highlights $\sigma>0.5$.
-Model Series Model Name VAN ALL MDC MPS MHC
+Model Series Model Name VAN PSC-ALL MCC MPS MHC
 GPT-4 Omni GPT-4o-Mini $0.1021$ $0.8498$ $0.5621$ $0.4602$ $0.2569$
@@
 Gemini-2.0-Flash $0.27$ $0.5893$ $0.5035$ $0.5683$ $0.1179$
-Gemini-1.5-Pro-002 $0.27$ $0.4082$ $0.4677$ $0.1559$ 00
+Gemini-1.5-Pro-002 $0.27$ $0.4082$ $0.4677$ $0.1559$ 0
 DeepSeek DeepSeek-V3 $0.3584$ $0.1021$ $0.5803$ $0.6208$ $0.766$
@@
 Qwen2.5-14B-Instruct $0.368$ $0.1179$ $0.1768$ $0.5803$ $0.2041$
-Qwen2.5-32B-Instruct $0.1021$ $0.3584$ $5890.0$ $0.2125$
+Qwen2.5-32B-Instruct $0.1021$ $0.3584$ $0.0589$ 0 $0.2125$
 Qwen2.5-72B-Instruct $0.6152$ $0.1768$ $0.5137$ $0.0589$ $0.4082$
 Qwen2.5-Coder-32B-Instruct $0.27$ $0.4125$ $0.4449$ $0.1559$ $0.2946$
-Appendix C Prompts
+Appendix D Prompts
 Our prompt instructions for LLMs follow the format designed by Crux, which is also adopted by LCB, utilizing a two-shot prompt for direct inference and a single-shot prompt for CoT inference to ensure models can follow the instruction and output format effectively.
@@
 Therefore, applying CoT prompts may lead to redundant reasoning and an unfair comparison with standard LLMs under CoT prompting.
-Considering the fairness, we use direct inference prompts for RM experiments in §4.3.2.
+Considering the fairness, we use direct inference prompts for LRM experiments in §3.4.
 Unlike LCB, we rename all example function names to f to avoid prompting models to capture the NL cues from the example.
@@
 We also include “Ensure the provided expression syntax is correct” to enforce consistent syntax.
-However, some responses from models like QwQ-32B, DeepSeek-R1 still generate syntactically incorrect outputs such as [ANSWER] f("a") == [/ANSWER] instead of [ANSWER] f("a") == "" [/ANSWER].
+However, some responses from models like QwQ-32B, DeepSeek-R1, still generate syntactically incorrect outputs such as [ANSWER] f("a") == [/ANSWER] instead of [ANSWER] f("a") == "" [/ANSWER].
 For those situations, we identify them as “Syntax Errors” and incorrect.
 We provide the complete prompts used for output prediction in both direct and CoT inference (Code 15 and 16) and for input prediction in both direct and CoT inference (Code 17 and 18).
-C.1 Output Prediction (Direct)
+D.1 Output Prediction (Direct)
 Code 15: Output Prediction Prompt with Direct Inference.
@@
 The output value for f("x9j")‘ is "x9ja", then output your prediction [ANSWER] f("x9j") == "x9ja" [/ANSWER].
-C.2 Output Prediction (CoT)
+D.2 Output Prediction (CoT)
 Code 16: Output Prediction Prompt with CoT-based Inference.
@@
 Thus, the output value for f("hi")‘ is "bhihia", then output your prediction [ANSWER] f("hi") == "bhihia" [/ANSWER].
-C.3 Input Prediction (Direct)
+D.3 Input Prediction (Direct)
 Code 17: Input Prediction Prompt with Direct Inference.
@@
 The input arguments for f that result in the output value of ‘"banana"‘ are ‘"ba", "nana"‘. Then, output your prediction [ANSWER] f("ba", "nana") == "banana" [/ANSWER].
-C.4 Input Prediction (CoT)
+D.4 Input Prediction (CoT)
 Code 18: Input Prediction Prompt with CoT-based Inference.
@@
 Thus, the input arguments for the function f that result in the output value ‘17‘ is ‘16‘. Then, output your prediction [ANSWER] f(16) == 17 [/ANSWER].
-Appendix D Supplementary Experimental Results
-D.1 Impact on Output Prediction
-In §4.2, we evaluate the robustness of LLMs on output prediction tasks under direct inference.
-Table 1 reports the absolute Pass@1 degradation across all perturbations.
-We use absolute values to highlight the overall impact of each perturbation and conduct cross-perturbation comparisons.
-For completeness, we also include the relative Pass@1 degradation in Table 10, reflecting the percentage drop from each model’s vanilla performance.
-Table 10: Relative Pass@1 degradation of LLMs under all perturbation types for output prediction tasks across Crux and LCB in direct inference ($N=3$). [To Absolute Version]
-Model Series Model Name VAN REN RTF GBC ALL MDC MPS MHC
-Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
-GPT-4 Omni GPT-4o-Mini $57.29$ $53.24$ $-1.09\%$ $-18.17\%$ $-3.13\%$ $-1.83\%$ $-19.05\%$ $-10.33\%$ $-27.85\%$ $-31.90\%$ $-32.58\%$ $-40.39\%$ $-23.20\%$ $-27.97\%$ $-11.56\%$ $-44.31\%$
-GPT-4o $71.33$ $64.51$ $-3.80\%$ $-14.89\%$ $-4.73\%$ $-7.12\%$ $-12.91\%$ $-4.96\%$ $-15.01\%$ $-28.37\%$ $-13.96\%$ $-29.13\%$ $-17.23\%$ $-24.27\%$ $-6.37\%$ $-26.65\%$
-Claude Series Claude-3.5-Haiku-20241022 $57.42$ $58.25$ $-6.89\%$ $-16.01\%$ $-3.12\%$ $+0.60\%$ $-12.70\%$ $-7.77\%$ $-24.02\%$ $-36.92\%$ $-13.79\%$ $-10.04\%$ $-11.25\%$ $-8.00\%$ $-27.50\%$ $-53.29\%$
-Claude-3.5-Sonnet-20241022 $71.46$ $73.83$ $-2.33\%$ $-22.71\%$ $-1.28\%$ $+1.60\%$ $-9.62\%$ $-6.69\%$ $-14.75\%$ $-34.12\%$ $-8.10\%$ $-9.61\%$ $-8.63\%$ $-10.65\%$ $-14.40\%$ $-43.45\%$
-Gemini Series Gemini-1.5-Flash $56.21$ $44.26$ $-6.30\%$ $-0.94\%$ $-2.08\%$ $-2.36\%$ $-8.15\%$ $-0.79\%$ $-18.38\%$ $-12.74\%$ $-14.75\%$ $-32.23\%$ $-28.47\%$ $-40.41\%$ $-7.34\%$ $-22.17\%$
-Gemini-2.0-Flash $65.04$ $65.97$ $-3.59\%$ $-27.43\%$ $-3.20\%$ $-1.16\%$ $-23.83\%$ $-13.29\%$ $-33.25\%$ $-42.09\%$ $-26.78\%$ $-51.90\%$ $-33.12\%$ $-50.32\%$ $-8.26\%$ $-19.51\%$
-Gemini-1.5-Pro-002 $67.37$ $56.16$ $-4.33\%$ $-6.07\%$ $-3.71\%$ $-8.43\%$ $-8.35\%$ $-6.82\%$ $-19.91\%$ $-23.17\%$ $-23.01\%$ $-32.96\%$ $-21.71\%$ $-32.84\%$ $-10.02\%$ $-32.47\%$
-DeepSeek DeepSeek-V3 $67.87$ $67.78$ $-3.93\%$ $-24.13\%$ $-1.96\%$ $-5.54\%$ $-6.81\%$ $-8.83\%$ $-12.89\%$ $-35.63\%$ $-16.57\%$ $-41.38\%$ $-10.13\%$ $-34.19\%$ $-10.74\%$ $-29.67\%$
-LLaMA Series LLaMA-3.1-8B-Instruct $35.96$ $34.66$ $+0.23\%$ $-12.45\%$ $-0.93\%$ $-4.82\%$ $-8.69\%$ $-13.65\%$ $-23.75\%$ $-19.68\%$ $-22.48\%$ $-28.51\%$ $-17.84\%$ $-23.09\%$ $-16.69\%$ $-39.16\%$
-LLaMA-3.1-70B-Instruct $56.08$ $44.89$ $-4.23\%$ $-3.10\%$ $-0.97\%$ $-3.88\%$ $-9.58\%$ $-5.43\%$ $-19.61\%$ $-17.21\%$ $-19.24\%$ $-27.91\%$ $-26.67\%$ $-38.76\%$ $-8.40\%$ $-32.25\%$
-LLaMA-3.1-405B-Instruct $63.54$ $50.66$ $-8.46\%$ $-7.69\%$ $-0.98\%$ $-8.79\%$ $-9.11\%$ $-5.63\%$ $-19.67\%$ $-16.90\%$ $-6.56\%$ $-14.70\%$ $-11.28\%$ $-19.92\%$ $-6.89\%$ $-25.69\%$
-LLaMA-3.3-70B-Instruct $59.92$ $48.50$ $-3.69\%$ $-2.87\%$ $-2.09\%$ $-7.03\%$ $-9.18\%$ $-11.76\%$ $-17.32\%$ $-20.09\%$ $-12.45\%$ $-19.80\%$ $-13.49\%$ $-25.39\%$ $-6.82\%$ $-20.95\%$
-Qwen Series Qwen2.5-7B-Instruct $43.29$ $41.41$ $-7.03\%$ $-17.48\%$ $-2.98\%$ $-7.73\%$ $-29.93\%$ $-9.08\%$ $-37.92\%$ $-30.92\%$ $-58.04\%$ $-38.32\%$ $-45.24\%$ $-19.83\%$ $-26.95\%$ $-55.63\%$
-Qwen2.5-14B-Instruct $47.79$ $49.48$ $-3.84\%$ $-19.97\%$ $-1.39\%$ $-2.81\%$ $-23.45\%$ $-6.89\%$ $-39.84\%$ $-29.96\%$ $-34.87\%$ $-41.35\%$ $-32.52\%$ $-33.90\%$ $-9.33\%$ $-22.22\%$
-Qwen2.5-32B-Instruct $60.04$ $59.64$ $-2.85\%$ $-18.09\%$ $-4.30\%$ $-0.82\%$ $-9.65\%$ $-2.33\%$ $-19.78\%$ $-34.89\%$ $-18.04\%$ $-32.67\%$ $-23.66\%$ $-23.57\%$ $-13.12\%$ $-30.81\%$
-Qwen2.5-72B-Instruct $60.08$ $54.91$ $-1.73\%$ $-15.46\%$ $-3.12\%$ $+0.76\%$ $-17.82\%$ $-3.68\%$ $-23.30\%$ $-25.22\%$ $-16.99\%$ $-12.67\%$ $-24.06\%$ $-26.24\%$ $-16.02\%$ $-40.56\%$
-Qwen2.5-Coder-32B-Instruct $66.96$ $56.65$ $-3.73\%$ $-14.86\%$ $-1.43\%$ $-3.56\%$ $-16.05\%$ $-4.91\%$ $-22.34\%$ $-27.89\%$ $-23.21\%$ $-36.61\%$ $-16.55\%$ $-30.34\%$ $-10.45\%$ $-21.74\%$
-Average $-3.98\%$ $-14.25\%$ $-2.44\%$ $-3.70\%$ $-13.82\%$ $-7.23\%$ $-22.92\%$ $-27.51\%$ $-21.26\%$ $-29.42\%$ $-21.47\%$ $-27.63\%$ $-12.40\%$ $-32.97\%$
-$\mathbf{-7.83}\%$ $\mathbf{-2.91}\%$ $\mathbf{-11.35}\%$ $\mathbf{-24.64}\%$ $\mathbf{-24.32}\%$ $\mathbf{-23.78}\%$ $\mathbf{-20.11}\%$
-D.2 Robustness Across Model Scales
-In §4.2.3, we compare the robustness against different model scales.
-To fairly aggregate model robustness across different perturbation types, we compute a weighted average of the relative Pass@1 degradation under four representative perturbations: ALL (structural), MDC and MPS (misleading messages), and MHC (hints).
-We select ALL as the representative structural perturbation because of the compounding effects of REN, RTF, and GBC captured in §4.2.2, which reflect the most challenging structural setting.
-Since MDC and MPS stem from the same perturbation strategy and often target overlapping AST node types, we assign them lower weights ($1/8$ each) to avoid over-representing similar effects.
-Overall, we define a weighted robustness metric with weights of $1/2$, $1/8$, $1/8$, and $1/4$ for ALL, MDC, MPS, and MHC, respectively, and summarize the results in Table 2.
-For completeness, we also report the results under uniform weighting ($1/4$ for each) in Table 11, which shows similar trends and confirms the stability of our findings in §4.2.3, as well as their absolute degradation versions in Table 12 and 13, respectively.
-Table 11: Relative Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/4 ALL, 1/4 MDC, 1/4 MPS, and 1/4 MHC. [To Absolute Version]
-Model Series Model Name VAN ALL ($1/4$) MDC ($1/4$) MPS ($1/4$) MHC ($1/4$) Average
-GPT-4 Omni GPT-4o-Mini 55.77 $-29.37\%$ $-35.51\%$ $-24.99\%$ $-23.83\%$ $-28.42\%$
-GPT-4o 68.78 $-20.01\%$ $-19.64\%$ $-19.87\%$ $-13.96\%$ $-18.37\%$
-Claude Series Claude-3.5-Haiku-20241022 57.73 $-28.85\%$ $-12.38\%$ $-10.03\%$ $-37.16\%$ $-22.11\%$
-Claude-3.5-Sonnet-20241022 72.35 $-22.01\%$ $-8.67\%$ $-9.39\%$ $-25.28\%$ $-16.34\%$
-Gemini Series Gemini-1.5-Flash 51.73 $-16.27\%$ $-21.30\%$ $-32.94\%$ $-12.89\%$ $-20.85\%$
-Gemini-2.0-Flash 65.39 $-36.56\%$ $-36.19\%$ $-39.56\%$ $-12.48\%$ $-31.20\%$
-Gemini-1.5-Pro-002 63.17 $-21.13\%$ $-26.73\%$ $-25.88\%$ $-18.43\%$ $-23.04\%$
-DeepSeek DeepSeek-V3 67.84 $-21.41\%$ $-25.86\%$ $-19.14\%$ $-17.83\%$ $-21.06\%$
-LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-22.23\%$ $-24.74\%$ $-19.81\%$ $-25.10\%$ $-22.97\%$
-LLaMA-3.1-70B-Instruct 51.89 $-18.71\%$ $-22.49\%$ $-31.20\%$ $-17.33\%$ $-22.43\%$
-LLaMA-3.1-405B-Instruct 58.72 $-18.63\%$ $-9.61\%$ $-14.51\%$ $-13.93\%$ $-14.17\%$
-LLaMA-3.3-70B-Instruct 55.64 $-18.35\%$ $-15.20\%$ $-17.95\%$ $-12.11\%$ $-15.90\%$
-Qwen Series Qwen2.5-7B-Instruct 42.59 $-35.30\%$ $-50.65\%$ $-35.72\%$ $-37.69\%$ $-39.84\%$
-Qwen2.5-14B-Instruct 48.42 $-36.14\%$ $-37.30\%$ $-33.04\%$ $-14.16\%$ $-30.16\%$
-Qwen2.5-32B-Instruct 59.89 $-25.44\%$ $-23.52\%$ $-23.63\%$ $-19.74\%$ $-23.08\%$
-Qwen2.5-72B-Instruct 58.14 $-24.02\%$ $-15.37\%$ $-24.88\%$ $-25.21\%$ $-22.37\%$
-Qwen2.5-Coder-32B-Instruct 63.10 $-24.42\%$ $-28.23\%$ $-21.72\%$ $-14.68\%$ $-22.26\%$
-Table 12: Absolute Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/2 ALL, 1/8 MDC, 1/8 MPS, and 1/4 MHC. [To Relative Version]
-Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Average
-GPT-4 Omni GPT-4o-Mini 55.77 $-16.34$ $-19.73$ $-13.89$ $-12.98$ $-15.62$
-GPT-4o 68.78 $-13.55$ $-13.27$ $-13.55$ $-9.28$ $-12.45$
-Claude Series Claude-3.5-Haiku-20241022 57.73 $-16.68$ $-7.14$ $-5.79$ $-21.50$ $-15.33$
-Claude-3.5-Sonnet-20241022 72.35 $-16.03$ $-6.28$ $-6.80$ $-18.45$ $-14.26$
-Gemini Series Gemini-1.5-Flash 51.73 $-8.57$ $-10.53$ $-16.71$ $-6.25$ $-9.26$
-Gemini-2.0-Flash 65.39 $-23.92$ $-23.72$ $-25.91$ $-8.18$ $-20.21$
-Gemini-1.5-Pro-002 63.17 $-13.27$ $-16.63$ $-16.05$ $-11.05$ $-13.48$
-DeepSeek DeepSeek-V3 67.84 $-14.52$ $-17.54$ $-12.98$ $-12.09$ $-14.10$
-LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-7.90$ $-8.76$ $-7.01$ $-8.84$ $-8.13$
-LLaMA-3.1-70B-Instruct 51.89 $-9.77$ $-11.44$ $-15.87$ $-8.37$ $-10.39$
-LLaMA-3.1-405B-Instruct 58.72 $-11.02$ $-5.39$ $-8.26$ $-7.61$ $-9.12$
-LLaMA-3.3-70B-Instruct 55.64 $-10.14$ $-8.26$ $-9.67$ $-6.36$ $-8.90$
-Qwen Series Qwen2.5-7B-Instruct 42.59 $-15.06$ $-21.66$ $-15.32$ $-15.92$ $-16.14$
-Qwen2.5-14B-Instruct 48.42 $-17.46$ $-18.09$ $-16.00$ $-6.91$ $-14.72$
-Qwen2.5-32B-Instruct 59.89 $-15.22$ $-14.07$ $-14.15$ $-11.81$ $-14.09$
-Qwen2.5-72B-Instruct 58.14 $-13.94$ $-8.99$ $-14.44$ $-14.36$ $-13.49$
-Qwen2.5-Coder-32B-Instruct 63.10 $-15.27$ $-17.49$ $-13.37$ $-8.99$ $-13.74$
-Table 13: Absolute Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/4 ALL, 1/4 MDC, 1/4 MPS, and 1/4 MHC. [To Relative Version]
-Model Series Model Name VAN ALL ($1/4$) MDC ($1/4$) MPS ($1/4$) MHC ($1/4$) Average
-GPT-4 Omni GPT-4o-Mini 55.77 $-16.34$ $-19.73$ $-13.89$ $-12.98$ $-15.73$
-GPT-4o 68.78 $-13.55$ $-13.27$ $-13.55$ $-9.28$ $-12.41$
-Claude Series Claude-3.5-Haiku-20241022 57.73 $-16.68$ $-7.14$ $-5.79$ $-21.50$ $-12.78$
-Claude-3.5-Sonnet-20241022 72.35 $-16.03$ $-6.28$ $-6.80$ $-18.45$ $-11.89$
-Gemini Series Gemini-1.5-Flash 51.73 $-8.57$ $-10.53$ $-16.71$ $-6.25$ $-10.52$
-Gemini-2.0-Flash 65.39 $-23.92$ $-23.72$ $-25.91$ $-8.18$ $-20.43$
-Gemini-1.5-Pro-002 63.17 $-13.27$ $-16.63$ $-16.05$ $-11.05$ $-14.25$
-DeepSeek DeepSeek-V3 67.84 $-14.52$ $-17.54$ $-12.98$ $-12.09$ $-14.28$
-LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-7.90$ $-8.76$ $-7.01$ $-8.84$ $-8.12$
-LLaMA-3.1-70B-Instruct 51.89 $-9.77$ $-11.44$ $-15.87$ $-8.37$ $-11.36$
-LLaMA-3.1-405B-Instruct 58.72 $-11.02$ $-5.39$ $-8.26$ $-7.61$ $-8.07$
-LLaMA-3.3-70B-Instruct 55.64 $-10.14$ $-8.26$ $-9.67$ $-6.36$ $-8.61$
-Qwen Series Qwen2.5-7B-Instruct 42.59 $-15.06$ $-21.66$ $-15.32$ $-15.92$ $-16.99$
-Qwen2.5-14B-Instruct 48.42 $-17.46$ $-18.09$ $-16.00$ $-6.91$ $-14.61$
-Qwen2.5-32B-Instruct 59.89 $-15.22$ $-14.07$ $-14.15$ $-11.81$ $-13.81$
-Qwen2.5-72B-Instruct 58.14 $-13.94$ $-8.99$ $-14.44$ $-14.36$ $-12.93$
-Qwen2.5-Coder-32B-Instruct 63.10 $-15.27$ $-17.49$ $-13.37$ $-8.99$ $-13.78$
-D.3 Effect of CoT Reasoning
-In §4.3.1, we analyze the effect of CoT prompting on model robustness using relative degradation aggregated across all perturbations (Table 3).
+Appendix E Supplementary Experimental Results
+E.1 Additional Performance Degradation Tables for Output Prediction (Direct Inference)
+In §3.2, we evaluate the robustness of LLMs on output prediction tasks under direct inference, using the relative Pass@1 degradation in Table 1.
+For completeness, we also include the absolute Pass@1 degradation in Table 13, reflecting the percentage drop from each model’s vanilla performance.
+Table 13: Absolute Pass@1 degradation of LLMs under all perturbation types for output prediction tasks on Crux and LCB using direct inference. [Relative Version] for the relative degradation.
+Model Series Model Name VAN PSC-ALL MCC MPS MHC Average
+Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
+GPT-4 Omni GPT-4o-Mini $57.3$ $53.2$ 16.0-16.0 17.0-17.0 18.7-18.7 21.5-21.5 13.3-13.3 14.9-14.9 6.6-6.6 23.6-23.6 15.7-15.7
+GPT-4o $71.3$ $64.5$ 10.7-10.7 18.3-18.3 10.0-10.0 18.8-18.8 12.3-12.3 15.7-15.7 4.5-4.5 17.2-17.2 12.4-12.4
+Claude Series Claude-3.5-Haiku-20241022 $57.4$ $58.2$ 13.8-13.8 21.5-21.5 7.9-7.9 5.8-5.8 6.5-6.5 4.7-4.7 15.8-15.8 31.0-31.0 12.8-12.8
+Claude-3.5-Sonnet-20241022 $71.5$ $73.8$ 10.5-10.5 25.2-25.2 5.8-5.8 7.1-7.1 6.2-6.2 7.9-7.9 10.3-10.3 32.1-32.1 11.9-11.9
+Gemini Series Gemini-1.5-Flash $56.2$ $44.3$ 10.3-10.3 5.6-5.6 8.3-8.3 14.3-14.3 16.0-16.0 17.9-17.9 4.1-4.1 9.8-9.8 10.5-10.5
+Gemini-2.0-Flash $65.0$ $66.0$ 21.6-21.6 27.8-27.8 17.4-17.4 34.2-34.2 21.5-21.5 33.2-33.2 5.4-5.4 12.9-12.9 20.4-20.4
+Gemini-1.5-Pro-002 $67.4$ $56.2$ 13.4-13.4 13.0-13.0 15.5-15.5 18.5-18.5 14.6-14.6 18.4-18.4 6.7-6.7 18.2-18.2 14.2-14.2
+DeepSeek DeepSeek-V3 $67.9$ $67.8$ 8.7-8.7 24.1-24.1 11.2-11.2 28.0-28.0 6.9-6.9 23.2-23.2 7.3-7.3 20.1-20.1 14.3-14.3
+LLaMA Series LLaMA-3.1-8B-Instruct $36.0$ $34.7$ 8.5-8.5 6.8-6.8 8.1-8.1 9.9-9.9 6.4-6.4 8.0-8.0 6.0-6.0 13.6-13.6 8.1-8.1
+LLaMA-3.1-70B-Instruct $56.1$ $44.9$ 11.0-11.0 7.7-7.7 10.8-10.8 12.5-12.5 15.0-15.0 17.4-17.4 4.7-4.7 14.5-14.5 11.4-11.4
+LLaMA-3.1-405B-Instruct $63.5$ $50.7$ 12.5-12.5 8.6-8.6 4.2-4.2 7.4-7.4 7.2-7.2 10.1-10.1 4.4-4.4 13.0-13.0 8.1-8.1
+LLaMA-3.3-70B-Instruct $59.9$ $48.5$ 10.4-10.4 9.7-9.7 7.5-7.5 9.6-9.6 8.1-8.1 12.3-12.3 4.1-4.1 10.2-10.2 8.6-8.6
+Qwen Series Qwen2.5-7B-Instruct $43.3$ $41.4$ 16.4-16.4 12.8-12.8 25.1-25.1 15.9-15.9 19.6-19.6 8.2-8.2 11.7-11.7 23.0-23.0 17.0-17.0
+Qwen2.5-14B-Instruct $47.8$ $49.5$ 19.0-19.0 14.8-14.8 16.7-16.7 20.5-20.5 15.5-15.5 16.8-16.8 4.5-4.5 11.0-11.0 14.6-14.6
+Qwen2.5-32B-Instruct $60.0$ $59.6$ 11.9-11.9 20.8-20.8 10.8-10.8 19.5-19.5 14.2-14.2 14.1-14.1 7.9-7.9 18.4-18.4 13.8-13.8
+Qwen2.5-72B-Instruct $60.1$ $54.9$ 14.0-14.0 13.8-13.8 10.2-10.2 7.0-7.0 14.5-14.5 14.4-14.4 9.6-9.6 22.3-22.3 12.9-12.9
+Qwen2.5-Coder-32B-Instruct $67.0$ $56.6$ 15.0-15.0 15.8-15.8 15.5-15.5 20.7-20.7 11.1-11.1 17.2-17.2 7.0-7.0 12.3-12.3 13.8-13.8
+Average 13.2-13.2 15.5-15.5 12.0-12.0 16.0-16.0 12.3-12.3 15.0-15.0 7.1-7.1 17.8-17.8
+$\mathbf{-14.0}$ $\mathbf{-13.5}$ $\mathbf{-13.3}$ $\mathbf{-11.1}$ $\mathbf{-13.0}$
+E.2 Additional Performance Degradation Tables for Output Prediction (CoT Inference)
+In §3.3, we analyze the effect of CoT prompting on model robustness using relative degradation aggregated across all perturbations (Table 2).
 We adopt the relative metric to mitigate the large variance between vanilla performances in direct and CoT inference modes, providing a fair comparison of performance degradation.
-For completeness, we also report the absolute degradation in Table 14, as well as the relative and absolute degradation on both Crux and LCB in Tables 15 and 16, respectively.
-Table 14: Comparison of absolute Pass@1 degradation aggregated over Crux and LCB for output prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Relative Version]
-Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average
+For completeness, we report the absolute degradation in Table 14, as well as the relative and absolute degradation on both Crux and LCB in Tables 15 and 16, respectively.
+Table 14: Comparison of absolute Pass@1 degradation aggregated over Crux and LCB for output prediction in different reasoning mode. See [Relative Version] for the relative degradation.
+Model Series Model Name VAN PSC-ALL MCC MPS MHC Average
 Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT
-GPT-4 Omni GPT-4o-Mini $55.77$ $81.47$ $-16.34$ $-10.71$ $-19.73$ $-8.84$ $-13.89$ $-8.91$ $-12.98$ $-2.19$ $-15.62$ $-7.75$
-GPT-4o $68.78$ $91.79$ $-13.55$ $-4.53$ $-13.27$ $-4.61$ $-13.55$ $-5.16$ $-9.28$ $-1.25$ $-12.45$ $-3.87$
-Claude Series Claude-3.5-Haiku-20241022 $57.73$ $72.95$ $-16.68$ $-15.56$ $-7.14$ $-7.74$ $-5.79$ $-6.33$ $-21.50$ $-10.40$ $-15.33$ $-10.43$
-Claude-3.5-Sonnet-20241022 $72.35$ $86.00$ $-16.03$ $-6.49$ $-6.28$ $-4.53$ $-6.80$ $-5.47$ $-18.45$ $-6.80$ $-14.26$ $-5.90$
-Gemini Series Gemini-1.5-Flash $51.73$ $75.22$ $-8.57$ $-13.76$ $-10.53$ $-16.34$ $-16.71$ $-31.74$ $-6.25$ $-1.56$ $-9.26$ $-15.34$
-Gemini-2.0-Flash $65.39$ $89.13$ $-23.92$ $-5.55$ $-23.72$ $-5.63$ $-25.91$ $-12.74$ $-8.18$ $-1.80$ $-20.21$ $-6.25$
-Gemini-1.5-Pro-002 $63.17$ $87.18$ $-13.27$ $-6.49$ $-16.63$ $-10.40$ $-16.05$ $-12.82$ $-11.05$ $-3.36$ $-13.48$ $-8.01$
-DeepSeek DeepSeek-V3 $67.84$ $89.52$ $-14.52$ $-8.91$ $-17.54$ $-15.87$ $-12.98$ $-16.81$ $-12.09$ $-3.13$ $-14.10$ $-10.81$
-LLaMA Series LLaMA-3.1-8B-Instruct $35.47$ $44.72$ $-7.90$ $-12.35$ $-8.76$ $-9.38$ $-7.01$ $-9.62$ $-8.84$ $-3.91$ $-8.13$ $-8.96$
-LLaMA-3.1-70B-Instruct $51.89$ $69.43$ $-9.77$ $-12.43$ $-11.44$ $-16.11$ $-15.87$ $-21.42$ $-8.37$ $-4.69$ $-10.39$ $-13.35$
-LLaMA-3.1-405B-Instruct $58.72$ $78.42$ $-11.02$ $-10.63$ $-5.39$ $-8.05$ $-8.26$ $-12.20$ $-7.61$ $-5.86$ $-9.12$ $-9.21$
-LLaMA-3.3-70B-Instruct $55.64$ $76.86$ $-10.14$ $-8.37$ $-8.26$ $-5.94$ $-9.67$ $-8.52$ $-6.36$ $-3.67$ $-8.90$ $-6.68$
-Qwen Series Qwen2.5-7B-Instruct $42.59$ $58.17$ $-15.06$ $-12.67$ $-21.66$ $-12.74$ $-15.32$ $-16.26$ $-15.92$ $-4.07$ $-16.14$ $-11.34$
-Qwen2.5-14B-Instruct $48.42$ $70.68$ $-17.46$ $-16.34$ $-18.09$ $-24.94$ $-16.00$ $-30.65$ $-6.91$ $-7.66$ $-14.72$ $-19.33$
-Qwen2.5-32B-Instruct $59.89$ $79.44$ $-15.22$ $-13.84$ $-14.07$ $-14.78$ $-14.15$ $-19.23$ $-11.81$ $-4.22$ $-14.09$ $-12.86$
-Qwen2.5-72B-Instruct $58.14$ $82.80$ $-13.94$ $-10.48$ $-8.99$ $-5.63$ $-14.44$ $-13.45$ $-14.36$ $-5.47$ $-13.49$ $-8.80$
-Qwen2.5-Coder-32B-Instruct $63.10$ $84.75$ $-15.27$ $-12.43$ $-17.49$ $-16.11$ $-13.37$ $-16.89$ $-8.99$ $-7.90$ $-13.74$ $-13.13$
-Average $\mathbf{-14.04}$ $\mathbf{-10.68}$ $\mathbf{-13.47}$ $\mathbf{-11.04}$ $\mathbf{-13.28}$ $\mathbf{-14.60}$ $\mathbf{-11.11}$ $\mathbf{-4.58}$ $\mathbf{-13.14}$ $\mathbf{-10.12}$
-Table 15: Relative Pass@1 degradation of LLMs under all perturbation types for output prediction tasks across Crux and LCB in CoT inference ($N=1$). [To Absolute Version]
-Model Series Model Name VAN ALL MDC MPS MHC
+GPT-4 Omni GPT-4o-Mini $55.8$ $81.5$ 29.4-29.4 13.1-13.1 35.5-35.5 10.8-10.8 25.0-25.0 11.0-11.0 23.8-23.8 2.7-2.7 28.4-28.4 9.4-9.4
+GPT-4o $68.8$ $91.8$ 20.0-20.0 4.9-4.9 19.6-19.6 5.0-5.0 19.9-19.9 5.6-5.6 14.0-14.0 1.4-1.4 18.4-18.4 4.2-4.2
+Claude Series Claude-3.5-Haiku-20241022 $57.7$ $72.9$ 28.9-28.9 21.2-21.2 12.4-12.4 10.6-10.6 10.0-10.0 8.7-8.7 37.2-37.2 14.2-14.2 22.1-22.1 13.7-13.7
+Claude-3.5-Sonnet-20241022 $72.3$ $86.0$ 22.0-22.0 7.4-7.4 8.7-8.7 5.3-5.3 9.4-9.4 6.3-6.3 25.3-25.3 7.8-7.8 16.3-16.3 6.7-6.7
+Gemini Series Gemini-1.5-Flash $51.7$ $75.2$ 16.3-16.3 18.3-18.3 21.3-21.3 21.6-21.6 32.9-32.9 42.1-42.1 12.9-12.9 2.1-2.1 20.8-20.8 21.0-21.0
+Gemini-2.0-Flash $65.4$ $89.1$ 36.6-36.6 6.2-6.2 36.2-36.2 6.3-6.3 39.6-39.6 14.1-14.1 12.5-12.5 2.0-2.0 31.2-31.2 7.1-7.1
+Gemini-1.5-Pro-002 $63.2$ $87.2$ 21.1-21.1 7.4-7.4 26.7-26.7 11.9-11.9 25.9-25.9 14.6-14.6 18.4-18.4 3.9-3.9 23.0-23.0 9.4-9.4
+DeepSeek DeepSeek-V3 $67.8$ $89.5$ 21.4-21.4 9.9-9.9 25.9-25.9 17.6-17.6 19.1-19.1 18.7-18.7 17.8-17.8 3.5-3.5 21.1-21.1 12.4-12.4
+LLaMA Series LLaMA-3.1-8B-Instruct $35.5$ $44.7$ 22.2-22.2 27.6-27.6 24.7-24.7 21.1-21.1 19.8-19.8 21.4-21.4 25.1-25.1 9.0-9.0 23.0-23.0 19.8-19.8
+LLaMA-3.1-70B-Instruct $51.9$ $69.4$ 18.7-18.7 18.0-18.0 22.5-22.5 23.3-23.3 31.2-31.2 31.1-31.1 17.3-17.3 6.8-6.8 22.4-22.4 19.8-19.8
+LLaMA-3.1-405B-Instruct $58.7$ $78.4$ 18.6-18.6 13.6-13.6 9.6-9.6 10.3-10.3 14.5-14.5 15.6-15.6 13.9-13.9 7.5-7.5 14.2-14.2 11.8-11.8
+LLaMA-3.3-70B-Instruct $55.6$ $76.9$ 18.4-18.4 10.8-10.8 15.2-15.2 7.7-7.7 17.9-17.9 11.0-11.0 12.1-12.1 4.7-4.7 15.9-15.9 8.6-8.6
+Qwen Series Qwen2.5-7B-Instruct $42.6$ $58.2$ 35.3-35.3 21.7-21.7 50.7-50.7 22.0-22.0 35.7-35.7 28.0-28.0 37.7-37.7 7.0-7.0 39.8-39.8 19.7-19.7
+Qwen2.5-14B-Instruct $48.4$ $70.7$ 36.1-36.1 22.9-22.9 37.3-37.3 35.3-35.3 33.0-33.0 43.4-43.4 14.2-14.2 10.9-10.9 30.2-30.2 28.1-28.1
+Qwen2.5-32B-Instruct $59.9$ $79.4$ 25.4-25.4 17.3-17.3 23.5-23.5 18.5-18.5 23.6-23.6 24.1-24.1 19.7-19.7 5.4-5.4 23.1-23.1 16.3-16.3
+Qwen2.5-72B-Instruct $58.1$ $82.8$ 24.0-24.0 12.3-12.3 15.4-15.4 6.9-6.9 24.9-24.9 16.3-16.3 25.2-25.2 6.5-6.5 22.4-22.4 10.5-10.5
+Qwen2.5-Coder-32B-Instruct $63.1$ $84.8$ 24.4-24.4 14.6-14.6 28.2-28.2 19.0-19.0 21.7-21.7 19.9-19.9 14.7-14.7 9.3-9.3 22.3-22.3 15.7-15.7
+Average 24.6-24.6 14.5-14.5 24.3-24.3 14.9-14.9 23.8-23.8 19.5-19.5 20.1-20.1 6.2-6.2 23.2-23.2 13.8-13.8
+$\mathbf{-20.9}$ $\mathbf{-20.8}$ $\mathbf{-22.2}$ $\mathbf{-14.9}$ $\mathbf{-19.7}$
+Table 15: Relative Pass@1 degradation of LLMs under all perturbations for output prediction tasks across Crux and LCB in CoT inference ($N=1$). [To Absolute Version].
+Model Series Model Name VAN PSC-ALL MCC MPS MHC Average
 Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
-GPT-4 Omni GPT-4o-Mini $79.88$ $84.13$ $-10.80\%$ $-16.87\%$ $-10.80\%$ $-10.92\%$ $-13.15\%$ $-7.44\%$ $-3.29\%$ $-1.74\%$
-GPT-4o $90.25$ $94.36$ $-2.77\%$ $-8.41\%$ $-3.60\%$ $-7.30\%$ $-4.71\%$ $-7.08\%$ $-2.49\%$ $+0.44\%$
-Claude Series Claude-3.5-Haiku-20241022 $71.62$ $75.16$ $-16.06\%$ $-29.72\%$ $-9.95\%$ $-11.67\%$ $-9.77\%$ $-6.94\%$ $-11.69\%$ $-18.33\%$
-Claude-3.5-Sonnet-20241022 $83.75$ $89.77$ $-4.33\%$ $-12.56\%$ $-5.97\%$ $-4.19\%$ $-4.48\%$ $-9.30\%$ $-6.12\%$ $-10.70\%$
-Gemini Series Gemini-1.5-Flash $74.62$ $76.20$ $-17.09\%$ $-20.27\%$ $-13.23\%$ $-35.62\%$ $-34.00\%$ $-55.62\%$ $-2.01\%$ $-2.19\%$
-Gemini-2.0-Flash $87.25$ $92.28$ $-5.01\%$ $-8.14\%$ $-6.02\%$ $-6.79\%$ $-7.59\%$ $-24.89\%$ $-2.29\%$ $-1.58\%$
-Gemini-1.5-Pro-002 $84.50$ $91.65$ $-6.51\%$ $-8.88\%$ $-10.80\%$ $-13.67\%$ $-11.98\%$ $-18.91\%$ $-4.44\%$ $-2.96\%$
-DeepSeek DeepSeek-V3 $88.00$ $92.07$ $-7.39\%$ $-14.06\%$ $-13.92\%$ $-23.81\%$ $-15.77\%$ $-23.58\%$ $-4.26\%$ $-2.27\%$
-LLaMA Series LLaMA-3.1-8B-Instruct $46.75$ $41.34$ $-28.34\%$ $-26.26\%$ $-19.52\%$ $-23.74\%$ $-22.19\%$ $-20.20\%$ $-5.61\%$ $-14.65\%$
-LLaMA-3.1-70B-Instruct $70.75$ $67.22$ $-15.19\%$ $-22.67\%$ $-18.90\%$ $-30.75\%$ $-24.56\%$ $-41.93\%$ $-5.65\%$ $-8.70\%$
-LLaMA-3.1-405B-Instruct $78.88$ $77.66$ $-9.83\%$ $-19.89\%$ $-7.29\%$ $-15.32\%$ $-10.14\%$ $-24.73\%$ $-3.96\%$ $-13.44\%$
-LLaMA-3.3-70B-Instruct $76.12$ $78.08$ $-8.21\%$ $-15.24\%$ $-5.42\%$ $-11.50\%$ $-5.75\%$ $-19.79\%$ $-1.81\%$ $-9.63\%$
-Qwen Series Qwen2.5-7B-Instruct $57.25$ $59.71$ $-17.47\%$ $-28.67\%$ $-23.80\%$ $-18.88\%$ $-30.35\%$ $-24.13\%$ $-7.42\%$ $-6.29\%$
-Qwen2.5-14B-Instruct $67.25$ $76.41$ $-20.26\%$ $-27.32\%$ $-36.06\%$ $-34.15\%$ $-43.49\%$ $-43.17\%$ $-11.71\%$ $-9.56\%$
-Qwen2.5-32B-Instruct $77.62$ $82.46$ $-14.01\%$ $-22.78\%$ $-16.26\%$ $-22.28\%$ $-22.38\%$ $-27.09\%$ $-6.60\%$ $-3.29\%$
-Qwen2.5-72B-Instruct $78.25$ $90.40$ $-8.31\%$ $-18.94\%$ $-7.67\%$ $-5.54\%$ $-16.61\%$ $-15.70\%$ $-5.27\%$ $-8.55\%$
-Qwen2.5-Coder-32B-Instruct $84.12$ $85.80$ $-10.10\%$ $-22.14\%$ $-16.79\%$ $-22.63\%$ $-13.97\%$ $-29.68\%$ $-8.02\%$ $-11.44\%$
-Average $-11.86\%$ $-18.99\%$ $-13.29\%$ $-17.57\%$ $-17.11\%$ $-23.54\%$ $-5.45\%$ $-7.35\%$
-$\mathbf{-14.53}\%$ $\mathbf{-14.90}\%$ $\mathbf{-19.52}\%$ $\mathbf{-6.16}\%$
-Table 16: Absolute Pass@1 degradation of LLMs under all perturbation types for output prediction tasks across Crux and LCB in CoT inference ($N=1$). [To Relative Version]
-Model Series Model Name VAN ALL MDC MPS MHC
+GPT-4 Omni GPT-4o-Mini $79.9$ $84.1$ $10.8\%$ $16.9\%$ $10.8\%$ $10.9\%$ $13.1\%$ $7.4\%$ $3.3\%$ $1.7\%$ $9.4\%$
+GPT-4o $90.2$ $94.4$ $2.8\%$ $8.4\%$ $3.6\%$ $7.3\%$ $4.7\%$ $7.1\%$ $2.5\%$ $0.4\%$ $4.2\%$
+Claude Series Claude-3.5-Haiku-20241022 $71.6$ $75.2$ $16.1\%$ $29.7\%$ $9.9\%$ $11.7\%$ $9.8\%$ $6.9\%$ $11.7\%$ $18.3\%$ $13.7\%$
+Claude-3.5-Sonnet-20241022 $83.8$ $89.8$ $4.3\%$ $12.6\%$ $6.0\%$ $4.2\%$ $4.5\%$ $9.3\%$ $6.1\%$ $10.7\%$ $6.7\%$
+Gemini Series Gemini-1.5-Flash $74.6$ $76.2$ $17.1\%$ $20.3\%$ $13.2\%$ $35.6\%$ $34.0\%$ $55.6\%$ $0\% 2.2$$2\%$ $21.0\%$
+Gemini-2.0-Flash $87.2$ $92.3$ $5.0\%$ $8.1\%$ $6.0\%$ $6.8\%$ $7.6\%$ $24.9\%$ $2.3\%$ $1.6\%$ $7.1\%$
+Gemini-1.5-Pro-002 $84.5$ $91.6$ $6.5\%$ $8.9\%$ $10.8\%$ $13.7\%$ $12.0\%$ $18.9\%$ $4.4\%$ $3.0\%$ $9.4\%$
+DeepSeek DeepSeek-V3 $88.0$ $92.1$ $7.4\%$ $14.1\%$ $13.9\%$ $23.8\%$ $15.8\%$ $23.6\%$ $4.3\%$ $2.3\%$ $12.4\%$
+LLaMA Series LLaMA-3.1-8B-Instruct $46.8$ $41.3$ $28.3\%$ $26.3\%$ $19.5\%$ $3.7\% 22$$2.2\%$ $20.2\%$ $5.6\%$ $14.6\%$ $19.8\%$
+LLaMA-3.1-70B-Instruct $70.8$ $67.2$ $15.2\%$ $22.7\%$ $18.9\%$ $30.7\%$ $24.6\%$ $41.9\%$ $5.7\%$ $8.7\%$ $19.8\%$
+LLaMA-3.1-405B-Instruct $78.9$ $77.7$ $9.8\%$ $19.9\%$ $7.3\%$ $15.3\%$ $10.1\%$ $24.7\%$ $4.0\%$ $3.4\% 11$$1.8\%$
+LLaMA-3.3-70B-Instruct $76.1$ $78.1$ $8.2\%$ $15.2\%$ $5.4\%$ $11.5\%$ $5.7\%$ $19.8\%$ $1.8\%$ $9.6\%$ $8.6\%$
+Qwen Series Qwen2.5-7B-Instruct $57.2$ $59.7$ $17.5\%$ $28.7\%$ $23.8\%$ $18.9\%$ $30.3\%$ $24.1\%$ $7.4\%$ $6.3\%$ $19.7\%$
+Qwen2.5-14B-Instruct $67.2$ $76.4$ $20.3\%$ $27.3\%$ $36.1\%$ $34.2\%$ $43.5\%$ $43.2\%$ $11.7\%$ $9.6\%$ $28.1\%$
+Qwen2.5-32B-Instruct $77.6$ $82.5$ $14.0\%$ $22.8\%$ $16.3\%$ $2.3\% 22$$2.4\%$ $27.1\%$ $6.6\%$ $3.3\%$ $16.3\%$
+Qwen2.5-72B-Instruct $78.2$ $90.4$ $8.3\%$ $18.9\%$ $7.7\%$ $5.5\%$ $16.6\%$ $15.7\%$ $5.3\%$ $8.5\%$ $10.5\%$
+Qwen2.5-Coder-32B-Instruct $84.1$ $85.8$ $10.1\%$ $22.1\%$ $16.8\%$ $22.6\%$ $14.0\%$ $29.7\%$ $8.0\%$ $11.4\%$ $15.7\%$
+Average $11.9\%$ $19.0\%$ $13.3\%$ $17.6\%$ $17.1\%$ $23.5\%$ $5.5\%$ $7.3\%$
+$\mathbf{-14.5}\%$ $\mathbf{-14.9}\%$ $\mathbf{-19.5}\%$ $\mathbf{-6.2}\%$ $\mathbf{-13.8}\%$
+Table 16: Absolute Pass@1 degradation of LLMs under all perturbations for output prediction tasks across Crux and LCB in CoT inference ($N=1$). [To Relative Version].
+Model Series Model Name VAN PSC-ALL MCC MPS MHC Average
 Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
-GPT-4 Omni GPT-4o-Mini $79.88$ $84.13$ $-8.62$ $-14.20$ $-8.62$ $-9.19$ $-10.50$ $-6.26$ $-2.62$ $-1.46$
-GPT-4o $90.25$ $94.36$ $-2.50$ $-7.93$ $-3.25$ $-6.89$ $-4.25$ $-6.68$ $-2.25$ $+0.42$
-Claude Series Claude-3.5-Haiku-20241022 $71.62$ $75.16$ $-11.50$ $-22.34$ $-7.12$ $-8.77$ $-7.00$ $-5.22$ $-8.38$ $-13.78$
-Claude-3.5-Sonnet-20241022 $83.75$ $89.77$ $-3.62$ $-11.27$ $-5.00$ $-3.76$ $-3.75$ $-8.35$ $-5.12$ $-9.60$
-Gemini Series Gemini-1.5-Flash $74.62$ $76.20$ $-12.75$ $-15.45$ $-9.88$ $-27.14$ $-25.38$ $-42.38$ $-1.50$ $-1.67$
-Gemini-2.0-Flash $87.25$ $92.28$ $-4.38$ $-7.52$ $-5.25$ $-6.26$ $-6.62$ $-22.96$ $-2.00$ $-1.46$
-Gemini-1.5-Pro-002 $84.50$ $91.65$ $-5.50$ $-8.14$ $-9.12$ $-12.53$ $-10.12$ $-17.33$ $-3.75$ $-2.71$
-DeepSeek DeepSeek-V3 $88.00$ $92.07$ $-6.50$ $-12.94$ $-12.25$ $-21.92$ $-13.88$ $-21.71$ $-3.75$ $-2.09$
-LLaMA Series LLaMA-3.1-8B-Instruct $46.75$ $41.34$ $-13.25$ $-10.86$ $-9.12$ $-9.81$ $-10.38$ $-8.35$ $-2.62$ $-6.05$
-LLaMA-3.1-70B-Instruct $70.75$ $67.22$ $-10.75$ $-15.24$ $-13.38$ $-20.67$ $-17.38$ $-28.18$ $-4.00$ $-5.85$
-LLaMA-3.1-405B-Instruct $78.88$ $77.66$ $-7.75$ $-15.45$ $-5.75$ $-11.90$ $-8.00$ $-19.21$ $-3.12$ $-10.44$
-LLaMA-3.3-70B-Instruct $76.12$ $78.08$ $-6.25$ $-11.90$ $-4.12$ $-8.98$ $-4.38$ $-15.45$ $-1.38$ $-7.52$
-Qwen Series Qwen2.5-7B-Instruct $57.25$ $59.71$ $-10.00$ $-17.12$ $-13.62$ $-11.27$ $-17.38$ $-14.41$ $-4.25$ $-3.76$
-Qwen2.5-14B-Instruct $67.25$ $76.41$ $-13.62$ $-20.88$ $-24.25$ $-26.10$ $-29.25$ $-32.99$ $-7.88$ $-7.31$
-Qwen2.5-32B-Instruct $77.62$ $82.46$ $-10.88$ $-18.79$ $-12.62$ $-18.37$ $-17.38$ $-22.34$ $-5.12$ $-2.71$
-Qwen2.5-72B-Instruct $78.25$ $90.40$ $-6.50$ $-17.12$ $-6.00$ $-5.01$ $-13.00$ $-14.20$ $-4.12$ $-7.72$
-Qwen2.5-Coder-32B-Instruct $84.12$ $85.80$ $-8.50$ $-19.00$ $-14.12$ $-19.42$ $-11.75$ $-25.47$ $-6.75$ $-9.81$
-Average $-8.40$ $-14.48$ $-9.62$ $-13.41$ $-12.38$ $-18.32$ $-4.04$ $-5.50$
-$\mathbf{-10.68}$ $\mathbf{-11.04}$ $\mathbf{-14.60}$ $\mathbf{-4.58}$
-D.4 Impact on Input Prediction
+GPT-4 Omni GPT-4o-Mini $79.9$ $84.1$ 8.6-8.6 14.2-14.2 8.6-8.6 9.2-9.2 10.5-10.5 6.3-6.3 2.6-2.6 1.5-1.5 7.7-7.7
+GPT-4o $90.2$ $94.4$ 2.5-2.5 7.9-7.9 3.2-3.2 6.9-6.9 4.2-4.2 6.7-6.7 2.2-2.2 +0.4+0.4 3.9-3.9
+Claude Series Claude-3.5-Haiku-20241022 $71.6$ $75.2$ 11.5-11.5 22.3-22.3 7.1-7.1 8.8-8.8 7.0-7.0 5.2-5.2 8.4-8.4 13.8-13.8 10.0-10.0
+Claude-3.5-Sonnet-20241022 $83.8$ $89.8$ 3.6-3.6 11.3-11.3 5.0-5.0 3.8-3.8 3.8-3.8 8.4-8.4 5.1-5.1 9.6-9.6 5.8-5.8
+Gemini Series Gemini-1.5-Flash $74.6$ $76.2$ 12.8-12.8 15.4-15.4 9.9-9.9 27.1-27.1 25.4-25.4 42.4-42.4 1.5-1.5 1.7-1.7 15.9-15.9
+Gemini-2.0-Flash $87.2$ $92.3$ 4.4-4.4 7.5-7.5 5.2-5.2 6.3-6.3 6.6-6.6 23.0-23.0 2.0-2.0 1.5-1.5 6.4-6.4
+Gemini-1.5-Pro-002 $84.5$ $91.6$ 5.5-5.5 8.1-8.1 9.1-9.1 12.5-12.5 10.1-10.1 17.3-17.3 3.8-3.8 2.7-2.7 8.3-8.3
+DeepSeek DeepSeek-V3 $88.0$ $92.1$ 6.5-6.5 12.9-12.9 12.2-12.2 21.9-21.9 13.9-13.9 21.7-21.7 3.8-3.8 2.1-2.1 11.2-11.2
+LLaMA Series LLaMA-3.1-8B-Instruct $46.8$ $41.3$ 13.2-13.2 10.9-10.9 9.1-9.1 9.8-9.8 10.4-10.4 8.4-8.4 2.6-2.6 6.1-6.1 8.8-8.8
+LLaMA-3.1-70B-Instruct $70.8$ $67.2$ 10.8-10.8 15.2-15.2 13.4-13.4 20.7-20.7 17.4-17.4 28.2-28.2 4.0-4.0 5.8-5.8 13.7-13.7
+LLaMA-3.1-405B-Instruct $78.9$ $77.7$ 7.8-7.8 15.4-15.4 5.8-5.8 11.9-11.9 8.0-8.0 19.2-19.2 3.1-3.1 10.4-10.4 9.2-9.2
+LLaMA-3.3-70B-Instruct $76.1$ $78.1$ 6.2-6.2 11.9-11.9 4.1-4.1 9.0-9.0 4.4-4.4 15.4-15.4 1.4-1.4 7.5-7.5 6.6-6.6
+Qwen Series Qwen2.5-7B-Instruct $57.2$ $59.7$ 10.0-10.0 17.1-17.1 13.6-13.6 11.3-11.3 17.4-17.4 14.4-14.4 4.2-4.2 3.8-3.8 11.4-11.4
+Qwen2.5-14B-Instruct $67.2$ $76.4$ 13.6-13.6 20.9-20.9 24.2-24.2 26.1-26.1 29.2-29.2 33.0-33.0 7.9-7.9 7.3-7.3 19.9-19.9
+Qwen2.5-32B-Instruct $77.6$ $82.5$ 10.9-10.9 18.8-18.8 12.6-12.6 18.4-18.4 17.4-17.4 22.3-22.3 5.1-5.1 2.7-2.7 13.0-13.0
+Qwen2.5-72B-Instruct $78.2$ $90.4$ 6.5-6.5 17.1-17.1 6.0-6.0 5.0-5.0 13.0-13.0 14.2-14.2 4.1-4.1 7.7-7.7 8.8-8.8
+Qwen2.5-Coder-32B-Instruct $84.1$ $85.8$ 8.5-8.5 19.0-19.0 14.1-14.1 19.4-19.4 11.8-11.8 25.5-25.5 6.8-6.8 9.8-9.8 13.3-13.3
+Average 8.4-8.4 14.5-14.5 9.6-9.6 13.4-13.4 12.4-12.4 18.3-18.3 4.0-4.0 5.5-5.5
+$\mathbf{-10.7}$ $\mathbf{-11.0}$ $\mathbf{-14.6}$ $\mathbf{-4.6}$ $\mathbf{-10.2}$
+E.3 Additional Performance Degradation Tables for Input Prediction
 In §3.5, we evaluate LLM robustness on input prediction using the Crux dataset.
-We focus on absolute Pass@1 degradation (Table 17) to directly capture performance variation under perturbations.
-For completeness, we also report relative degradation in Table 5.
-Table 17: Comparison of absolute Pass@1 degradation under Crux for input prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Relative Version]
-Model Series Model Name VAN ALL MDC MPS MHC
+We focus on relative Pass@1 degradation (Table 5) to capture performance variation under perturbations.
+For completeness, we also report absolute degradation in Table 17.
+Table 17: Absolute Pass@1 degradation for input prediction tasks across different reasoning modes on Crux. [To Relative Version].
+Model Series Model Name VAN PSC-ALL MCC MPS MHC
 Direct CoT Direct CoT Direct CoT Direct CoT Direct CoT
-GPT-4 Omni GPT-4o-Mini $57.67$ $69.52$ $-17.21$ $-13.65$ $+0.75$ $+0.73$ $+1.00$ $+0.73$ $-17.58$ $-6.02$
-GPT-4o $68.04$ $79.12$ $-9.04$ $-5.88$ $+3.83$ $+1.00$ $+0.96$ $+2.12$ $-5.67$ $-2.25$
-Claude Series Claude-3.5-Haiku-20241022 $54.67$ $63.62$ $-9.29$ $-10.62$ $+3.83$ $+2.62$ $+4.25$ $+1.50$ $-24.79$ $-19.00$
-Claude-3.5-Sonnet-20241022 $73.33$ $80.12$ $-11.67$ $-7.75$ $+0.08$ $-2.75$ $-1.25$ $-2.12$ $-18.21$ $-13.62$
-Gemini Series Gemini-1.5-Flash $57.62$ $73.62$ $-17.38$ $-17.50$ $+5.88$ $-4.00$ $+3.46$ $-3.12$ $-17.83$ $-8.12$
-Gemini-2.0-Flash $70.50$ $84.88$ $-29.12$ $-11.38$ $-0.54$ $-2.38$ $-8.08$ $-10.25$ $-15.17$ $-3.62$
-Gemini-1.5-Pro-002 $70.87$ $81.88$ $-14.79$ $-12.25$ $-1.75$ $-5.75$ $-2.67$ $-7.88$ $-13.25$ $-6.25$
-DeepSeek DeepSeek-V3 $69.08$ $82.88$ $-11.46$ $-6.62$ $+0.77$ $-0.25$ $+1.08$ $+0.75$ $-3.88$ $+2.62$
-LLaMA Series LLaMA-3.1-8B-Instruct $37.12$ $41.62$ $-11.33$ $-25.75$ $+1.44$ $+0.62$ $+1.13$ $-1.50$ $-2.67$ $-10.38$
-LLaMA-3.1-70B-Instruct $62.00$ $66.38$ $-17.54$ $-11.87$ $+3.04$ $+3.00$ $+1.00$ $+1.12$ $-10.79$ $-6.75$
-LLaMA-3.1-405B-Instruct $66.79$ $75.00$ $-10.21$ $-6.75$ $+3.92$ $+2.75$ $+2.00$ $+1.75$ $-5.21$ $-4.25$
-LLaMA-3.3-70B-Instruct $63.38$ $76.50$ $-18.62$ $-12.00$ $+0.79$ $-0.25$ $-2.58$ $+0.50$ $-13.83$ $-8.50$
-Qwen Series Qwen2.5-7B-Instruct $38.83$ $51.38$ $-17.25$ $-25.88$ $+5.71$ $+1.62$ $+4.75$ $+3.87$ $-3.63$ $-7.38$
-Qwen2.5-14B-Instruct $50.54$ $60.75$ $-19.42$ $-9.13$ $+4.38$ $+6.62$ $+2.46$ $+3.12$ $-4.42$ $-0.75$
-Qwen2.5-32B-Instruct $63.50$ $74.12$ $-16.54$ $-15.12$ $+4.92$ $+0.38$ $+1.42$ $-3.00$ $-13.08$ $-8.12$
-Qwen2.5-72B-Instruct $64.46$ $73.81$ $-17.31$ $-7.29$ $+1.71$ $+1.19$ $-2.38$ $+1.81$ $-16.46$ $-5.06$
-Qwen2.5-Coder-32B-Instruct $74.25$ $78.25$ $-21.21$ $-16.12$ $-0.75$ 0.00 $-1.79$ $+0.25$ $-27.08$ $-26.12$
-Average $\mathbf{-15.85}$ $\mathbf{-12.68}$ $\mathbf{+2.24}$ $\mathbf{+0.30}$ $\mathbf{+0.28}$ $\mathbf{-0.61}$ $\mathbf{-12.56}$ $\mathbf{-7.86}$
-Appendix E Case Studies
-E.1 Case Study 1: How does DeepSeek-V3 be misled by ALL
+GPT-4 Omni GPT-4o-Mini $57.8$ $69.5$ 17.3-17.3 13.6-13.6 +0.7+0.7 +0.7+0.7 +0.8+0.8 +0.7+0.7 17.7-17.7 6.0-6.0
+GPT-4o $68.0$ $79.1$ 9.0-9.0 5.9-5.9 +3.8+3.8 +1.0+1.0 +1.0+1.0 +2.1+2.1 5.7-5.7 2.2-2.2
+Claude Series Claude-3.5-Haiku-20241022 $54.7$ $63.6$ 9.3-9.3 10.6-10.6 +3.7+3.7 +2.6+2.6 +4.2+4.2 +1.5+1.5 24.8-24.8 19.0-19.0
+Claude-3.5-Sonnet-20241022 $73.3$ $80.1$ 11.7-11.7 7.8-7.8 +0.1+0.1 2.8-2.8 1.3-1.3 2.1-2.1 18.2-18.2 13.6-13.6
+Gemini Series Gemini-1.5-Flash $57.6$ $73.6$ 17.2-17.2 17.5-17.5 +5.9+5.9 4.0-4.0 +3.5+3.5 3.1-3.1 17.8-17.8 8.1-8.1
+Gemini-2.0-Flash $70.5$ $84.9$ 29.7-29.7 11.4-11.4 0.5-0.5 2.4-2.4 8.1-8.1 10.2-10.2 15.2-15.2 3.6-3.6
+Gemini-1.5-Pro-002 $71.0$ $81.9$ 14.9-14.9 12.1-12.1 1.8-1.8 5.8-5.8 2.7-2.7 7.9-7.9 13.4-13.4 6.2-6.2
+DeepSeek DeepSeek-V3 $69.1$ $82.9$ 11.5-11.5 6.6-6.6 +0.7+0.7 0.2-0.2 +1.2+1.2 +0.8+0.8 4.0-4.0 +2.6+2.6
+LLaMA Series LLaMA-3.1-8B-Instruct $37.1$ $41.6$ 11.4-11.4 25.8-25.8 +1.3+1.3 +0.6+0.6 +1.1+1.1 1.5-1.5 2.7-2.7 10.4-10.4
+LLaMA-3.1-70B-Instruct $62.1$ $66.4$ 17.6-17.6 11.8-11.8 +2.9+2.9 +3.0+3.0 +1.0+1.0 +1.2+1.2 10.8-10.8 6.6-6.6
+LLaMA-3.1-405B-Instruct $66.8$ $75.0$ 10.2-10.2 6.8-6.8 +3.9+3.9 +2.8+2.8 +2.0+2.0 +1.9+1.9 5.3-5.3 4.2-4.2
+LLaMA-3.3-70B-Instruct $63.3$ $76.5$ 18.5-18.5 11.9-11.9 +0.8+0.8 0.2-0.2 2.5-2.5 +0.5+0.5 13.8-13.8 8.5-8.5
+Qwen Series Qwen2.5-7B-Instruct $38.8$ $51.4$ 17.2-17.2 25.9-25.9 +5.7+5.7 +1.7+1.7 +4.8+4.8 +3.9+3.9 3.7-3.7 7.4-7.4
+Qwen2.5-14B-Instruct $50.4$ $60.8$ 19.3-19.3 9.1-9.1 +4.3+4.3 +6.6+6.6 +2.7+2.7 +3.2+3.2 4.3-4.3 0.8-0.8
+Qwen2.5-32B-Instruct $63.5$ $74.1$ 16.5-16.5 15.1-15.1 +4.7+4.7 +0.4+0.4 +1.4+1.4 3.0-3.0 13.1-13.1 8.1-8.1
+Qwen2.5-72B-Instruct $64.6$ $74.1$ 17.5-17.5 8.0-8.0 +1.7+1.7 +0.9+0.9 2.4-2.4 +1.5+1.5 16.5-16.5 5.4-5.4
+Qwen2.5-Coder-32B-Instruct $74.2$ $78.4$ 21.2-21.2 16.1-16.1 0.8-0.8 0.1-0.1 1.8-1.8 +0.1+0.1 27.1-27.1 26.2-26.2
+Average $\mathbf{-15.9}$ $\mathbf{-12.6}$ $\mathbf{+2.2}$ $\mathbf{+0.3}$ $\mathbf{+0.3}$ $\mathbf{-0.6}$ $\mathbf{-12.6}$ $\mathbf{-7.8}$
+Appendix F Case Studies
+F.1 Case Study 1: How is DeepSeek-V3 misled by PSC-ALL
 This case (Code 19 and 20) demonstrates how garbage code can distort model reasoning.
@@
 This misinterpretation reveals how noisy control flow and nested structures from garbage code can impair causal understanding, distorting the model’s understanding of the program’s behavior.
-Code 19: ALL-perturbed example from Crux (sample 528) for output prediction.
+Code 19: PSC-ALL-perturbed example from Crux (sample 528) for output prediction.
 1Var_1 = 17
@@
 Thus, the output value for f(’papeluchis’)‘ is ‘0‘, then output your prediction [ANSWER] f(’papeluchis’) == 0 [/ANSWER].
-E.2 Case Study 2: How does GPT-4o be misled by MDC
+F.2 Case Study 2: How is GPT-4o misled by MCC
 This case (Code 22) reveals that GPT-4o implicitly relies on NL information.
@@
 Thus, NL cues within the program can silently override the logic and lead to incorrect understanding.
-Code 22: MDC-perturbed example from Crux (sample 1) for output prediction.
+Code 22: MCC-perturbed example from Crux (sample 1) for output prediction.
 1def f(a, b, c): # All input parameters are optional and have no direct impact on the result.
@@
 [ANSWER] f((1, ), (1, ), (1, 2)) == {} [/ANSWER]
-E.3 Case Study 3: How does Qwen2.5-72B-Ins be misled by MPS
+F.3 Case Study 3: How is Qwen2.5-72B-Ins misled by MPS
 This case (Code 24) highlights how MPS can subtly manipulate model reasoning.
-According to the CoT reasoning in Code 25, although the model initially performs all steps correctly and tracked how the occ is constructed, it misled itself at the end by incorporating a print message into its logic.
-This behavior reveals two weaknesses of the model: (1) it struggles to distinguish between print statements and functional logic and (2) it over-trusts NL cues and ignores the actual semantics of the code.
+According to the CoT reasoning in Code 25, although the model initially performs all steps correctly and tracks how the occ is constructed, it misleads itself at the end by incorporating a print message into its logic.
+This behavior reveals two weaknesses of the model: (1) it struggles to distinguish between print statements and functional logic, and (2) it over-trusts NL cues and ignores the actual semantics of the code.
 Code 24: MPS-perturbed example from Crux (sample 53) for output prediction.
@@
 [ANSWER] f("URW rNB") == [] [/ANSWER]
-E.4 Case Study 4: How does MHC override Claude-3.5-Sonnet’s answer
+F.4 Case Study 4: How does MHC override Claude-3.5-Sonnet’s answer
 This case (Code 26) demonstrates how reasoning-level perturbations (MHC) can subtly mislead Claude-3.5-Sonnet’s final output.
@@
 [ANSWER] alternatingSubarray(nums = [2, 3, 4, 3, 4]) == 4 [/ANSWER]
-E.5 Case Study 5: How does GPT-4o gain from reasoning shortcuts
+F.5 Case Study 5: How does GPT-4o gain from reasoning shortcuts
 In §3.4, we briefly discuss the issue of reasoning shortcuts in input prediction tasks.
 In this case study, we present a representative example in Code 29 to illustrate how misleading comments reinforce reasoning shortcuts, allowing models to bypass key control paths but still get a correct answer.
-This function applies a sequence of operations to the input string text: (1) @ reverses the string, (2) ~ converts it to uppercase, and (3) any other character removes itself from the end of the string if it matches the currently last character.
-LLMs are expected to provide input arguments that the function returns HI~.
-In the vanilla thought, as shown in Code 30, GPT-4o mentions, “Text is hi~, ~ makes it HI~, @ reverses it to ~IH ”, it wrongly assumes ~IH is the final output because it applies the reversal after the uppercasing step, resulting in an incorrect answer.
-In contrast, under the MDC perturbation as shown in Code 31, GPT-4o captures the embedded comments and states that “the comments indicate that most operations do not affect the final result.” Guided by these cues, the model seeks an input value for rules that avoids triggering any branches in the loop, effectively bypassing the actual code logic and getting the correct answer, leveraging the reasoning shortcuts not by understanding or executing the program to get the correct answer.
+This function applies a sequence of operations to the input string text: (1) @ reverses the string, (2) ˜ converts it to uppercase, and (3) any other character removes itself from the end of the string if it matches the currently last character.
+LLMs are expected to provide input arguments that the function returns HI˜.
+In the vanilla reasoning, as shown in Code 30, GPT-4o mentions, “Text is hi˜, ˜ makes it HI˜, @ reverses it to ˜IH ”, it wrongly assumes ˜IH is the final output because it applies the reversal after the uppercasing step, resulting in an incorrect answer.
+In contrast, under the MCC perturbation as shown in Code 31, GPT-4o captures the embedded comments and states that “the comments indicate that most operations do not affect the final result.” Guided by these cues, the model seeks an input value for rules that avoids triggering any branches in the loop, effectively bypassing the actual code logic and getting the correct answer by leveraging reasoning shortcuts rather than truly understanding or executing the program.
 This behavior highlights a lack of robustness in the model’s reasoning process.
@@
 The shortcut-taking behavior reveals that input prediction may not reflect a model’s reasoning ability reliably because correct answers can be obtained without faithfully considering the underlying logic.
-Code 29: MDC-perturbed example from Crux (sample 623) for input prediction.
+Code 29: MCC-perturbed example from Crux (sample 623) for input prediction.
 1def f(text, rules): # The inputs have no impact on the final result.
@@
 [ANSWER] f(’HI~’, []) == HI~’ [/ANSWER]
-E.6 Case Study 6: How does QwQ-32B Collapse
-In §5.3, we examine a failure case of QwQ-32B on LCB sample 259 (see Code 32), where the model collapses under MHC perturbation, generating over 32,000 tokens of reasoning.
-We manually segmented the reasoning process into six stages (see Figure 4):
+F.6 Case Study 6: How does QwQ-32B collapse
+In §3.4, we examine a failure case of QwQ-32B on LCB sample 259 (see Code 32), where the model collapses under MHC perturbation, generating over 32k tokens of reasoning.
+Figure 3 shows a quadratic growth of confusing tokens before the collapse.
+We manually segment the reasoning process into six stages (see Figure 3):
   • In Stage 1, the model executes the two for loops step-by-step, demonstrates stable reasoning with occasional self-reflection, and correctly predicts the output as ‘-1.’
@@
 16""" Suggesting answer: minimumSum(nums = [6, 5, 4, 3, 4, 5]) == -1 """
-Figure 4: Breakdown of QwQ-32B’s reasoning trajectory under MHC perturbation on LCB sample 259.
-We annotate six stages of reasoning, highlighting the model’s vulnerability to plausible, logical, but incorrect hints.
-Appendix F Limitation
+Appendix G Limitation
 This study is subject to several limitations.
-Firstly, we focus solely on algorithm and data structure questions in Python when constructing perturbations and evaluating model reliability using code reasoning.
-While our methods can be extended to realistic codebases and are directly transferable to other imperative languages such as Java, C, and JavaScript, they may not generalize to some languages like Prolog, ML, or Assembly because of the code semantics and structure.
-Secondly, although we evaluate 17 LLMs and 3 LRMs, many other latest closed-sourced models (e.g., GPT-4.5, o3, o4-Mini, Gemini-2.5, and Claude-3.7) and proprietary open-sourced models (e.g., Mixtral, Falcon, Vicuna) are not included due to budget constraints.
-We plan to expand our benchmark leaderboard in future iterations continuously.
+Firstly, we mainly focus on algorithms and data structure questions in Python when constructing perturbations and evaluating model reliability using code reasoning.
+While our methods can be extended to realistic codebases and are directly transferable to other imperative languages, such as C/C++ and Java, they may not generalize to some languages like Prolog, ML, or Assembly because of the code semantics and structure.
+Secondly, although we evaluate 17 LLMs and 5 LRMs, many other latest closed-source models (e.g., GPT-4.5, o3, o4-Mini, Gemini-2.5, and Claude-3.7) and open-source models (e.g., Mixtral, Falcon, Vicuna) are not included due to budget constraints.
 Thirdly, we adopt Pass@1 as the primary evaluation metric.